### v1

# Strong Screening Rules for Group-based SLOPE Models

Abstract
Tuning the regularization parameter in penalized regression models is an expensive task, requiring multiple models to be fit along a path of parameters. Strong screening rules drastically reduce computational costs by lowering the dimensionality of the input prior to fitting. We develop strong screening rules for group-based Sorted L-One Penalized Estimation (SLOPE) models: Group SLOPE and Sparse-group SLOPE. The developed rules are applicable for the wider family of group-based OWL models, including OSCAR. Our experiments on both synthetic and real data show that the screening rules significantly accelerate the fitting process. The screening rules make it accessible for group SLOPE and sparse-group SLOPE to be applied to high-dimensional datasets, particularly those encountered in genetics.
1 Introduction
As the amount of data collected increases, the emergence of high-dimensional data, where the number of features ($p$) is much larger than the number of observations ($n$), is becoming increasingly common in fields ranging from genetics to finance. Performing regression and discovering relevant features on these datasets is a challenging task, as classical statistical methods tend to break down. The most popular approach to meeting this challenge is the Lasso [30], which has given rise to the general penalized regression framework
$\hat{\beta}(\lambda)=\operatorname*{arg\,min}_{\beta\in\mathbb{R}^{p}}\left\{f(\beta)+\lambda J(\beta;v)\right\},$ (1)
where $f$ is the loss function, $J$ is a convex penalty norm, $v\succeq 0$ are penalty parameters, and $\lambda>0$ is the regularization parameter.
A key aspect of fitting a penalized model is to tune the value of the regularization parameter, which determines the level of sparsity in the fitted model. Several approaches exist for tuning this parameter, including cross-validation [8, 15] and exact solution path algorithms [10], but can often be computationally expensive. Generally, the model is fit along an $l$-length path $\lambda_{1}\geq\ldots\geq\lambda_{l}\geq 0$. Finding approaches for speeding up the model fitting process can lead to large computational savings when tuning. Screening rules are popular methods that reduce the dimensions of the input dimensionality by discarding irrelevant features before the fitting process. Discarding features, especially in high-dimensional settings, has a tremendous impact on the computational costs.
Denoting the active set of non-zero parameters at $\lambda_{k+1}$ by $\mathcal{A}_{v}(\lambda_{k+1})=\{i\in\{1,\dots,p\}:\hat{\beta}_{i}(\lambda_{k+1})\neq 0\}$, the goal of a screening rule is to use the solution at $\lambda_{k}$ to recover a screened set of features, $\mathcal{S}_{v}(\lambda_{k+1})$, which is a superset of $\mathcal{A}_{v}(\lambda_{k+1})$. The screened set can then be used as input for calculating the fitted values. By finding the smallest possible screened set, that still contains the active set, the number of irrelevant features included in the optimization process is reduced.
There are two types of screening rules: safe and heuristic. Safe rules, as suggested by the name, provide guarantees that any variables discarded are in fact inactive at the solution. The seminal work of El Ghaoui et al. [11] introduced the Safe Feature Elimination Rule (SAFE) for the lasso, in which safe regions were constructed to eliminate non-active features. Other examples of safe rules are given in [2, 25, 34], including sample screening [28]. Heuristic rules, on the other hand, tend to discard more variables, but this increased screening efficiency can lead to active variables being incorrectly discarded. Examples of heuristic screening rules are given in [1, 13, 31]. Amongst the proposed heuristic rules, of particular interest is the strong screening rule framework developed for the lasso [31]. As strong screening rules can lead to violations, Tibshirani et al. [31] complemented their rules with a check of the Karush–Kuhn–Tucker (KKT) [16] optimality conditions to ensure no violations occur. There are also hybrid screening regimes, which combine safe and heuristic rules [32, 37].
A strong screening rule is formulated through the KKT conditions for Equation 1, given by:
$\mathbf{0}\in\nabla f(\beta)+\lambda\partial J(\beta;v).$ (2)
If the gradient $\nabla f(\hat{\beta}(\lambda_{k+1}))$ was available, the set $\mathcal{A}_{v}(\lambda_{k+1})$ could be identified by checking the subdifferential of the norm at zero. That is, identifying for which variables
$-\nabla f(\beta)\in\lambda\partial J(\mathbf{0};v)=\{x\in\mathbb{R}^{p}:J^{*}(x;\lambda v)\leq 1\},$ (3)
where $J^{*}$ refers to the dual norm of $J$. The subdifferential at zero, $\partial J(\mathbf{0};v)$, can also be seen to be the unit ball of its dual norm. As the gradient at $\lambda_{k+1}$ is not available in practice, an approximation of the gradient is used to find a screened subset of the features, $\mathcal{S}_{v}(\lambda_{k+1})$. The screened variables are combined with the previously active ones to form the reduced input space, $\mathcal{E}_{v}(\lambda_{k+1})=\mathcal{S}_{v}(\lambda_{k+1})\cup\mathcal{A}_{v}(\lambda_{k})$. The fitted values are then calculated by solving Equation 1 using only the variables in $\mathcal{E}_{v}(\lambda_{k+1})$. Optimality of the fitted values is checked for using the KKT conditions (Equation 2), with violations added into $\mathcal{E}_{v}$. This ensures features are not discarded that should be active.
1.1 Contributions
Over the years, several adaptive extensions to the lasso have been proposed, including the Sorted L-One Penalized Estimation (SLOPE) method [4]. SLOPE applies the sorted $\ell_{1}$ norm $J_{\text{slope}}(\beta;v)=\sum_{i=1}^{p}v_{i}|\beta|_{(i)}$, where $v_{1}\geq\ldots\geq v_{p}\geq 0,|\beta|_{(1)}\geq\ldots\geq|\beta|_{(p)}$. One key advantage of SLOPE is its ability to control the variable false discovery rate (FDR) under orthogonal data. Both safe and strong rules have been proposed for SLOPE [3, 12, 17], as well as exact solution path algorithms [9, 24]. As SLOPE is a non-separable penalty, safe screening rules require repeated screening during the fitting process [17]. This can be a computational bottleneck, as the calculation of the safe regions required for screening are often expensive.
Strong screening rules, on the other hand, are usually more time efficient as the extra cost of the screening and KKT checks tend to be outweighed by the significant savings of fitting on a reduced input space [31]. Motivated by this, we introduce strong screening rules for two group-based extensions of SLOPE: Group SLOPE (gSLOPE) [6] and Sparse-group SLOPE (SGS) [14].
Group SLOPE (gSLOPE) is an extension of SLOPE for selecting groups, defined by the norm
$J_{\text{gslope}}(\beta;w)=\sum_{g=1}^{m}\sqrt{p_{g}}w_{g}\|\beta^{(g)}\|_{2},$ (4)
such that $\beta^{(g)}\in\mathbb{R}^{p_{g}}$ is a vector of the group coefficients and $p_{g}$ denotes the group sizes [6]. The norm has ordered penalty weights $w_{1}\geq\ldots\geq w_{g}\geq 0$ (described in Appendix A.1) which are matched to $\sqrt{p_{1}}\|\beta^{(1)}\|_{2}\geq\ldots\geq\sqrt{p_{m}}\|\beta^{(m)}\|_{2}$.
Sparse-group SLOPE (SGS) was proposed as an extension to SLOPE and gSLOPE, applying both variable and group penalization for bi-level selection [14]. For $\alpha\in[0,1]$ with variable weights $v$ and group weights $w$ (described in Appendix B.1), the SGS norm is defined as
$J_{\text{sgs}}(\beta;\alpha,v,w)=\alpha J_{\text{slope}}(\beta;v)+(1-\alpha)J_{\text{gslope}}(\beta;w).$
Both of these approaches were developed for the analysis of features that naturally belong to groups. This situation is frequently encountered in the analysis of genetics data, where genes are grouped into pathways for the completion of a specific biological task. The gSLOPE approach selects groups of features, whereas SGS performs bi-level selection. Both approaches have been shown to inherit SLOPE’s ability to control FDR under orthogonal data: gSLOPE controls the group FDR [6] and SGS controls both variable and group FDRs [14].
Section 2 introduces a sparse-group screening rule framework. Based on this, strong screening rules for gSLOPE and SGS are developed and presented in Sections 3 and 4. The proofs of the theorems and propositions presented in Sections 2 and 3 are provided in the Appendices A.2 and B.3. Through a series of simulations and the analysis of real datasets (Sections 5), the proposed screening rules are shown to significantly reduce computational runtime. Additionally, due to the reduced input dimensionality offered by screening rules, issues with convergence in the analysis of large datasets are shown to be alleviated. These improvements are achieved without affecting the solution optimality.
The screening rules proposed in this manuscript are further applicable to group-based Ordered Weighted $\ell_{1}$ (OWL) models. SLOPE is a particular case of an OWL model through the definition of its weights. As the proposed screening rules only require that the penalty sequences ($v,w$) are ordered, they can be used for any group-based OWL models; for example, the Octagonal Shrinkage and Clustering Algorithm for Regression (OSCAR) model [5]. Results for group-based OSCAR models are provided in Appendix E.
Notation.
The feature space is divided into a set of $m$ non-overlapping groups, where $\mathcal{G}_{1},\dots,\mathcal{G}_{m}$ denote the set of variable indices for each group. The set of active groups is denoted by $\mathcal{A}_{g}=\{i\in\{1,\dots,m\}:\|\hat{\beta}^{(i)}\|_{2}\neq 0\}$ whereas $\mathcal{Z}=\{i\in\{1,\dots,m\}:\|\hat{\beta}^{(i)}\|_{2}=0\}$ presents the inactive groups. The set of variable indices of the active and inactive groups are denoted by $\mathcal{G}_{\mathcal{A}}$ and $\mathcal{G}_{\mathcal{Z}}$, respectively. The cardinality of a vector $x$ is denoted by $\operatorname{card}(x)$. The operators $(\cdot)_{\downarrow}$ and $(\cdot)_{|\downarrow|}$ sort a vector into decreasing and decreasing absolute form. We use $\preceq$ to denote element-wise less than or equal to. The operator $\mathcal{O}(\cdot)$ returns the index of a vector sorted in decreasing absolute form. The cumulative summation operator applied on a vector is denoted by $\text{cumsum}(x)=[x_{1},x_{1}+x_{2},\ldots,\sum_{i=1}^{\operatorname{card}(x)}x_{i}]$.
2 Screening rule framework
Algorithm 1 Sparse-group screening framework
  Input: $\lambda\in\mathbb{R}^{l}$, $\mathbf{X}\in\mathbb{R}^{n\times p},y\in\mathbb{R}^{n}$
  for $k=1$ to $l-1$ do
     $\mathcal{S}_{g}(\lambda_{k+1})$leftarrow group screening on full input
     $\mathcal{S}_{v}(\lambda_{k+1})$leftarrow variable screening on $g\in\mathcal{S}_{g}(\lambda_{k+1})$
     $\mathcal{E}_{v}\leftarrow\mathcal{S}_{v}(\lambda_{k+1})\cup\mathcal{A}_{v}(\lambda_{k})$
     compute $\hat{\beta}_{\mathcal{E}_{v}}(\lambda_{k+1})$
     $\mathcal{K}_{v}\leftarrow$ variable KKT violations on $\hat{\beta}(\lambda_{k+1})$
     while $\operatorname{card}(\mathcal{K}_{v})>0$ do
        $\mathcal{E}_{v}\leftarrow\mathcal{E}_{v}\cup\mathcal{K}_{v}$
        compute $\hat{\beta}_{\mathcal{E}_{v}}(\lambda_{k+1})$
        $\mathcal{K}_{v}\leftarrow$ variable KKT violations on $\hat{\beta}(\lambda_{k+1})$
     end while
  end for
  Output: $\hat{\beta}(\lambda_{1}),\ldots,\hat{\beta}(\lambda_{l})\in\mathbb{R}^{p\times l}$
Sparse-group models apply both variable and group penalization, such that Equation 1 is extended to include an additional convex penalty function. Examples of sparse-group models include the Sparse-group Lasso (SGL) [29] and SGS [14]. Safe screening rules are available for SGL which perform bi-level screening, i.e. both group and variable screening [21, 33], but there are no strong rules which perform bi-level screening. The strong screening rule framework proposed by Tibshirani et al. [31] does not fully extend to sparse-group or non-separable penalties, whilst the strong screening rule derived for SGL in Liang et al. [19] applies only group-level screening.
Framework.
We introduce a new framework (Algorithm 1) for applying strong screening rules to sparse-group penalties that allow for bi-level screening, based on the framework introduced by Tibshirani et al. [31]. In the sparse-group framework, a screened set of groups is computed, $\mathcal{S}_{g}$, from the full input. An additional layer of screening is then performed to compute $\mathcal{S}_{v}$ using $\mathcal{S}_{g}$. This forms the reduced input set for fitting, $\mathcal{E}_{v}$. The aim of finding $\mathcal{E}_{v}$ is the same as for the lasso and SLOPE strong rules, but the framework first discards irrelevant groups. KKT checks are performed on $\mathcal{E}_{v}$ to ensure no violations have occurred. The framework applied to gSLOPE and SGS is presented in Appendix D.
Based on the framework described in Algorithm 1, the screening rules for SGS are derived (Section 4). The screening rule for gSLOPE (Section 3) follows a similar framework, where $\mathcal{S}_{v}$ is taken as all variables within the groups of $\mathcal{S}_{g}$. The KKT checks are then performed only on the groups.
By applying bi-level screening for SGS, a substantially larger proportion of variables are discarded, than if just group screening were applied (Figure 1). Using only group screening for SGS, without the additional variable screening step, resulted in the number of variables in the screened set scaling poorly with increased dimensionality. This demonstrates that bi-level screening effectively manages the scaling of SGS under rising dimensions.
(a) Applied to synthetic data (Section 5.1). Data generated under a linear model for $p=500,5000$. The results are averaged over $100$ repetitions.
(b) Applied to two real genetics datasets (Section 5.2), with groups formed using pathways. The results are averaged over the nine pathway collections.
Figure 1: The proportion of variables in $\mathcal{S}_{v}$ relative to the full input for SGS, shown for group and bi-level screening, plotted as a function of the regularization path, with $95\%$ confidence intervals.
3 Group SLOPE
The strong screening rule for gSLOPE presented in this section is formulated by computing the zero condition of its subdifferential (as per Equation 3) derived in Theorem 3.1. To derive the subdifferential, we define the operator
$[b]_{\mathcal{G},q}:=(p_{1}^{q}\|b^{(1)}\|_{2},\dots,p_{m}^{q}\|b^{(m)}\|_{2})^{\top}.$
In particular, $[b]_{\mathcal{G}_{\mathcal{Z}},-0.5}$ is the operator applied only to the inactive groups using the quotient $q=-0.5$.
Theorem 3.1 (gSLOPE subdifferential).
The subdifferential for gSLOPE is given by
$\partial J_{\text{gslope}}(\beta;w)=\begin{dcases}\{x\in\mathbb{R}^{\operatorname{card}{\mathcal{G}_{\mathcal{Z}}}}:{[x]}_{\mathcal{G}_{\mathcal{Z}},-0.5}\in\partial J_{\text{slope}}(0;w_{\mathcal{Z}})\},&\text{at 0.}\\ w_{g}\sqrt{p_{g}}\frac{\beta^{(g)}}{\|\beta^{(g)}\|_{2}},&\text{otherwise.}\end{dcases}$
The choice of $q=-0.5$ leads to $J_{\text{gslope}}^{*}(x;w)=J_{\text{slope}}^{*}([x]_{\mathcal{G},-0.5})$, which allows the gSLOPE subdifferential to be written in terms of the SLOPE one [6]. Combining the KKT conditions (Equation 2) with the gSLOPE subdifferential (Theorem 3.1) reveals that a group is zero if $h(\lambda):=([\nabla f(\hat{\beta}(\lambda))]_{\mathcal{G},-0.5})_{\downarrow}\in\partial J_{\text{slope}}(\mathbf{0};\lambda w_{\mathcal{Z}})$. Using the subdifferential of SLOPE [17], this is given by
$\text{cumsum}(h(\lambda)-\lambda w_{\mathcal{Z}})\preceq\mathbf{0}.$ (5)
This condition is checked efficiently using the algorithm proposed for the SLOPE strong rule (Algorithm A1) leading to the strong rule for gSLOPE (Proposition 3.2). The algorithm assumes that the indices for the inactive predictors will be ordered last in the input $c$ and the features $|\hat{\beta}|_{\downarrow}$ [17].
Proposition 3.2 (Strong screening rule for gSLOPE).
Taking $c=h(\lambda_{k+1})$ and $\phi=\lambda_{k+1}w$ as inputs for Algorithm A1 returns a superset $\mathcal{S}_{g}(\lambda_{k+1})$ of the active set $\mathcal{A}_{g}(\lambda_{k+1})$.
The gradient at path value $k+1$ is not available for the computation of $h(\lambda_{k+1})$, so an approximation is required that does not lead to any violations in Algorithm A1. By the cumsum condition in this algorithm, an approximation for a group $g$ is sought such that $h_{g}(\lambda_{k+1})\leq h_{g}(\lambda_{k})+R_{g}$, where $R_{g}\geq 0$ needs to be determined. An approximation is found by assuming that $h_{g}(\lambda_{k+1})$ is a Lipschitz function of $\lambda_{k+1}$ with respect to the $\ell_{1}$ norm, that is,
$\left|h_{g}(\lambda_{k+1})-h_{g}(\lambda_{k})\right|\leq w_{g}|\lambda_{k+1}-\lambda_{k}|.$
By again noting that $J_{\text{gslope}}^{*}(x)=J_{\text{slope}}^{*}([x]_{\mathcal{G},-0.5})$, it can be seen that the assumption is equivalent to those used to derive strong rules for the lasso and SLOPE [17, 31]. By the reverse triangle inequality,
$|h_{g}(\lambda_{k+1})|\leq|h_{g}(\lambda_{k})|+\lambda_{k}w_{g}-\lambda_{k+1}w_{g},$
leading to the choice $R_{g}=\lambda_{k}w_{g}-\lambda_{k+1}w_{g}$ and the gradient approximation strong screening rule.
Proposition 3.3 (Gradient approximation strong screening rule for gSLOPE).
Taking $c=h(\lambda_{k})+\lambda_{k}w-\lambda_{k+1}w$ and $\phi=\lambda_{k+1}w$ as inputs for Algorithm A1, and assuming that for any $k\in\{1,\dots,l-1\}$,
$\left|h_{g}(\lambda_{k+1})-h_{g}(\lambda_{k})\right|\leq w_{g}|\lambda_{k+1}-\lambda_{k}|,\;\forall g=1,\dots,m,$
and $\mathcal{O}(h(\lambda_{k+1}))=\mathcal{O}(h(\lambda_{k}))$, then the algorithm returns a superset $\mathcal{S}_{g}(\lambda_{k+1})$ of the active set $\mathcal{A}_{g}(\lambda_{k+1})$.
4 Sparse-group SLOPE
This section presents the group and variable screening rules for SGS. They are derived using the SGS KKT conditions, formulated in terms of SLOPE and gSLOPE (by the sum rule of subdifferentials):
$-\nabla f(\beta)\in\lambda\alpha\partial J_{\text{slope}}(\beta;v)+\lambda(1-\alpha)\partial J_{\text{gslope}}(\beta;w).$ (6)
4.1 Group screening
For inactive groups, the KKT conditions (Equation 6) for SGS are
$(\nabla f(\beta)+\lambda\alpha\partial J_{\text{slope}}(\mathbf{0};v))_{\mathcal{G}_{\mathcal{Z}}}\in\lambda(1-\alpha)\partial J_{\text{gslope}}(\mathbf{0};w_{\mathcal{Z}}),$
[…]
The problem reduces to that of the gSLOPE screening rule (Section 3), with input given by $c=([\nabla f(\beta)+\lambda\alpha\partial J_{\text{slope}}(\mathbf{0};v)]_{\mathcal{G},-0.5})_{\downarrow}$ and $\phi=\lambda(1-\alpha)w$. To determine the form of the quantity $\partial J_{\text{slope}}(\mathbf{0};v)$, the term inside the $[\cdot]_{\mathcal{G}_{\mathcal{Z}},-0.5}$ operator needs to be as small as possible for Equation 7 to be satisfied. This term is found to be the soft thresholding operator, $S(\nabla f(\beta),\lambda\alpha):=\text{sign}(\nabla f(\beta))(|\nabla f(\beta)|-\lambda\alpha)_{+}$ (with the derivation presented in Appendix B.2).
By using the soft-thresholding operator, a valuable connection between SGS and SGL is made, as this operator is used in the gradient update step for SGL [29]. Such a connection has the potential to lead to new and more efficient optimization approaches for SGS that are more closely related to those used to solve SGL, similar to the recently developed coordinate descent algorithm for SLOPE [18].
Using this function, the (non-approximated) strong group screening rule for SGS is given in Proposition B.1. Using a similar Lipschitz assumption as for the gSLOPE rule gives the gradient approximation strong group screening rule for SGS (Proposition 4.1).
Proposition 4.1 (Gradient approximation strong group screening rule for SGS).
Let $\tilde{h}(\lambda):=([S(\nabla f(\hat{\beta}(\lambda)),\lambda\alpha)]_{\mathcal{G},-0.5})_{\downarrow}$. Taking $c=\tilde{h}(\lambda_{k})+\lambda_{k}(1-\alpha)w-\lambda_{k+1}(1-\alpha)w$ and $\phi=\lambda_{k+1}(1-\alpha)w$ as inputs for Algorithm A1, and assuming that for any $k\in\{1,\dots,l-1\}$,
$\left|\tilde{h}_{g}(\lambda_{k+1})-\tilde{h}_{g}(\lambda_{k})\right|\leq(1-\alpha)w_{g}|\lambda_{k+1}-\lambda_{k}|,\forall g=1,\dots,m,$
and $\mathcal{O}(\tilde{h}(\lambda_{k+1}))=\mathcal{O}(\tilde{h}(\lambda_{k}))$, then the algorithm returns a superset $\mathcal{S}_{g}(\lambda_{k+1})$ of $\mathcal{A}_{g}(\lambda_{k+1})$.
4.2 Variable screening
Whilst the group screening for SGS returns a superset, $\mathcal{S}_{g}$, of the active group set, exploiting the sparse-group penalization of SGS allows for the screened set to be further reduced by also applying variable screening. The KKT conditions (Equation 6) for a zero variable, $j\in\mathcal{G}_{g}$, in an active group, $g\in\mathcal{A}_{g}$, are
$-\nabla_{j}f(\beta)\in\lambda\alpha\partial J_{\text{slope}}(\mathbf{0};v_{j}).$ (8)
The gSLOPE subdifferential term vanishes as the numerator is zero in Theorem 3.1. The problem reduces to that of SLOPE variable screening, applied only to the variables that belong to groups in $\mathcal{A}_{g}$ and scaled by $\alpha$. The gradient approximated rule is formalized in Proposition 4.2 (the non-approximated version is presented in Proposition B.2).
Proposition 4.2 (Gradient approximation strong variable screening rule for SGS).
Let $\bar{h}(\lambda)=(\nabla f(\hat{\beta}(\lambda)))_{|\downarrow|}$. Taking $c=|\bar{h}(\lambda_{k+1})|+\lambda_{k}\alpha v-\lambda_{k+1}\alpha v$ and $\phi=\lambda_{k+1}\alpha v$ for only the variables in the groups in $\mathcal{A}_{g}(\lambda_{k+1})$ as inputs for Algorithm A1 , and assuming that for any $k\in\{1,\dots,l-1\}$,
$\left|\bar{h}_{j}(\lambda_{k+1})-\bar{h}_{j}(\lambda_{k})\right|\leq\alpha v_{j}|\lambda_{k+1}-\lambda_{k}|,\forall j\in\mathcal{G}_{\mathcal{A}_{g}(\lambda_{k+1})},$
and $\mathcal{O}(\bar{h}(\lambda_{k+1}))=\mathcal{O}(\bar{h}(\lambda_{k}))$, then the algorithm returns a superset $\mathcal{S}_{v}(\lambda_{k+1})$ of $\mathcal{A}_{v}(\lambda_{k+1})$.
In practice $\mathcal{A}_{g}(\lambda_{k+1})$ is not available, as this is exactly what we are trying to superset with any screening rule. However, Proposition 4.1 states that it is contained within $\mathcal{S}_{g}(\lambda_{k+1})$, so that the screened set can be used instead and complemented with KKT checks to ensure there are no violations, which would be performed anyway. For the checks, the group KKT conditions (Equation 6) are checked. For any active groups or groups violating the check (indicating that they should be active), the variable KKT conditions (Equation 8) are checked on the variables within those groups, to identify those that need to be added to $\mathcal{E}_{v}$.
5 Results
In this section, the effectiveness of the screening rules for gSLOPE and SGS is illustrated through the analysis of both synthetic (Section 5.1) and real data (Section 5.2). Any mentions of $\mathcal{E},\mathcal{A}$ in reference to group and variable metrics refers to $\mathcal{E}_{g},\mathcal{A}_{g}$ and $\mathcal{E}_{v},\mathcal{A}_{v}$ respectively. For SGS, $\mathcal{E}_{g}$ is the groups with representation in $\mathcal{E}_{v}$ (the groups fitted). Computational information is presented in Appendix F.1.
5.1 Synthetic data analysis
Set up.
For the synthetic data, a multivariate Gaussian design matrix, $\mathbf{X}\sim\mathcal{N}(\mathbf{0},\boldsymbol{\Sigma})\in\mathbb{R}^{400\times p}$, was used and the within-group correlation set to $\Sigma_{i,j}=\rho$, where $i$ and $j$ belong to the same group. The correlation and number of features were varied between $\rho\in\{0,0.3,0.6,0.9\}$ and $p\in\{500,1625,2750,3875,5000\}$, producing $20$ simulation cases. Each simulation case was repeated $100$ times. Two models were considered: linear and logistic. For the linear model, the output was generated as $y=\mathbf{X}\beta+\mathcal{N}(0,1)$ and for the logistic model the class probabilities were calculated using $\sigma(\mathbf{X}\beta+\mathcal{N}(0,1))$, where $\sigma$ is the sigmoid function. Groups of sizes between $3$ and $25$ were considered, of which $15\%$ were set to active. Within each group, $30\%$ of the variables were set to active with the signal sampled from $\beta\sim\mathcal{N}(0,5)$.
gSLOPE and SGS were each fit along a log-linear path of $l=50$ regularization parameters using warm starts. The first path value $\lambda_{1}$ was set to the exact value at which the first group enters the model (given in Appendix A.3 for gSLOPE and Appendix B.4 for SGS). The final regularization value was set to $\lambda_{50}=0.05\lambda_{1}$. The data was $\ell_{2}$ standardized and for the linear model an intercept was fitted. Both models had FDR-control parameters set to $0.05$, and $\alpha=0.95$ for SGS. Even though the models were fit using the adaptive three operator splitting algorithm (ATOS) [26], it should be noted that the screening rules are independent of the choice of fitting algorithm.
Primarily, the results for the linear model are presented, and the corresponding plots for the logistic model are presented in Appendix F.3.2.
Screening efficiency.
By comparing the sizes of the fitting set $(\mathcal{E})$ to the active set ($\mathcal{A})$, we can understand the difference between the minimum dimensionality needed against the dimensionality provided by the screening rules. As expected, these sizes increase as $\lambda$ decreases (Figure 2). In particular, the gap between sets widens as more features enter the model. The size of $\mathcal{E}$ remains far below the size of the full input space even at the lowest value of $\lambda$, showing that the screening rules have a benefit along the whole path. This finding is found to be independent of $\rho$, $p$, and model fitted (see Appendix F.3).
Figure 2: The number of groups/variables in $\mathcal{E},\mathcal{A}$ for both gSLOPE and SGS as a function of the regularization path for the linear model with $p=2750,\rho=0.6,m=197$. The results are averaged over $100$ repetitions, with the shaded regions corresponding to $95\%$ confidence intervals.
Figure 3: The proportion of groups/variables in $\mathcal{E},\mathcal{A}$, relative to the full input, shown for gSLOPE and SGS. This is shown as a function of the correlation ($\rho$), averaged over all cases of the input dimension ($p$), with $100$ repetitions for each $p$, for both linear and logistic models, with standard errors shown.
The performance of the screening rules are generally similar across linear and logistic models (Figure 3). As the within-group correlation increases, the size of both the screened and active sets decreases, causing the variables to move in the same direction within the same group. Due to the shape of the convex penalty of SLOPE, highly correlated features are clustered together [36]. This leads to smaller active sets, and in turn, also smaller screened sets. Additionally, it can be seen that the size difference between the screened and active sets decreases with correlation. Across the different correlation values, the linear model had smaller set sizes than the logistic one.
Runtime performance.
A key metric of performance for a screening rule is the time taken to fit a path of models. Increasing the value of $p$ demonstrates the computational cost reduction of applying a screening rule (Figure 4). For the smaller values of $p$, the runtime tends to be similar whether screening is applied or not. However, once the input dimension increases, the benefit of screening can be seen clearly. Additionally, increasing the within-group correlation tends to lead to smaller active and screened sets, but the benefit of screening appears to be robust under different correlations.
Figure 4: Runtime (in seconds) for fitting $50$ models along a path, shown for screening against no screening as a function of $p$, broken down into different correlation cases, for the linear model. The results are averaged over $100$ repetitions, with standard errors shown.
Figure 5: Runtime (in seconds) for fitting $100$ models along a path for screening against no screening for gSLOPE and SGS, averaged across the nine collections of pathways for each dataset, with standard errors shown.
A clear improvement with regards to runtime can be seen aggregating the results across all simulation cases (Table 1). In particular, for both models considered, screening halves the runtime for fitting $50$ models for SGS.
5.2 Real data experiments
Table 1: Runtime (in seconds) for fitting $50$ models along a path, shown for screening against no screening, for the linear and logistic models. The results are averaged across all cases of the correlation ($\rho$) and dimensionality ($p$), with standard errors shown.
Method Type Screen (s) No screen (s)
gSLOPE Linear $1016\pm 21$ $1623\pm 27$
gSLOPE Logistic $814\pm 8$ $1409\pm 11$
SGS Linear $735\pm 15$ $1830\pm 34$
SGS Logistic $407\pm 2$ $859\pm 6$
Table 2: The sizes of the sets $\mathcal{A}$ and $\mathcal{E}$ (with standard errors), averaged across $100$ path values and pathway collections for each of the two datasets. SGS var/grp shows the corresponding metrics for $\mathcal{A}_{v},\mathcal{E}_{v}$ and $\mathcal{A}_{g},\mathcal{E}_{g}$ for the bi-level screening procedure.
Method Dataset $\operatorname{card}(\mathcal{A})$ $\operatorname{card}(\mathcal{E})$
gSLOPE Cancer $66\pm 1$ $120\pm 2$
gSLOPE Colitis $32\pm 1$ $56\pm 2$
SGS grp Cancer $50\pm 1$ $115\pm 2$
SGS grp Colitis $18\pm 1$ $87\pm 3$
SGS var Cancer $71\pm 1$ $212\pm 4$
SGS var Colitis $25\pm 1$ $209\pm 6$
Datasets.
The screening rules were applied to two gene expression datasets with binary responses. The datasets were previously analyzed to measure the predictive performance of sparse-group regression models on large datasets [14, 29]. The cancer dataset contains gene data for $60$ breast cancer patients who had been treated with tamoxifen for 5 years, classified into binary labels depending on whether the cancer recurred [20]. The colitis dataset uses transcriptional profiles in peripheral blood mononuclear cells of $127$ patients to determine which patients have an inflammatory bowel disease (ulcerative colitis or Crohn’s). Of the $127$ samples, $42$ of them are controls [7].
The genes of both datasets were assigned to pathways (groups) that were obtained from the Molecular Signature Database. [^1] https://www.gsea-msigdb.org/gsea/msigdb/human/collections.jsp Nine collections of pathways were downloaded and as the pathways included different genes, nine unique design matrices for each dataset were created, each with a unique grouping structure. This gave $18$ design matrices in total to which the screening rules were applied.
Both gSLOPE and SGS were fitted with their FDR-control parameters set to $0.01$ and for SGS $\alpha=0.99$. Each model was fitted along a path of $100$ regularization parameters, with $\lambda_{1}$ set as described in Appendices A.3 and B.4, and $\lambda_{100}=0.01\lambda_{1}$. The data was $\ell_{2}$ standardized and no intercept was used.
Results.
For both gSLOPE and SGS, the screening rules lead to considerably faster runtimes (Figure 5). In particular, the use of the screening rules significantly benefits the gSLOPE approach. With the reduced input feature space, any issues with convergence of the fitting algorithm are substantially decreased. For the cancer dataset, $186$ model fits (out of the $900$) did not reach convergence with no screening, compared to only $16$ with screening. A similar trend occurred for the colitis dataset. As gSLOPE applies no variable penalization, it is forced to fit all variables within a group. For datasets with large groups, such as those considered here, this leads to a problematic fitting process which can include many noisy variables.
The analysis of the real data further illustrates the benefits of the bi-level screening to the runtime and performance of SGS. Figure 1(b) illustrates that for these large genetics datasets, the bi-level screening allows the input dimensionality for SGS to be reduced to a much greater extent than by just group screening.
Both screening rules can drastically reduce the input dimensionality as seen in Table 2. For the colitis dataset, the SGS screening rule reduces the input dimensionality to just $2\%$ of the full space. This comes without the cost of solution consistency. The estimated model coefficients with and without screening were very close to each other, with their $\ell_{2}$ distances of order $10^{-8}$ (Appendices F.2 and F.4).
5.3 KKT violations
The screening rules rely on assumptions that can fail. In these cases, KKT checks are implemented to ensure that no active variables are excluded from $\mathcal{E}_{v}$. A KKT violation occurs when a variable is not included in the screening set, $\mathcal{S}_{v}$, but should be according to the KKT optimality conditions (Equation 2), in which case it is added into the fitting set, $\mathcal{E}_{v}$. KKT violations are very rare for gSLOPE (Figure 6), occurring on the simulated data infrequently toward the start of the path. On the real data, no KKT violations occurred for gSLOPE.
(a) Group violations for gSLOPE and variable violations for SGS, fitted to linear models, averaged over all cases of $p$ and the correlation ($\rho$).
(b) Variable violations for SGS applied to the cancer and colitis datasets, averaged over the nine pathway collections.
Figure 6: The proportion of KKT violations relative to the full input space, as a function of the regularization path.
For SGS, KKT violations are more common, as the additional layer of screening carries with it additional assumptions. Additionally, the subdifferential term in Equation 7 was chosen to be as small as possible, leading to tighter screened sets. The increased violations are reflective of this. On both the simulated and real datasets, SGS recorded an increasing number of KKT violations as the fitted models became denser (Figure 6). The shape of the increasing number of KKT violations mirrors the log-linear shape of the regularization path. The trend of increasing KKT violations with denser models is also observed for the SLOPE strong screening rule [17].
Despite the increased number of KKT violations for SGS, which amounts to the additional computational cost of performing the KKT checks and repeated sub-fitting in Algorithm 1, the overall screening procedure still results in substantial runtime improvements, as evidenced in Tables 1 and 2. This shows that screening rules can provide improvements to computational cost, even with assumptions that are often violated.
6 Discussion
In this manuscript, we have developed strong screening rules for group-based SLOPE models: Group SLOPE and Sparse-group SLOPE, neither of which have any previous screening rules. The screening rule for gSLOPE screens out irrelevant groups before fitting. The screening rules for SGS perform bi-level screening, based on our proposed sparse-group screening framework.
The two proposed screening rules differ from the existing SLOPE screening rules both in construction and in final outcome. gSLOPE only performs screening through the groups, forcing it to keep all variables within the screened groups, even noisy ones, in the optimization process. SGS, similarly to SLOPE and in contrast to gSLOPE, performs variable screening after its group screening step. As additional assumptions are required for SGS to perform bi-level selection, its computational efficiency is not as effective as the one observed with the SLOPE screening rules [17]. The superiority of SGS in variable selection and predictive performance when group information is available overcomes the additional computational burden [14].
Through comprehensive analysis of synthetic and real data, we illustrate that the screening rules lead to dramatic improvements in the runtime of gSLOPE and SGS models, as well as for group-based OSCAR models (Appendix E). This is achieved without affecting model accuracy. This is particularly important in datasets where $p\gg n$, such as genetics ones, which is the main motivation behind SLOPE [4]. The screening rules presented in this manuscript allow group-based SLOPE, and by extension, group-based OWL models, to achieve computational fitting times more in line with lasso-based models, making them more widely accessible.
Limitations.
One of the limitations of our proposed screening rules for SGS is the number of assumptions made. A future direction includes the exploration of alternative strong screening rules that require fewer assumptions. Alternatively, the development of safe rules, which guarantee that only inactive variables are discarded, could be an avenue for future exploration, potentially incorporating safe and strong rules together in a hybrid rule [32, 37]. A comparison between safe and the proposed strong rules developed in this manuscript could provide further insight into which type of screening is most beneficial for non-separable penalties.
Appendix
[…]
Appendix A Group SLOPE
A.1 Penalty weights
The penalty weights for gSLOPE were derived to provide group FDR-control under orthogonal designs [6]. For the FDR-control parameter $q_{g}\in(0,1)$, they are given by (where the indexing corresponds to the sorted groups)
$w_{i}^{\text{max}}=\max_{j=1,\dots,m}\left\{\frac{1}{\sqrt{p_{j}}}F^{-1}_{\chi_{p_{j}}}(1-q_{g}i/m)\right\},\;\text{for}\;i=1,\dots,m,$
where $F_{\chi_{p_{j}}}$ is the cumulative distribution function of a $\chi$ distribution with $p_{j}$ degrees of freedom. A relaxation to this sequence is applied in [6], to give
$w_{i}^{\text{mean}}=\overline{F}^{-1}_{\chi_{p_{j}}}(1-q_{g}i/m),\;\text{where}\;\overline{F}_{\chi_{p_{j}}}(x):=\frac{1}{m}\sum_{j=1}^{m}F_{\chi_{p_{j}}}(\sqrt{p_{j}}x).$ (9)
The mean sequence weights defined in Equation 9 are used for all gSLOPE numerical simulations in this manuscript (shown in Figure A1).
Figure A1: The gSLOPE weights, $w$, shown for Figure 4 for $p=500,m=100,q_{g}=0.05$.
A.2 Theory
Proof of Theorem 3.1.
The proof is similar to that of Theorem 2.7 in [6], where the subdifferential of gSLOPE is derived under equal groups. It is derived here under more general terms. The subdifferential needs to be derived under two cases:
  1. 1.
    Inactive groups, $\mathcal{G}_{\mathcal{Z}}$.
  2. 2.
    Active groups, $\mathcal{G}_{\mathcal{A}}$.
Case 1: For inactive groups, we consider the subdifferential at zero. The subdifferential of a norm at zero is given by the dual norm of the unit ball [27],
$\partial J_{\text{gslope}}(\mathbf{0};w)=\mathbf{B}_{J_{\text{gslope}}^{*}(\mathbf{0};w)}[0,1]=\{x:J_{\text{gslope}}^{*}(x;w)\leq 1\}.$
The dual norm for gSLOPE is given by [6]
$J_{\text{gslope}}^{*}(x;w)=J_{\text{slope}}^{*}([x]_{\mathcal{G},-0.5}).$
Hence, the dual norm unit ball is
$\mathbf{B}_{J_{\text{gslope}}^{*}(\mathbf{0};w)}[0,1]=\{x:[x]_{\mathcal{G},-0.5}\in\mathbf{B}_{J_{\text{slope}}^{*}(\mathbf{0};w)}[0,1]\},$
where $\mathbf{B}_{J_{\text{slope}}^{*}(\mathbf{0};w)}[0,1]=\{x\in\mathbb{R}^{m}:\text{cumsum}(|x|_{\downarrow}-w)\preceq\mathbf{0}\}$ is the unit ball of the dual norm to $J_{\text{slope}}$ [4]. Using this, the subdifferential at zero for the inactive groups, $\mathcal{Z}$, is given by
$\partial J_{\text{gslope}}(\mathbf{0};w_{\mathcal{Z}})=\{x\in\mathbb{R}^{\operatorname{card}(\mathcal{G}_{\mathcal{Z}})}:[x]_{\mathcal{G}_{\mathcal{Z}},-0.5}\in\partial J_{\text{slope}}(\mathbf{0};w_{\mathcal{Z}})\}.$
Case 2: Without loss of generality, denote the group index $s$ such that $\|\beta^{(g)}\|_{2}=0$ for $g>s$ (inactive groups) and $\|\beta^{(g)}\|_{2}\neq 0$ for $g\leq s$ (active groups). In other words, $g\in\mathcal{G}_{\mathcal{A}}$ if $g\leq s$. Define a set $D=\{d\in\mathbb{R}^{p}:\|\beta^{(1)}+d^{(1)}\|_{2}>\ldots>\|\beta^{(s)}+d^{(s)}\|_{2},\|\beta^{(s)}+d^{(s)}\|_{2}>\|d^{(g)}\|_{2},g>s\}$. By definition of a subdifferential, if $x\in\partial J_{\text{gslope}}(\beta;w)$, then for all $d\in D$
$\sum_{g=1}^{m}\sqrt{p_{g}}w_{g}\|\beta^{(g)}+d^{(g)}\|_{2}\geq\sum_{g=1}^{m}\sqrt{p_{g}}w_{g}\|\beta^{(g)}\|_{2}+x^{\top}d.$
Splitting this up into whether the groups are active (whether $g\leq s$):
$\sum_{g=1}^{s}\sqrt{p_{g}}w_{g}\|\beta^{(g)}+d^{(g)}\|_{2}+\sum_{g=s+1}^{m}\sqrt{p_{g}}w_{g}\|d^{(g)}\|_{2}\geq$ $\sum_{g=1}^{s}\sqrt{p_{g}}w_{g}\|\beta^{(g)}\|_{2}$ (10)
$+\sum_{g=1}^{s}x^{(g)T}d^{(g)}+\sum_{g=s+1}^{m}x^{(g)T}d^{(g)}.$
Now, for $g\in\mathcal{G}_{\mathcal{A}}$, define a new set $D_{g}=\{d\in D:d^{(j)}\equiv\mathbf{0},j\neq g\}$. Taking $d\in D_{g}$, Equation 10 becomes
$\sqrt{p_{g}}w_{g}\|\beta^{(g)}+d^{(g)}\|_{2}\geq\sqrt{p_{g}}w_{g}\|\beta^{(g)}\|_{2}+x^{(g)T}d^{(g)}.$
Since the set $\{d^{(g)}:d\in D_{g}\}$ is open in $\mathbb{R}^{p_{g}}$ and contains zero, by Corollary G.1 in [6], it follows that $x^{(g)}\in\partial f_{g}(b^{(g)})$ for $f_{g}:\mathbb{R}^{p_{g}}\rightarrow\mathbb{R},f_{g}(x)=w_{g}\sqrt{p_{g}}\|x\|_{2}$. Now, for $g\leq s$, $f_{g}$ is differentiable in $\beta^{(g)}$, giving
$x^{(g)}=w_{g}\sqrt{p_{g}}\frac{\beta^{(g)}}{\|\beta^{(g)}\|_{2}},$
proving the result. ∎
Proof of Proposition 3.2.
Suppose we have $\mathcal{B}\neq\emptyset$ after running the algorithm. Then, plugging in $h(\lambda_{k+1})=([\nabla f(\hat{\beta}(\lambda_{k+1}))]_{\mathcal{G},-0.5})_{\downarrow}$ gives
$\text{cumsum}\Bigl{(}\bigl{(}([\nabla f(\hat{\beta}(\lambda_{k+1}))]_{\mathcal{G},-0.5})_{\downarrow}\bigr{)}_{\mathcal{B}}-\lambda_{k+1}w_{\mathcal{B}}\Bigr{)}\prec\mathbf{0},$
so that by the gSLOPE subdifferential (Theorem 3.1) all groups in $\mathcal{B}$ are inactive. This is valid by the KKT conditions (Equation 2), as we know that $-\nabla f(\hat{\beta}(\lambda_{k+1}))\in\partial J_{\text{gslope}}(\mathbf{0};w)$. Hence, $\mathcal{S}_{g}(\lambda_{k+1})$ will contain the active set $\mathcal{A}_{g}(\lambda_{k+1})$. ∎
Proof of Proposition 3.3.
Since $\text{cumsum}(y)\succeq\text{cumsum}(x)\iff y\succeq x$ [17], we only need to show for a group $g$,
$|h_{g}(\lambda_{k+1})|\leq|h_{g}(\lambda_{k})|+\lambda_{k}w_{g}-\lambda_{k+1}w_{g}.$
Applying the reverse triangle inequality to the Lipschitz assumption gives
$|h_{g}(\lambda_{k+1})|-|h_{g}(\lambda_{k})|\leq\left|h_{g}(\lambda_{k+1})-h_{g}(\lambda_{k})\right|\leq\lambda_{k}w_{g}-\lambda_{k+1}w_{g}$
$\implies$ $|h_{g}(\lambda_{k+1})|\leq|h_{g}(\lambda_{k})|+\lambda_{k}w_{g}-\lambda_{k+1}w_{g},$
proving the result. ∎
A.3 Path start derivation
The aim is to find the value of $\lambda$ at which the first group enters the model. When all features are zero, the gSLOPE KKT conditions (Equation 2) are
$\mathbf{0}\in\nabla f(\mathbf{0})+\lambda\partial J_{\text{gslope}}(\mathbf{0};w).$
This is satisfied when
$[\nabla f(\mathbf{0})]_{\mathcal{G},-0.5}\in\partial J_{\text{slope}}(\mathbf{0};\lambda w)\implies\text{cumsum}\bigl{(}([\nabla f(\mathbf{0})]_{\mathcal{G},-0.5})_{\downarrow}-\lambda w\bigr{)}\preceq\mathbf{0}.$
Rearranging this gives
$\lambda\succeq\text{cumsum}\bigl{(}([\nabla f(\mathbf{0})]_{\mathcal{G},-0.5})_{\downarrow}\bigr{)}\oslash\text{cumsum}(w),$
[…]
$\lambda_{1}=\max\left\{\text{cumsum}\bigl{(}([\nabla f(\mathbf{0})]_{\mathcal{G},-0.5})_{\downarrow}\bigr{)}\oslash\text{cumsum}(w)\right\}.$
This can be verified by noting that $\lambda_{1}=J^{*}_{\text{gslope}}(\nabla f(\mathbf{0});w)$ [22]. Now, $J^{*}_{\text{gslope}}(x;w)=J^{*}_{\text{slope}}([x]_{\mathcal{G},-0.5};w)$ [6]. The dual norm of SLOPE is given by [23]
$J^{*}_{\text{slope}}(x;w)=\max\left\{\text{cumsum}(|x|_{\downarrow})\oslash\text{cumsum}(w)\right\}.$
Therefore, $\lambda_{1}$ is as before.
Appendix B Sparse-group SLOPE
B.1 Penalty weights
The penalty weights for SGS provide variable and group FDR-control simultaneously, under orthogonal designs [14]. They are given by (where the indexing corresponds to the sorted variables/groups)
$v_{i}^{\text{max}}=\max_{j=1,\dots,m}\left\{\frac{1}{\alpha}F_{\mathcal{N}}^{-1}\left(1-\frac{q_{v}i}{2p}\right)-\frac{1}{3\alpha}(1-\alpha)a_{j}w_{j}\right\},\;i=1,\dots,p,$ (11)
$w_{i}^{\text{max}}=\max_{j=1,\dots,m}\left\{\frac{F_{\text{FN}}^{-1}(1-\frac{q_{g}i}{m})-\alpha\sum_{k\in\mathcal{G}_{j}}v_{k}}{(1-\alpha)p_{j}}\right\},\;i=1,\dots,m,$ (12)
where $F_{\chi_{p_{j}}}$ is the cumulative distribution function of a $\chi$ distribution with $p_{j}$ degrees of freedom, $F_{\mathcal{N}}$ is the cumulative distribution function of a folded Gaussian distribution, and $a_{j}$ is a quantity that requires estimation. The estimator $\hat{a}_{j}=\lfloor\alpha p_{j}\rfloor$ is proposed in [14]. As with gSLOPE (Appendix A.1), a relaxation is possible, giving the weights
$v_{i}^{\text{mean}}=\overline{F}_{\mathcal{N}}^{-1}\left(1-\frac{q_{v}i}{2p}\right),\;\text{where}\;\overline{F}_{\mathcal{N}}(x):=\frac{1}{m}\sum_{j=1}^{m}F_{\mathcal{N}}\left(\alpha x+\frac{1}{3}(1-\alpha)a_{j}w_{j}\right),$ (13)
$w_{i}^{\text{mean}}=\overline{F}_{\text{FN}}^{-1}\left(1-\frac{q_{g}i}{p}\right),\;\text{where}\;\overline{F}_{\text{FN}}(x):=\frac{1}{m}\sum_{j=1}^{m}F_{\text{FN}}\left((1-\alpha)p_{j}x+\alpha\sum_{k\in\mathcal{G}_{j}}v_{k}\right).$ (14)
The mean weights defined in Equations 13 and 14 are used for all SGS numerical simulations in this manuscript (shown in Figure A2).
Figure A2: The SGS weights, $(v,w)$, shown for Figure 4 for $p=500,m=100,q_{v}=0.05,q_{g}=0.05,\alpha=0.95$.
B.2 Derivation of soft thresholding operator
To determine the form of the quantity $\partial J_{\text{slope}}(\mathbf{0};v)$, consider that for Equation 7 to be satisfied, the term inside the []delimited-[][$\cdot$][ ⋅ ] operator needs to be as small as possible. Now,
$\partial J_{\text{slope}}(\mathbf{0};v)=\{y:\text{cumsum}(|y|)\preceq\text{cumsum}(v)\}.$
Note that $\text{cumsum}(y)\preceq\text{cumsum}(x)\iff y\preceq x$. We consider the cases:
  1. 1.
    $\nabla_{i}f(\beta)>\lambda\alpha v_{i}$: choose $y_{i}=-v_{i}$.
  2. 2.
    $\nabla_{i}f(\beta)<-\lambda\alpha v_{i}$: choose $y_{i}=v_{i}$.
  3. 3.
    $\nabla_{i}f(\beta)\in[-\lambda\alpha v_{i},\lambda\alpha v_{i}]$: choose $y_{i}=\nabla_{i}f(\beta)/\alpha\lambda$.
Hence, the term becomes
$S(\nabla f(\beta),\lambda\alpha):=\text{sign}(\nabla f(\beta))(|\nabla f(\beta)|-\lambda\alpha)_{+},$
which is the soft thresholding operator.
B.3 Theory
Proposition B.1 (Strong group screening rule for SGS).
Let $\tilde{h}(\lambda):=([S(\nabla f(\beta),\lambda\alpha)]_{\mathcal{G},-0.5})_{\downarrow}$. Then taking $c=\tilde{h}(\lambda_{k+1})$ and $\phi=(1-\alpha)\lambda_{k+1}w$ as inputs for Algorithm A1 returns a superset $\mathcal{S}_{g}(\lambda_{k+1})$ of the active set $\mathcal{A}_{g}(\lambda_{k+1})$.
Proof of Proposition B.1.
The proof is similar to that of Proposition 3.2. Suppose we have $\mathcal{B}\neq\emptyset$ after running the algorithm. Then,
$\text{cumsum}(\tilde{h}_{\mathcal{B}}(\lambda_{k+1})-\lambda_{k+1}(1-\alpha)w_{\mathcal{B}})$ $\prec\mathbf{0}$
$\implies\text{cumsum}\Bigl{(}\bigl{(}[S(\nabla f(\beta),\lambda_{k+1}\alpha)]_{\mathcal{G},-0.5})_{\downarrow}\bigr{)}_{\mathcal{B}}-\lambda_{k+1}(1-\alpha)w_{\mathcal{B}}\Bigr{)}$ $\prec,\mathbf{0}$
so that by the SGS subdifferential (Equation 7), all groups in $\mathcal{B}$ are inactive. Hence, $\mathcal{S}_{g}(\lambda_{k+1})$ will contain the active set $\mathcal{A}_{g}(\lambda_{k+1})$. ∎
Proof of Proposition 4.1.
The proof is identical to that of Proposition 3.3, replacing $h_{g}(\cdot)$ with $\tilde{h}_{g}(\cdot)$ and $\lambda_{k+1}w$ by $\lambda_{k+1}(1-\alpha)w$. ∎
Proposition B.2 (Strong variable screening rule for SGS).
Let $\bar{h}(\lambda)=|(\nabla f(\hat{\beta}(\lambda)))|_{\downarrow}$. Then taking $c=\bar{h}(\lambda_{k+1})$ and $\phi=\lambda_{k+1}\alpha v$ for only the variables contained in the groups in $\mathcal{A}_{g}(\lambda_{k+1})$ in Algorithm A1 returns a superset $\mathcal{S}_{v}(\lambda_{k+1})$ of the active set $\mathcal{A}_{v}(\lambda_{k+1})$.
Proof.
Suppose we have $\mathcal{B}\neq\emptyset$ after running the algorithm. Then, we have
$\text{cumsum}(\bar{h}_{\mathcal{B}}(\lambda_{k+1})-\lambda_{k+1}\alpha v_{\mathcal{B}})\prec\mathbf{0}\implies\text{cumsum}\Bigl{(}\bigl{(}|\nabla f(\hat{\beta}(\lambda_{k+1}))|_{\downarrow}\bigr{)}_{\mathcal{B}}-\lambda_{k+1}\alpha v_{\mathcal{B}}\Bigr{)}\prec\mathbf{0},$ (15)
so that by the SGS subdifferential for non-zero groups (Equation 8), all variables in $\mathcal{B}$ are inactive. Hence, $\mathcal{S}_{v}(\lambda_{k+1})$ will contain the active set $\mathcal{A}_{v}(\lambda_{k+1})$. ∎
Proof for Proposition 4.2.
The proof is identical to that of Proposition 3.3, replacing $h_{g}(\cdot)$ with $\bar{h}_{g}(\cdot)$, $\lambda_{k+1}v$ with $\lambda_{k+1}\alpha v$, and considering only variables in the groups contained in $\mathcal{A}_{g}(\lambda_{k+1})$. ∎
B.4 Path start derivation
The aim is to find the value of $\lambda$ at which the first variable enters the model. When all features are zero, the SGS KKT conditions (Equation 6) are
$-\nabla f(\mathbf{0})\in\lambda(1-\alpha)\partial J_{\text{gslope}}(\mathbf{0};w)+\lambda\alpha\partial J_{\text{slope}}(\mathbf{0};v)$
$\implies$ $-\frac{1}{\lambda}\nabla f(\mathbf{0})-(1-\alpha)\partial J_{\text{gslope}}(\mathbf{0};w)\in\alpha\partial J_{\text{slope}}(\mathbf{0};v)$
$\implies$ $\text{cumsum}\left(\left|-\frac{1}{\lambda}\nabla f(\mathbf{0})-(1-\alpha)\partial J_{\text{gslope}}(\mathbf{0};w)\right|_{\downarrow}-\alpha v\right)\preceq\mathbf{0}.$
By the reverse triangle inequality and ordering of the group weights
$\frac{1}{\lambda}\text{cumsum}\left(|\nabla f(\mathbf{0})|_{\downarrow}\right)\preceq\text{cumsum}((1-\alpha)|\partial J_{\text{gslope}}(\mathbf{0};w)|-\alpha v)$
$\implies$ $\lambda\succeq\text{cumsum}(|\nabla f(\mathbf{0})|_{\downarrow})\oslash\text{cumsum}((1-\alpha)|\partial J_{\text{gslope}}(\mathbf{0};w)|-\alpha v).$
Now, note that for $x\in J_{\text{gslope}}(\mathbf{0};w)$, it holds
$\text{cumsum}([x]_{\mathcal{G},-0.5}-w)\preceq\mathbf{0}\implies[x]_{\mathcal{G},-0.5}\preceq w\implies\|x^{(g)}\|_{2}\leq\sqrt{p_{g}}w_{g},\forall g\in\mathcal{G}.$
This is satisfied at the upper limit at $x=\tau\omega$, where $\tau$ and $\omega$ are expanded vectors of the group sizes ($\sqrt{p_{g}})$ and penalty weights ($w_{g}$) to $p$ dimensions, so that each variable within the same group is assigned the same value. Hence,
$\lambda_{1}=\max\left\{\text{cumsum}(|\nabla f(\mathbf{0})|_{\downarrow})\oslash\text{cumsum}((1-\alpha)\tau\omega-\alpha v)\right\}.$
Appendix C SLOPE Algorithm
Algorithm A1 Cumsum algorithm from [17]
  Input: $c\in\mathbb{R}^{p},\phi\in\mathbb{R}^{p}$, where ϕ1subscriptitalic-$\phi_{1}\geq\cdots\geq$ $\phi_{p}\geq 0$
  $\mathcal{S},\mathcal{B}\leftarrow\varnothing$
  for $i=1$ to $p$ do
     $\mathcal{B}\leftarrow\mathcal{B}\cup\{i\}$
     if $\text{cumsum}(c_{\mathcal{B}}-\phi_{\mathcal{B}})\geq 0$ then
        $\mathcal{S}\leftarrow\mathcal{S}\cup\mathcal{B}$
        $\mathcal{B}\leftarrow\varnothing$
     end if
  end for
  Output: $\mathcal{S}$
Appendix D Screening rule framework
D.1 Group SLOPE algorithm
For the following is performed for $k=1,\ldots,l-1$:
  1. 1.
    Set $\mathcal{E}_{g}=\mathcal{S}_{g}(\lambda_{k+1})\cup\mathcal{A}_{g}(\lambda_{k})$, where $\mathcal{S}_{g}(\lambda_{k+1})$ is obtained using Proposition 3.3.
  2. 2.
    Compute $\hat{\beta}(\lambda_{k+1})$ by Equation 1 with the gSLOPE norm using only the groups in $\mathcal{E}_{g}.$ For any groups not in $\mathcal{E}_{g}$, $\hat{\beta}(\lambda_{k+1})$ is set to zero.
  3. 3.
    Check the KKT conditions (Equation 2) for all groups at this solution.
  4. 4.
    If there are no violations, we are done and keep $\hat{\beta}(\lambda_{k+1})$. Otherwise, add the violating groups into $\mathcal{E}$ and return to Step 2.
D.2 SGS algorithm
For the following is performed for $k=1,\ldots,l-1$:
  1. 1.
    Group screen step: Calculate $\mathcal{S}_{g}(\lambda_{k+1})$ using Proposition 4.1.
  2. 2.
    Variable screen step: Set $\mathcal{E}_{v}=\mathcal{S}_{v}(\lambda_{k+1})\cup\mathcal{A}_{v}(\lambda_{k})$, where $\mathcal{S}_{v}(\lambda_{k+1})$ is obtained using Proposition 4.2 with only the groups in $\mathcal{S}_{g}(\lambda_{k+1})$.
  3. 3.
    Compute $\hat{\beta}(\lambda_{k+1})$ by Equation 1 with the SGS norm using only the features in $\mathcal{E}_{v}$. For features not in $\mathcal{E}_{v}$, $\hat{\beta}(\lambda_{k+1})$ is set to zero.
  4. 4.
    Check the KKT conditions (Equation 6) for all features at this solution.
  5. 5.
    If there are no violations, we are done and keep $\hat{\beta}(\lambda_{k+1})$, otherwise add in the violating variables into $\mathcal{E}_{g}$ and return to Step 3.
Appendix E Group-based OSCAR
This section presents an extension of the proposed screening rules for group-based OSCAR models. Firstly the OSCAR models are presented, and subsequently, the performance of the screening rules on synthetic data is presented.
E.1 Model definition
The Ordered Weighted $\ell_{1}$ (OWL) framework is defined as [35]
$\hat{\beta}=\operatorname*{arg\,min}_{\beta}\left\{\nabla f(\beta)+\lambda J_{\text{owl}}(\beta;v)\right\},$
where $J_{\text{owl}}(\beta;v)=\sum_{i=1}^{p}v_{i}|\beta|_{(i)}$, $|\beta|_{(1)}\geq\ldots\geq|\beta|_{(p)}$, and $v$ are non-negative non-increasing weights. SLOPE is a special case of OWL where the weights are taken to be the Benjamini-Hochberg critical values [4]. Octagonal Shrinkage and Clustering Algorithm for Regression (OSCAR) [5] is a further special case of OWL (often referred to as OWL with linear decay) where for a variable $i$, the weights are taken to be $v_{i}=\sigma_{1}+\sigma_{2}(p-i)$, and $\sigma_{1},\sigma_{2}$ are to be set. In Bao et al. [3] they are set to $\sigma_{1}=d_{i}\|\mathbf{X}^{\top}y\|_{\infty},\sigma_{2}=\sigma_{1}/p$, where $d_{i}=i\times e^{-2}$.
Group OSCAR (gOSCAR) and Sparse-group OSCAR (SGO) are defined using the frameworks provided by gSLOPE [6] and SGS [14], respectively, but instead using the weights (for a variable $i$ and a group $g$)
$v_{i}=\sigma_{1}+\sigma_{2}(p-i),\;w_{g}=\sigma_{1}+\sigma_{3}(m-g),\;\sigma_{3}=\sigma_{1}/m.$ (16)
with the weights visualised in Figure A3.
Figure A3: The SGO weights, $(v,w)$, shown for Figure 4 for $p=500,m=100,q_{v}=0.05,q_{g}=0.05,\alpha=0.95$.
E.2 Results
Similar observations and conclusions made for the screening rules of gSLOPE and SGS are made for gOSCAR and SGO (Figures A4 - A8).
Figure A4 illustrates the effectiveness of the bi-level screening of SGO, similar to the effectiveness observed for SGS. Figures A5 - A6 showcase the efficiency of the screening rules on the proportion of the selected groups/variables. The screening rules are found to be effective across different data characteristics, with the running time of the models significantly decreasing (Figure A7). Similarly to SGS, KKT violations are more common compared to the ones observed for gOSCAR (Figure A8). This is due to the additional assumptions made at the additional screening layer of SGO and SGS. Similarly to Figure 6(a), the shape of the increasing number of KKT violations mirrors the log-linear shape of the regularization path.
Figure A4: The proportion of variables in $\mathcal{S}_{v}$ relative to the full input for SGO, shown for group and bi-level screening plotted as a function of the regularization path, applied to the synthetic data (Section 5.1). The data are generated under a linear model for $p=500,5000$. The results are averaged over 100 repetitions and 95% confidence intervals are shown (the SGO equivalent of Figure 1 (a)).
Figure A5: The number of groups/variables in $\mathcal{E},\mathcal{A}$ for both gOSCAR and SGO as a function of the regularization path for the linear model with $p=2750,\rho=0.6,m=197$. The results are averaged over $100$ repetitions, with the shaded regions corresponding to $95\%$ confidence intervals (the gOSCAR/SGO equivalent of Figure 2).
Figure A6: The proportion of groups/variables in $\mathcal{E},\mathcal{A}$, relative to the full input, shown for gOSCAR and SGO. This is shown as a function of the correlation ($\rho$), averaged over all cases of the input dimension ($p$), with $100$ repetitions for each $p$, for both linear and logistic models, with standard errors shown (the gOSCAR/SGO equivalent of Figure 3).
Figure A7: Runtime (in seconds) for fitting $50$ models along a path, shown for screening against no screening as a function of $p$, broken down into different correlation cases, for the linear model. The results are averaged over $100$ repetitions, with standard errors shown (the OSCAR equivalent of Figure 4).
Figure A8: The proportion of KKT violations relative to the full input space, as a function of the regularization path. Group violations for gOSCAR and variable violations for SGS, fitted to linear models, averaged over all cases of $p$ and $\rho$ (the gOSCAR/SGO equivalent of Figure 6(a)).
Appendix F Results
F.1 Computational information
The simulated experiments were executed on a high-performance computing cluster (x86-64 Linux GNU) and the real data analysis was conducted on a Apple Macbook Air (M1, 8GB).
F.2 Model Comparison
This section presents the accuracy of the models with and without screening, by comparing the $\ell_{2}$ distances observed between the screened and non-screened fitted values.
Synthetic data.
For the linear model, the maximum $\ell_{2}$ distances observed between the screened and non-screened fitted values were of order $10^{-5}$ for gSLOPE and $10^{-8}$ for SGS (Table F.3). Across the different cases, 98000 models were fit in total for each approach (excluding the models for $\lambda_{1}$, where no screening is applied). Of these model fits, there were no instances for gSLOPE where $\mathcal{E}$ was not a superset of $\mathcal{A}$. There was only one instance (out of the 98000) that this occurred for SGS, where $\mathcal{E}$ was missing a single variable contained in $\mathcal{A}$ (which had a non-screen fitted value of $\hat{\beta}=-0.004$).
For the logistic model, the maximum $\ell_{2}$ distances observed between the screened and non-screened fitted values were of order $10^{-8}$ for both gSLOPE and SGS (Table F.3.2). Across the different cases, 98000 models were fit in total for each approach (excluding the models for $\lambda_{1}$, where no screening is applied). Of these model fits, there were no instances for gSLOPE or SGS where $\mathcal{E}$ was not a superset of $\mathcal{A}$.
Real data.
In the real data analysis, the estimated coefficients with and without screening were very close to each other (with $\ell_{2}$ to be of order $10^{-8}$) for both SGS and gSLOPE (Table F.4). No instances occurred where $\mathcal{E}$ was not a superset of $\mathcal{A}$.
F.3 Additional results from the simulation study
Table A1: Variable screening metrics for SGS using linear and logistic models for the simulation study presented in Section 5.1. The number of variables in $\mathcal{A}_{v},\mathcal{S}_{v},\mathcal{E}_{v}$, and $\mathcal{K}_{v}$ are shown, averaged across all $20$ cases of the correlation ($\rho$) and $p$. Standard errors are shown.
Method Type $\operatorname{card}(\mathcal{A}_{v})$ $\operatorname{card}(\mathcal{S}_{v})$ $\operatorname{card}(\mathcal{E}_{v})$ $\operatorname{card}(\mathcal{K}_{v})$
SGS Linear $179\pm 3$ $313\pm 5$ $363\pm 6$ $51\pm 1$
SGS Logistic $230\pm 3$ $405\pm 5$ $472\pm 6$ $66\pm 1$
Table A2: General and group screening metrics for SGS and gSLOPE using linear and logistic models for the simulation study presented in Section 5.1. General metrics: the runtime (in seconds) for screening against no screening, the number of fitting iterations for screening against no screening, and the $\ell_{2}$ distance between the fitted values obtained with screening and no screening. Group screening metrics: the number of groups in $\mathcal{A}_{g},\mathcal{S}_{g},\mathcal{E}_{g}$, and $\mathcal{K}_{g}$. The results are averaged across all $20$ cases of the correlation ($\rho$) and $p$. Standard errors are shown.
Method Type Runtime (s)
Runtime (s) screen
$\operatorname{card}(\mathcal{A}_{g})$ $\operatorname{card}(\mathcal{S}_{g})$ $\operatorname{card}(\mathcal{E}_{g})$ $\operatorname{card}(\mathcal{K}_{g})$ Num it
Num it screen
$\ell_{2}$ dist
to no screen
gSLOPE Linear $1016\pm 21$ $1623\pm 27$ $55\pm 1$ $76\pm 1$ $76\pm 1$ $0.006\pm 0.004$ $333\pm 6$ $351\pm 6$ $2\times 10^{-6}\pm 1\times 10^{-6}$
gSLOPE Logistic $814\pm 8$ $1409\pm 11$ $71\pm 1$ $97\pm 1$ $97\pm 1$ $0.014\pm 0.014$ $78\pm 1$ $83\pm 1$ $1\times 10^{-8}\pm 1\times 10^{-8}$
SGS Linear $735\pm 15$ $1830\pm 34$ $61\pm 1$ $84\pm 1$ $91\pm 1$ $21\pm 1$ $91\pm 3$ $708\pm 12$ $2\times 10^{-9}\pm 3\times 10^{-9}$
SGS Logistic $407\pm 2$ $859\pm 6$ $84\pm 1$ $107\pm 1$ $118\pm 1$ $22\pm 0.3$ $7\pm 0.2$ $51\pm 0.8$ $4\times 10^{-9}\pm 3\times 10^{-10}$
F.3.1 Additional results for the linear model
(a) $p=500,\rho=0$
(b) $p=500,\rho=0.9$
(c) $p=5000,\rho=0$
(d) $p=5000,\rho=0.9$
Figure A9: The number of groups/variables in $\mathcal{E},\mathcal{A}$ as a function of the regularization path for the linear model with SGS and gSLOPE, shown for different values of the correlation ($\rho$) and $p$. The results are averaged over $100$ repetitions, with $95\%$ confidence intervals shown.
Table A3: Variable screening metrics for SGS using a linear model for the simulation study presented in Section 5.1. The number of variables in $\mathcal{A}_{v},\mathcal{S}_{v},\mathcal{E}_{v}$, and $\mathcal{K}_{v}$ are shown. The results are shown for different values of $p$, averaged across $\rho\in\{0,0.3,0.6,0.9\}$. Standard errors are shown.
Method $p$ $\operatorname{card}(\mathcal{A}_{v})$ $\operatorname{card}(\mathcal{S}_{v})$ $\operatorname{card}(\mathcal{E}_{v})$ $\operatorname{card}(\mathcal{K}_{v})$
SGS $500$ $19\pm 1$ $24\pm 1$ $28\pm 1$ $4\pm 0.2$
SGS $1625$ $83\pm 3$ $138\pm 5$ $161\pm 6$ $22\pm 1$
SGS $2750$ $188\pm 7$ $316\pm 10$ $370\pm 12$ $54\pm 2$
SGS $3875$ $268\pm 9$ $470\pm 14$ $548\pm 16$ $78\pm 3$
SGS $5000$ $334\pm 10$ $618\pm 17$ $712\pm 20$ $95\pm 3$
Table A4: General and group screening metrics for SGS and gSLOPE using linear models for the simulation study presented in Section 5.1. General metrics: the runtime (in seconds) for screening against no screening, the number of fitting iterations for screening against no screening, and the $\ell_{2}$ distance between the fitted values obtained with screening and no screening. Group screening metrics: the number of groups in $\mathcal{A}_{g},\mathcal{S}_{g},\mathcal{E}_{g}$, and $\mathcal{K}_{g}$. The results are shown for different values of $p$, averaged across $\rho\in\{0,0.3,0.6,0.9\}$. Standard errors are shown.
Method $p$ Runtime (s)
Runtime (s) screen
$\operatorname{card}(\mathcal{A}_{g})$ $\operatorname{card}(\mathcal{S}_{g})$ $\operatorname{card}(\mathcal{E}_{g})$ $\operatorname{card}(\mathcal{K}_{g})$ Num it
Num it screen
$\ell_{2}$ dist
to no screen
gSLOPE $500$ $89\pm 1$ $144\pm 1$ $9\pm 0.2$ $10\pm 0.3$ $10\pm 0.3$ $0.005\pm 0.005$ $47\pm 1$ $55\pm 1$ $6\times 10^{-6}\pm 7\times 10^{-6}$
gSLOPE $1625$ $231\pm 5$ $453\pm 5$ $26\pm 1$ $36\pm 1$ $36\pm 1$ $0.005\pm 0.006$ $203\pm 6$ $222\pm 6$ $1\times 10^{-6}\pm 1\times 10^{-6}$
gSLOPE $2750$ $1061\pm 15$ $2296\pm 25$ $56\pm 2$ $75\pm 2$ $75\pm 2$ $0.004\pm 0.005$ $270\pm 9$ $350\pm 9$ $5\times 10^{-7}\pm 7\times 10^{-7}$
gSLOPE $3875$ $1765\pm 83$ $2800\pm 113$ $82\pm 3$ $114\pm 3$ $114\pm 3$ $0.006\pm 0.008$ $549\pm 19$ $546\pm 18$ $3\times 10^{-7}\pm 5\times 10^{-7}$
gSLOPE $5000$ $1937\pm 63$ $2422\pm 66$ $102\pm 3$ $147\pm 4$ $147\pm 4$ $0.007\pm 0.012$ $594\pm 21$ $581\pm 19$ $2\times 10^{-7}\pm 3\times 10^{-7}$
SGS $500$ $94\pm 1$ $133\pm 2$ $9\pm 0.2$ $11\pm 0.3$ $12\pm 0.3$ $3\pm 0.2$ $28\pm 1$ $74\pm 3$ $5\times 10^{-10}\pm 4\times 10^{-10}$
SGS $1625$ $416\pm 8$ $1129\pm 19$ $25\pm 1$ $37\pm 1$ $41\pm 1$ $12\pm 0.3$ $62\pm 7$ $511\pm 21$ $2\times 10^{-9}\pm 2\times 10^{-9}$
SGS $2750$ $639\pm 14$ $2137\pm 47$ $62\pm 2$ $82\pm 2$ $89\pm 3$ $20\pm 1$ $80\pm 6$ $791\pm 28$ $2\times 10^{-9}\pm 6\times 10^{-9}$
SGS $3875$ $939\pm 31$ $2862\pm 96$ $93\pm 3$ $124\pm 3$ $136\pm 4$ $30\pm 1$ $112\pm 8$ $1049\pm 34$ $5\times 10^{-9}\pm 1\times 10^{-8}$
SGS $5000$ $1586\pm 66$ $2891\pm 128$ $119\pm 4$ $164\pm 3$ $180\pm 4$ $39\pm 1$ $171\pm 11$ $1118\pm 37$ $2\times 10^{-9}\pm 4\times 10^{-9}$
F.3.2 Additional results for the logistic model
This section presents additional results for the logistic model. Similar trends to the ones observed for the linear model are seen.
Figure A10: The number of groups/variables in $\mathcal{E},\mathcal{A}$ as a function of the regularization path for the logistic model with $p=2750,\rho=0.6,m=197$, shown for gSLOPE and SGS. The results are averaged over $100$ repetitions, with $95\%$ confidence intervals shown. This figure is the equivalent of Figure 2 for the logistic model.
(a) $p=500,\rho=0$
(b) $p=500,\rho=0.9$
(c) $p=5000,\rho=0$
(d) $p=5000,\rho=0.9$
Figure A11: The number of groups/variables in $\mathcal{E},\mathcal{A}$ as a function of the regularization path for the logistic model with SGS and gSLOPE, shown for different values of the correlation ($\rho$) and $p$. The results are averaged over $100$ repetitions, with $95\%$ confidence intervals shown.
Figure A12: Runtime (in seconds) for screening against no screening as a function of $p$, broken down into different correlation cases, for the logistic model. The results are averaged over $100$ repetitions, with standard errors shown.
Table A5: Variable screening metrics for SGS using a logistic model for the simulation study presented in Section 5.1. The number of variables in $\mathcal{A}_{v},\mathcal{S}_{v},\mathcal{E}_{v}$, and $\mathcal{K}_{v}$ are shown. The results are shown for different values of $p$, averaged across $\rho\in\{0,0.3,0.6,0.9\}$. Standard errors are shown.
Method $p$ $\operatorname{card}(\mathcal{A}_{v})$ $\operatorname{card}(\mathcal{S}_{v})$ $\operatorname{card}(\mathcal{E}_{v})$ $\operatorname{card}(\mathcal{K}_{v})$
SGS $500$ $53\pm 2$ $71\pm 3$ $89\pm 4$ $19\pm 1$
SGS $1625$ $157\pm 5$ $248\pm 8$ $291\pm 9$ $44\pm 1$
SGS $2750$ $247\pm 7$ $420\pm 11$ $491\pm 13$ $71\pm 2$
SGS $3875$ $316\pm 9$ $571\pm 14$ $663\pm 16$ $92\pm 3$
SGS $5000$ $375\pm 10$ $717\pm 16$ $824\pm 19$ $107\pm 3$
Table A6: General and group screening metrics for SGS and gSLOPE using logistic models for the simulation study presented in Section 5.1. General metrics: the runtime (in seconds) for screening against no screening, the number of fitting iterations for screening against no screening, and the $\ell_{2}$ distance between the fitted values obtained with screening and no screening. Group screening metrics: the number of groups in $\mathcal{A}_{g},\mathcal{S}_{g},\mathcal{E}_{g}$, and $\mathcal{K}_{g}$. The results are shown for different values of $p$, averaged across $\rho\in\{0,0.3,0.6,0.9\}$. Standard errors are shown.
Method $p$ Runtime (s)
Runtime (s) screen
$\operatorname{card}(\mathcal{A}_{g})$ $\operatorname{card}(\mathcal{S}_{g})$ $\operatorname{card}(\mathcal{E}_{g})$ $\operatorname{card}(\mathcal{K}_{g})$ Num it
Num it screen
$\ell_{2}$ dist
to no screen
gSLOPE $500$ $49\pm 1$ $76\pm 1$ $26\pm 1$ $31\pm 1$ $31\pm 1$ $0.003\pm 0.004$ $31\pm 1$ $40\pm 1$ $4\times 10^{-8}\pm 5\times 10^{-8}$
gSLOPE $1625$ $138\pm 3$ $203\pm 3$ $43\pm 1$ $54\pm 2$ $54\pm 2$ $0.004\pm 0.006$ $78\pm 2$ $79\pm 1$ $1\times 10^{-8}\pm 5\times 10^{-9}$
gSLOPE $2750$ $987\pm 11$ $1641\pm 16$ $72\pm 2$ $95\pm 3$ $95\pm 3$ $0.008\pm 0.013$ $87\pm 2$ $89\pm 1$ $1\times 10^{-8}\pm 5\times 10^{-9}$
gSLOPE $3875$ $1441\pm 26$ $2398\pm 31$ $98\pm 3$ $135\pm 3$ $135\pm 3$ $0.031\pm 0.054$ $95\pm 2$ $98\pm 1$ $7\times 10^{-9}\pm 3\times 10^{-9}$
gSLOPE $5000$ $1454\pm 29$ $2727\pm 40$ $118\pm 3$ $168\pm 4$ $168\pm 4$ $0.022\pm 0.041$ $101\pm 2$ $109\pm 1$ $4\times 10^{-9}\pm 1\times 10^{-9}$
SGS $500$ $118\pm 1$ $113\pm 1$ $28\pm 1$ $33\pm 1$ $38\pm 1$ $8\pm 0.2$ $6\pm 0.3$ $29\pm 1$ $8\times 10^{-9}\pm 7\times 10^{-10}$
SGS $1625$ $248\pm 2$ $538\pm 9$ $50\pm 2$ $59\pm 2$ $64\pm 2$ $11\pm 0.4$ $7\pm 1$ $63\pm 3$ $5\times 10^{-9}\pm 1\times 10^{-9}$
SGS $2750$ $374\pm 2$ $868\pm 12$ $85\pm 3$ $104\pm 2$ $115\pm 3$ $21\pm 1$ $7\pm 0.4$ $57\pm 2$ $3\times 10^{-9}\pm 5\times 10^{-10}$
SGS $3875$ $558\pm 4$ $1280\pm 19$ $116\pm 3$ $148\pm 3$ $164\pm 3$ $30\pm 1$ $8\pm 0.4$ $54\pm 1$ $2\times 10^{-9}\pm 2\times 10^{-10}$
SGS $5000$ $737\pm 5$ $1498\pm 19$ $141\pm 4$ $188\pm 3$ $209\pm 4$ $40\pm 1$ $8\pm 0.4$ $54\pm 1$ $1\times 10^{-9}\pm 2\times 10^{-10}$
F.4 Additional results from the real data analysis
Table A7: Variable screening metrics for SGS applied to real data in 5.2. The number of variables in $\mathcal{A}_{v},\mathcal{S}_{v},\mathcal{E}_{v}$, and $\mathcal{K}_{v}$ are shown. The results are averaged across the nine pathway collections, with standard errors shown.
Dataset Method Input
$\operatorname{card}(\mathcal{A}_{v})$ $\operatorname{card}(\mathcal{S}_{v})$ $\operatorname{card}(\mathcal{E}_{v})$ $\operatorname{card}(\mathcal{K}_{v})$
SGS Cancer $5651$ $71\pm 1$ $211\pm 4$ $212\pm 4$ $1\pm 0.04$
SGS Colitis $10259$ $25\pm 1$ $208\pm 6$ $209\pm 6$ $1\pm 0.07$
Table A8: General and group screening metrics for SGS and gSLOPE applied to real data in 5.2. General metrics: the runtime (in seconds) for screening against no screening, the number of fitting iterations for screening against no screening (with the number of occasions of failed convergence given in brackets), and the $\ell_{2}$ distance between the fitted values obtained with screening and no screening. Group screening metrics: the number of groups in $\mathcal{A}_{g},\mathcal{S}_{g},\mathcal{E}_{g}$, and $\mathcal{K}_{g}$. The results are averaged across the nine pathway collections, with standard errors shown.
Method Dataset Input
$\operatorname{card}(\mathcal{A}_{g})$ $\operatorname{card}(\mathcal{S}_{g})$ $\operatorname{card}(\mathcal{E}_{g})$ $\operatorname{card}(\mathcal{K}_{g})$ Num it
Num it screen (num failed)
$\ell_{2}$ dist
to no screen
gSLOPE Cancer $592$ $66\pm 1$ $120\pm 2$ $120\pm 2$ $0\pm 0$ $1719\pm 63(16)$ $4453\pm 103(186)$ $7\times 10^{-8}\pm 7\times 10^{-9}$
gSLOPE Colitis $655$ $32\pm 1$ $56\pm 2$ $56\pm 2$ $0\pm 0$ $1215\pm 73(28)$ $3904\pm 127(140)$ $3\times 10^{-7}\pm 5\times 10^{-8}$
SGS Cancer $592$ $50\pm 1$ $167\pm 2$ $115\pm 2$ $67\pm 0.8$ $72\pm 3(0)$ $127\pm 4(0)$ $1\times 10^{-8}\pm 1\times 8^{-10}$
SGS Colitis $655$ $18\pm 0.6$ $136\pm 4$ $87\pm 3$ $61\pm 2$ $156\pm 9(0)$ $1348\pm 62(0)$ $4\times 10^{-8}\pm 6\times 10^{-9}$

### v2

@@
 We develop strong screening rules for group-based Sorted L-One Penalized Estimation (SLOPE) models: Group SLOPE and Sparse-group SLOPE.
-The developed rules are applicable for the wider family of group-based OWL models, including OSCAR.
+The developed rules are applicable to the wider family of group-based OWL models, including OSCAR.
 Our experiments on both synthetic and real data show that the screening rules significantly accelerate the fitting process.
@@
 Performing regression and discovering relevant features on these datasets is a challenging task, as classical statistical methods tend to break down.
-The most popular approach to meeting this challenge is the Lasso [30], which has given rise to the general penalized regression framework
-$\hat{\beta}(\lambda)=\operatorname*{arg\,min}_{\beta\in\mathbb{R}^{p}}\left\{f(\beta)+\lambda J(\beta;v)\right\},$ (1)
-where $f$ is the loss function, $J$ is a convex penalty norm, $v\succeq 0$ are penalty parameters, and $\lambda>0$ is the regularization parameter.
-A key aspect of fitting a penalized model is to tune the value of the regularization parameter, which determines the level of sparsity in the fitted model.
-Several approaches exist for tuning this parameter, including cross-validation [8, 15] and exact solution path algorithms [10], but can often be computationally expensive.
-Generally, the model is fit along an $l$-length path $\lambda_{1}\geq\ldots\geq\lambda_{l}\geq 0$.
-Finding approaches for speeding up the model fitting process can lead to large computational savings when tuning.
-Screening rules are popular methods that reduce the dimensions of the input dimensionality by discarding irrelevant features before the fitting process.
-Discarding features, especially in high-dimensional settings, has a tremendous impact on the computational costs.
-Denoting the active set of non-zero parameters at $\lambda_{k+1}$ by $\mathcal{A}_{v}(\lambda_{k+1})=\{i\in\{1,\dots,p\}:\hat{\beta}_{i}(\lambda_{k+1})\neq 0\}$, the goal of a screening rule is to use the solution at $\lambda_{k}$ to recover a screened set of features, $\mathcal{S}_{v}(\lambda_{k+1})$, which is a superset of $\mathcal{A}_{v}(\lambda_{k+1})$.
-The screened set can then be used as input for calculating the fitted values.
-By finding the smallest possible screened set, that still contains the active set, the number of irrelevant features included in the optimization process is reduced.
+The most popular approach to meeting this challenge is the lasso (Tibshirani, 1996), which has given rise to the general penalized regression framework
+$\hat{\beta}(\lambda)\in\operatorname*{arg\,min}_{\beta\in\mathbb{R}^{p}}\left\{f(\beta)+\lambda J(\beta;v)\right\},$ (1)
+where $f$ is a differentiable and convex loss function, $J$ is a convex penalty norm, $v$ are penalty weights, and $\lambda>0$ is the regularization parameter.
+A key aspect of fitting a penalized model is to tune the value of $\lambda$ along an $l$-length path $\lambda_{1}\geq\ldots\geq\lambda_{l}\geq 0$.
+Several approaches exist for tuning this parameter, including cross-validation (Homrighausen and McDonald, 2018; Chetverikov et al., 2021) and exact solution path algorithms (Efron et al., 2004), but these can be computationally expensive.
+Screening rules help alleviate these costs by discarding variables that are inactive at the optimal solution, thus reducing input dimensionality prior to optimization.
+Denote the active set of coefficients at path point $\lambda_{k+1}$, for $k\in[l-1]:=\{1,\ldots,l-1\}$, by $\mathcal{A}_{v}(\lambda_{k+1})=\{i\in[p]:\hat{\beta}_{i}(\lambda_{k+1})\neq 0\}$.
+The goal of a (sequential) screening rule is to use the solution at $\lambda_{k}$ to recover a screened set of features, $\mathcal{S}_{v}(\lambda_{k+1})$, which is a superset of $\mathcal{A}_{v}(\lambda_{k+1})$.
+The screened set is then used as input for calculating the fitted values, leading to significant computational savings.
 There are two types of screening rules: safe and heuristic.
-Safe rules, as suggested by the name, provide guarantees that any variables discarded are in fact inactive at the solution.
-The seminal work of El Ghaoui et al.
-[11] introduced the Safe Feature Elimination Rule (SAFE) for the lasso, in which safe regions were constructed to eliminate non-active features.
-Other examples of safe rules are given in [2, 25, 34], including sample screening [28].
-Heuristic rules, on the other hand, tend to discard more variables, but this increased screening efficiency can lead to active variables being incorrectly discarded.
-Examples of heuristic screening rules are given in [1, 13, 31].
-Amongst the proposed heuristic rules, of particular interest is the strong screening rule framework developed for the lasso [31].
-As strong screening rules can lead to violations, Tibshirani et al.
-[31] complemented their rules with a check of the Karush–Kuhn–Tucker (KKT) [16] optimality conditions to ensure no violations occur.
-There are also hybrid screening regimes, which combine safe and heuristic rules [32, 37].
-A strong screening rule is formulated through the KKT conditions for Equation 1, given by:
+Safe rules are guaranteed to only discard inactive variables and mostly follow the Safe Feature Elimination (SAFE) framework (El Ghaoui et al., 2010), in which safe regions for variables are constructed.
+Other notable examples include Slores (Wang et al., 2014), the dome test (Xiang and Ramadge, 2012), and Dual Polytope Projections (DPP) (Wang et al., 2013), as well as sample screening (Shibagaki et al., 2016) and other examples given in Ogawa et al.
+(2013); Fercoq et al.
+(2015); Atamturk and Gomez (2020).
+Heuristic rules tend to follow the strong screening rule framework, proposed by Tibshirani et al.
+(2010), and discard considerably more variables than safe rules.
+However, they can incorrectly discard active variables, so are complemented by a check of the Karush–Kuhn–Tucker (KKT) (Kuhn and Tucker, 1950) stationarity conditions.
+Other strong rules include Blitz (Johnson and Guestrin, 2015), SIS (Fan and Lv, 2008), and ExSIS (Ahmed and Bajwa, 2019).
+Hybrid schemes exist, using both safe and heuristic rules (Zeng et al., 2021; Wang and Breheny, 2022).
+A strong screening rule is formulated through the KKT stationarity conditions for Equation 1, given by
 $\mathbf{0}\in\nabla f(\beta)+\lambda\partial J(\beta;v).$ (2)
-If the gradient $\nabla f(\hat{\beta}(\lambda_{k+1}))$ was available, the set $\mathcal{A}_{v}(\lambda_{k+1})$ could be identified by checking the subdifferential of the norm at zero.
-That is, identifying for which variables
-$-\nabla f(\beta)\in\lambda\partial J(\mathbf{0};v)=\{x\in\mathbb{R}^{p}:J^{*}(x;\lambda v)\leq 1\},$ (3)
-where $J^{*}$ refers to the dual norm of $J$.
-The subdifferential at zero, $\partial J(\mathbf{0};v)$, can also be seen to be the unit ball of its dual norm.
-As the gradient at $\lambda_{k+1}$ is not available in practice, an approximation of the gradient is used to find a screened subset of the features, $\mathcal{S}_{v}(\lambda_{k+1})$.
-The screened variables are combined with the previously active ones to form the reduced input space, $\mathcal{E}_{v}(\lambda_{k+1})=\mathcal{S}_{v}(\lambda_{k+1})\cup\mathcal{A}_{v}(\lambda_{k})$.
-The fitted values are then calculated by solving Equation 1 using only the variables in $\mathcal{E}_{v}(\lambda_{k+1})$.
-Optimality of the fitted values is checked for using the KKT conditions (Equation 2), with violations added into $\mathcal{E}_{v}$.
-This ensures features are not discarded that should be active.
-1.1 Contributions
-Over the years, several adaptive extensions to the lasso have been proposed, including the Sorted L-One Penalized Estimation (SLOPE) method [4].
+If the gradient were available, the active set could be identified exactly by checking the subdifferential of the norm at zero:
+$\nabla f(\beta)\in\lambda\partial J(\mathbf{0};v)=\{x\in\mathbb{R}^{p}:J^{*}(x;\lambda v)\leq 1\},$ (3)
+where $J^{*}$ is the dual norm of $J$ and $\partial J(\mathbf{0};v)$ is the unit ball of the dual norm.
+So, $J^{*}(\nabla f(\beta);\lambda v)\leq 1$ indicates that $\beta=0$.
+As the gradient at $\lambda_{k+1}$ is not available, a model-specific approximation of the gradient is derived to find a screened subset of the features, $\mathcal{S}_{v}(\lambda_{k+1})$, such that $\mathcal{A}_{v}(\lambda_{k+1})\subset\mathcal{S}_{v}(\lambda_{k+1})$.
+1.1 Screening Approaches for SLOPE
+As the lasso is inconsistent under certain scenarios (Zou, 2006), several adaptive extensions have been proposed, including the Sorted L-One Penalized Estimation (SLOPE) model (Bogdan et al., 2015).
 SLOPE applies the sorted $\ell_{1}$ norm $J_{\text{slope}}(\beta;v)=\sum_{i=1}^{p}v_{i}|\beta|_{(i)}$, where $v_{1}\geq\ldots\geq v_{p}\geq 0,|\beta|_{(1)}\geq\ldots\geq|\beta|_{(p)}$.
 One key advantage of SLOPE is its ability to control the variable false discovery rate (FDR) under orthogonal data.
-Both safe and strong rules have been proposed for SLOPE [3, 12, 17], as well as exact solution path algorithms [9, 24].
-As SLOPE is a non-separable penalty, safe screening rules require repeated screening during the fitting process [17].
-This can be a computational bottleneck, as the calculation of the safe regions required for screening are often expensive.
-Strong screening rules, on the other hand, are usually more time efficient as the extra cost of the screening and KKT checks tend to be outweighed by the significant savings of fitting on a reduced input space [31].
-Motivated by this, we introduce strong screening rules for two group-based extensions of SLOPE: Group SLOPE (gSLOPE) [6] and Sparse-group SLOPE (SGS) [14].
-Group SLOPE (gSLOPE) is an extension of SLOPE for selecting groups, defined by the norm
+Additional powerful properties include: it clusters strongly correlated features, it finds the minimum total squared error loss across different sparsity levels, removing the need for prior knowledge of sparsity, and it is asymptotically minimax (Figueiredo and Nowak, 2014; Su and Candès, 2016).
+All of these useful properties have meant that SLOPE has found widespread use in machine learning and genetics (Gossmann et al., 2015; Virouleau et al., 2017; Kremer et al., 2020; Frommlet et al., 2022; Riccobello et al., 2023).
+Both safe (Bao et al., 2020; Elvira and Herzet, 2021) and strong (Larsson et al., 2020) rules have been proposed for SLOPE, as well as exact solution path algorithms (Nomura, 2020; Dupuis and Tardivel, 2023).
+As SLOPE is a non-separable penalty, safe screening requires repeated screening during optimization, which is expensive due to the repeated dual norm evaluations required for the safe regions (Larsson et al., 2020).
+Group-based SLOPE Models
+In genetics, the analysis of grouped features is frequently encountered, as genes are grouped into pathways for the completion of a specific biological task.
+To use this grouping information, SLOPE has been extended to group (gSLOPE) and sparse-group (SGS) regression.
+For a set of $m$ non-overlapping groups, $\mathcal{G}_{1},\dots,\mathcal{G}_{m}$ of sizes $p_{1},\ldots,p_{m}$, Group SLOPE (gSLOPE) (Gossmann et al., 2015; Brzyski et al., 2019) is given by
 $J_{\text{gslope}}(\beta;w)=\sum_{g=1}^{m}\sqrt{p_{g}}w_{g}\|\beta^{(g)}\|_{2},$ (4)
-such that $\beta^{(g)}\in\mathbb{R}^{p_{g}}$ is a vector of the group coefficients and $p_{g}$ denotes the group sizes [6].
+such that $\beta^{(g)}\in\mathbb{R}^{p_{g}}$ is a vector of the group coefficients.
 The norm has ordered penalty weights $w_{1}\geq\ldots\geq w_{g}\geq 0$ (described in Appendix A.1) which are matched to $\sqrt{p_{1}}\|\beta^{(1)}\|_{2}\geq\ldots\geq\sqrt{p_{m}}\|\beta^{(m)}\|_{2}$.
-Sparse-group SLOPE (SGS) was proposed as an extension to SLOPE and gSLOPE, applying both variable and group penalization for bi-level selection [14].
-For $\alpha\in[0,1]$ with variable weights $v$ and group weights $w$ (described in Appendix B.1), the SGS norm is defined as
+Sparse-group SLOPE (SGS) (Feser and Evangelou, 2023) was further proposed as a convex combination of SLOPE and gSLOPE for concurrent variable and group selection.
+For $\alpha\in[0,1]$, with weights $(v,w)$ (described in Appendix B.1), the norm is given by
 $J_{\text{sgs}}(\beta;\alpha,v,w)=\alpha J_{\text{slope}}(\beta;v)+(1-\alpha)J_{\text{gslope}}(\beta;w).$
-Both of these approaches were developed for the analysis of features that naturally belong to groups.
-This situation is frequently encountered in the analysis of genetics data, where genes are grouped into pathways for the completion of a specific biological task.
-The gSLOPE approach selects groups of features, whereas SGS performs bi-level selection.
-Both approaches have been shown to inherit SLOPE’s ability to control FDR under orthogonal data: gSLOPE controls the group FDR [6] and SGS controls both variable and group FDRs [14].
-Section 2 introduces a sparse-group screening rule framework.
-Based on this, strong screening rules for gSLOPE and SGS are developed and presented in Sections 3 and 4.
-The proofs of the theorems and propositions presented in Sections 2 and 3 are provided in the Appendices A.2 and B.3.
-Through a series of simulations and the analysis of real datasets (Sections 5), the proposed screening rules are shown to significantly reduce computational runtime.
-Additionally, due to the reduced input dimensionality offered by screening rules, issues with convergence in the analysis of large datasets are shown to be alleviated.
-These improvements are achieved without affecting the solution optimality.
-The screening rules proposed in this manuscript are further applicable to group-based Ordered Weighted $\ell_{1}$ (OWL) models.
-SLOPE is a particular case of an OWL model through the definition of its weights.
-As the proposed screening rules only require that the penalty sequences ($v,w$) are ordered, they can be used for any group-based OWL models; for example, the Octagonal Shrinkage and Clustering Algorithm for Regression (OSCAR) model [5].
-Results for group-based OSCAR models are provided in Appendix E.
-Notation.
-The feature space is divided into a set of $m$ non-overlapping groups, where $\mathcal{G}_{1},\dots,\mathcal{G}_{m}$ denote the set of variable indices for each group.
-The set of active groups is denoted by $\mathcal{A}_{g}=\{i\in\{1,\dots,m\}:\|\hat{\beta}^{(i)}\|_{2}\neq 0\}$ whereas $\mathcal{Z}=\{i\in\{1,\dots,m\}:\|\hat{\beta}^{(i)}\|_{2}=0\}$ presents the inactive groups.
-The set of variable indices of the active and inactive groups are denoted by $\mathcal{G}_{\mathcal{A}}$ and $\mathcal{G}_{\mathcal{Z}}$, respectively.
+Both approaches control the FDR under orthogonal data: gSLOPE at the group-level (Brzyski et al., 2019) and SGS at both levels (Feser and Evangelou, 2023).
+SGS has been found to outperform other methods at selection and prediction (Feser and Evangelou, 2023).
+1.2 Contributions
+No screening rules exist for group-based SLOPE models.
+The strong screening rule framework introduced by Tibshirani et al.
+(2010) facilitated the extension from the lasso to the group lasso by requiring only knowledge of the dual norm.
+However, this framework is restricted to separable penalties, making it unsuitable for SLOPE.
+Instead, in Larsson et al.
+(2020), the subdifferential of SLOPE is used to derive strong rules.
+Motivated by this, we propose a new strong screening framework for sparse-group norms (Section 2), used to develop screening for gSLOPE (Section 3) and SGS (Section 4) (with proofs of the results provided in Appendices A.3 and B.3).
+The framework applies two screening layers, drastically reducing dimensionality (Figure 1).
+The screening requires knowledge of the subdifferentials, necessitating a general derivation for gSLOPE (Theorem 3.1).
+The choice of strong screening over safe is motivated by two main reasons.
+First, strong rules discard significantly more variables than safe rules (Tibshirani et al., 2010).
+Second, safe rules require calculating the dual norm set.
+For SGS, this is a sum of convex sets.
+Determining if a point lies in this set requires knowledge of the summation’s decomposability, which is a difficult task (Wang and Ye, 2014).
+This challenge is addressed in Section 4.1.
+The analysis of synthetic and real datasets shows that our proposed screening rules considerably improve runtime (Section 6).
+The reduced input dimensionality from the screening also eases convergence issues with large datasets.
+These improvements are achieved without affecting solution optimality.
+Figure 1: The proportion of variables in $\mathcal{S}_{v}$ relative to $p$ for group-only and bi-level screening applied to SGS, plotted along the regularization path with 95% confidence intervals.
+Synthetic data was generated under a linear model for $p=500,5000$ (Section 6.1), with results averaged over 100 repetitions.
+As the proposed screening rules only require that the penalty sequences ($v,w$) are ordered, they apply to the wider class of group-based Ordered Weighted $\ell_{1}$ (OWL) models.
+Popular special cases of OWL models include SLOPE and the Octagonal Shrinkage and Clustering Algorithm for Regression (OSCAR) model (Bondell and Reich, 2008) (the description of group-based OSCAR models is provided in Section 5).
+Notation
+The sets of active and inactive groups are given by $\mathcal{A}_{g}=\{g\in[m]:\|\hat{\beta}^{(g)}\|_{2}\neq 0\}$ and $\mathcal{Z}=\{g\in[m]:\|\hat{\beta}^{(g)}\|_{2}=0\}$.
+Their corresponding set of variable indices are denoted by $\mathcal{G}_{\mathcal{A}}$ and $\mathcal{G}_{\mathcal{Z}}$.
 The cardinality of a vector $x$ is denoted by $\operatorname{card}(x)$.
 The operators $(\cdot)_{\downarrow}$ and $(\cdot)_{|\downarrow|}$ sort a vector into decreasing and decreasing absolute form.
-We use $\preceq$ to denote element-wise less than or equal to.
-The operator $\mathcal{O}(\cdot)$ returns the index of a vector sorted in decreasing absolute form.
+We use $\preceq$ to denote element-wise inequality signs.
+The operator $\mathcal{O}(\cdot)$ returns the index of a vector sorted into decreasing absolute form.
 The cumulative summation operator applied on a vector is denoted by $\text{cumsum}(x)=[x_{1},x_{1}+x_{2},\ldots,\sum_{i=1}^{\operatorname{card}(x)}x_{i}]$.
-2 Screening rule framework
+2 SPARSE-GROUP STRONG SCREENING
+Sparse-group models, such as SGS and the sparse-group lasso (SGL) (Simon et al., 2013), apply both variable and group penalization so that bi-level screening is possible.
+Safe rules that perform bi-level screening exist for SGL (Wang and Ye, 2014; Ndiaye et al., 2016a), but there are no such strong rules.
+The strong screening framework (Tibshirani et al., 2010) does not extend to sparse-group or non-separable norms, and the strong rule derived for SGL in Liang et al.
+(2022) applies only group-level screening.
+Framework
+We introduce a new sparse-group framework (Algorithm 1), based on the strong framework by Tibshirani et al.
+(2010), for applying strong screening to sparse-group norms, allowing for bi-level screening.
+By applying bi-level screening for SGS, a substantially larger proportion of variables are discarded than with group screening alone (Figure 1).
+First, a screened set of groups is computed, $\mathcal{S}_{g}$.
+An additional layer of screening is then performed to compute $\mathcal{S}_{v}$ using $\mathcal{S}_{g}$.
+This, combined with the previously active variables, forms the reduced input set for fitting, $\mathcal{E}_{v}$.
+KKT checks are performed on $\mathcal{E}_{v}$ to ensure no violations have occurred (Appendices A.4 and B.4).
+Based on this framework, the SGS rules are derived in Section 4.
+The screening for gSLOPE (Section 3) also uses this framework but does not perform the variable screening and instead takes $\mathcal{S}_{v}$ as all variables in the groups of $\mathcal{S}_{g}$.
+The KKT checks are then performed only on the groups.
+Appendix D describes the implementation of the framework for gSLOPE and SGS.
+The main cost of the framework is calculating the fitted values.
+For $t$ iterations of ATOS (a proximal algorithm used to fit SGS, see Section 6.1), the convergence rate is $O(1/t)$ (Pedregosa and Gidel, 2018) and we expect a time complexity of $O(tp^{2})$ for a proximal algorithm (Zhao and Huo, 2023).
 Algorithm 1 Sparse-group screening framework
-  Input: $\lambda\in\mathbb{R}^{l}$, $\mathbf{X}\in\mathbb{R}^{n\times p},y\in\mathbb{R}^{n}$
+  Input: $(\lambda_{1},\ldots,\lambda_{l})\in\mathbb{R}^{l}$, $\mathbf{X}\in\mathbb{R}^{n\times p},y\in\mathbb{R}^{n}$
   for $k=1$ to $l-1$ do
@@
   end for
-  Output: $\hat{\beta}(\lambda_{1}),\ldots,\hat{\beta}(\lambda_{l})\in\mathbb{R}^{p\times l}$
-Sparse-group models apply both variable and group penalization, such that Equation 1 is extended to include an additional convex penalty function.
-Examples of sparse-group models include the Sparse-group Lasso (SGL) [29] and SGS [14].
-Safe screening rules are available for SGL which perform bi-level screening, i.e.
-both group and variable screening [21, 33], but there are no strong rules which perform bi-level screening.
-The strong screening rule framework proposed by Tibshirani et al.
-[31] does not fully extend to sparse-group or non-separable penalties, whilst the strong screening rule derived for SGL in Liang et al.
-[19] applies only group-level screening.
-Framework.
-We introduce a new framework (Algorithm 1) for applying strong screening rules to sparse-group penalties that allow for bi-level screening, based on the framework introduced by Tibshirani et al.
-[31].
-In the sparse-group framework, a screened set of groups is computed, $\mathcal{S}_{g}$, from the full input.
-An additional layer of screening is then performed to compute $\mathcal{S}_{v}$ using $\mathcal{S}_{g}$.
-This forms the reduced input set for fitting, $\mathcal{E}_{v}$.
-The aim of finding $\mathcal{E}_{v}$ is the same as for the lasso and SLOPE strong rules, but the framework first discards irrelevant groups.
-KKT checks are performed on $\mathcal{E}_{v}$ to ensure no violations have occurred.
-The framework applied to gSLOPE and SGS is presented in Appendix D.
-Based on the framework described in Algorithm 1, the screening rules for SGS are derived (Section 4).
-The screening rule for gSLOPE (Section 3) follows a similar framework, where $\mathcal{S}_{v}$ is taken as all variables within the groups of $\mathcal{S}_{g}$.
-The KKT checks are then performed only on the groups.
-By applying bi-level screening for SGS, a substantially larger proportion of variables are discarded, than if just group screening were applied (Figure 1).
-Using only group screening for SGS, without the additional variable screening step, resulted in the number of variables in the screened set scaling poorly with increased dimensionality.
-This demonstrates that bi-level screening effectively manages the scaling of SGS under rising dimensions.
-(a) Applied to synthetic data (Section 5.1). Data generated under a linear model for $p=500,5000$. The results are averaged over $100$ repetitions.
-(b) Applied to two real genetics datasets (Section 5.2), with groups formed using pathways. The results are averaged over the nine pathway collections.
-Figure 1: The proportion of variables in $\mathcal{S}_{v}$ relative to the full input for SGS, shown for group and bi-level screening, plotted as a function of the regularization path, with $95\%$ confidence intervals.
+  Output: $\hat{\beta}(\lambda_{1}),\ldots,\hat{\beta}(\lambda_{l})\in\mathbb{R}^{p}$
 3 GROUP SLOPE
-The strong screening rule for gSLOPE presented in this section is formulated by computing the zero condition of its subdifferential (as per Equation 3) derived in Theorem 3.1.
-To derive the subdifferential, we define the operator
+The strong rule for gSLOPE is formulated by checking the zero condition of the subdifferential (as per Equation 3) derived in Theorem 3.1. To derive the subdifferential, we define the operator
 $[b]_{\mathcal{G},q}:=(p_{1}^{q}\|b^{(1)}\|_{2},\dots,p_{m}^{q}\|b^{(m)}\|_{2})^{\top}.$
@@
 The subdifferential for gSLOPE is given by
-$\partial J_{\text{gslope}}(\beta;w)=\begin{dcases}\{x\in\mathbb{R}^{\operatorname{card}{\mathcal{G}_{\mathcal{Z}}}}:{[x]}_{\mathcal{G}_{\mathcal{Z}},-0.5}\in\partial J_{\text{slope}}(0;w_{\mathcal{Z}})\},&\text{at 0.}\\ w_{g}\sqrt{p_{g}}\frac{\beta^{(g)}}{\|\beta^{(g)}\|_{2}},&\text{otherwise.}\end{dcases}$
-The choice of $q=-0.5$ leads to $J_{\text{gslope}}^{*}(x;w)=J_{\text{slope}}^{*}([x]_{\mathcal{G},-0.5})$, which allows the gSLOPE subdifferential to be written in terms of the SLOPE one [6].
-Combining the KKT conditions (Equation 2) with the gSLOPE subdifferential (Theorem 3.1) reveals that a group is zero if $h(\lambda):=([\nabla f(\hat{\beta}(\lambda))]_{\mathcal{G},-0.5})_{\downarrow}\in\partial J_{\text{slope}}(\mathbf{0};\lambda w_{\mathcal{Z}})$.
-Using the subdifferential of SLOPE [17], this is given by
-$\text{cumsum}(h(\lambda)-\lambda w_{\mathcal{Z}})\preceq\mathbf{0}.$ (5)
+$\partial J_{\text{gslope}}(\beta;w)=\begin{dcases}\begin{array}[]{l}\big{\{}x\in\mathbb{R}^{\operatorname{card}\mathcal{G}_{\mathcal{Z}}}:\\ {[x]}_{\mathcal{G}_{\mathcal{Z}},-0.5}\in\partial J_{\text{slope}}(0;w_{\mathcal{Z}})\big{\}},\end{array}\text{at 0.}&\\ \left\{w_{g}\sqrt{p_{g}}\frac{\beta^{(g)}}{\|\beta^{(g)}\|_{2}}\right\},\;\;\;\;\;\;\;\;\;\text{otherwise.}&\end{dcases}$
+The choice of $q=-0.5$ leads to $J_{\text{gslope}}^{*}(x;w)=J_{\text{slope}}^{*}([x]_{\mathcal{G},-0.5})$, which allows the gSLOPE subdifferential to be written in terms of the SLOPE one (Brzyski et al., 2019).
+Combining the KKT conditions at zero (Equation 3) with the gSLOPE subdifferential (Theorem 3.1) reveals that a group is inactive if $h(\lambda):=([\nabla f(\hat{\beta}(\lambda))]_{\mathcal{G},-0.5})_{\downarrow}\in\partial J_{\text{slope}}(\mathbf{0};\lambda w)$.
+Using the subdifferential of SLOPE (Appendix A.2) (Larsson et al., 2020), this is given by
+$\text{cumsum}(h(\lambda)-\lambda w)\preceq\mathbf{0}.$ (5)
 This condition is checked efficiently using the algorithm proposed for the SLOPE strong rule (Algorithm A1) leading to the strong rule for gSLOPE (Proposition 3.2).
-The algorithm assumes that the indices for the inactive predictors will be ordered last in the input $c$ and the features $|\hat{\beta}|_{\downarrow}$ [17].
+The algorithm assumes that the indices for the inactive predictors will be ordered last in the input $c$ and the features $|\hat{\beta}|_{\downarrow}$ (Larsson et al., 2020).
 Proposition 3.2 (Strong screening rule for gSLOPE).
 Taking $c=h(\lambda_{k+1})$ and $\phi=\lambda_{k+1}w$ as inputs for Algorithm A1 returns a superset $\mathcal{S}_{g}(\lambda_{k+1})$ of the active set $\mathcal{A}_{g}(\lambda_{k+1})$.
-The gradient at path value $k+1$ is not available for the computation of $h(\lambda_{k+1})$, so an approximation is required that does not lead to any violations in Algorithm A1.
-By the cumsum condition in this algorithm, an approximation for a group $g$ is sought such that $h_{g}(\lambda_{k+1})\leq h_{g}(\lambda_{k})+R_{g}$, where $R_{g}\geq 0$ needs to be determined.
+The gradient at path index $k+1$ is not available for the computation of $h(\lambda_{k+1})$, so an approximation is required that does not lead to any violations in Algorithm A1.
+By the cumsum condition in this algorithm, an approximation for a group $g\in[m]$ is sought such that $h_{g}(\lambda_{k+1})\leq h_{g}(\lambda_{k})+R_{g}$, where $R_{g}\geq 0$ needs to be determined.
 An approximation is found by assuming that $h_{g}(\lambda_{k+1})$ is a Lipschitz function of $\lambda_{k+1}$ with respect to the $\ell_{1}$ norm, that is,
 $\left|h_{g}(\lambda_{k+1})-h_{g}(\lambda_{k})\right|\leq w_{g}|\lambda_{k+1}-\lambda_{k}|.$
-By again noting that $J_{\text{gslope}}^{*}(x)=J_{\text{slope}}^{*}([x]_{\mathcal{G},-0.5})$, it can be seen that the assumption is equivalent to those used to derive strong rules for the lasso and SLOPE [17, 31].
+By again noting that $J_{\text{gslope}}^{*}(x)=J_{\text{slope}}^{*}([x]_{\mathcal{G},-0.5})$, it can be seen that the assumption is equivalent to the Lipschitz assumptions used for the lasso and SLOPE strong rules (Tibshirani et al., 2010; Larsson et al., 2020).
 By the reverse triangle inequality,
 $|h_{g}(\lambda_{k+1})|\leq|h_{g}(\lambda_{k})|+\lambda_{k}w_{g}-\lambda_{k+1}w_{g},$
-leading to the choice $R_{g}=\lambda_{k}w_{g}-\lambda_{k+1}w_{g}$ and the gradient approximation strong screening rule.
+leading to the choice $R_{g}=\lambda_{k}w_{g}-\lambda_{k+1}w_{g}$ and the gradient approximation strong screening rule (Proposition 3.3).
+To apply gSLOPE screening in practice (Section 6.1), Proposition 3.4 describes the calculation of the first path value.
 Proposition 3.3 (Gradient approximation strong screening rule for gSLOPE).
-Taking $c=h(\lambda_{k})+\lambda_{k}w-\lambda_{k+1}w$ and $\phi=\lambda_{k+1}w$ as inputs for Algorithm A1, and assuming that for any $k\in\{1,\dots,l-1\}$,
-$\left|h_{g}(\lambda_{k+1})-h_{g}(\lambda_{k})\right|\leq w_{g}|\lambda_{k+1}-\lambda_{k}|,\;\forall g=1,\dots,m,$
+Taking $c=h(\lambda_{k})+\lambda_{k}w-\lambda_{k+1}w$ and $\phi=\lambda_{k+1}w$ as inputs for Algorithm A1, and assuming that for any $k\in[l-1]$,
+$\left|h_{g}(\lambda_{k+1})-h_{g}(\lambda_{k})\right|\leq w_{g}|\lambda_{k+1}-\lambda_{k}|,\;\forall g=[m],$
 and $\mathcal{O}(h(\lambda_{k+1}))=\mathcal{O}(h(\lambda_{k}))$, then the algorithm returns a superset $\mathcal{S}_{g}(\lambda_{k+1})$ of the active set $\mathcal{A}_{g}(\lambda_{k+1})$.
+Proposition 3.4 (gSLOPE path start).
+For gSLOPE, the path value at which the first group enters the model is given by
+$\lambda=\max\left\{\operatorname{cumsum}\bigl{(}([\nabla f(\mathbf{0})]_{\mathcal{G},-0.5})_{\downarrow}\bigr{)}\oslash\operatorname{cumsum}(w)\right\},$
+[…]
 4 SPARSE-GROUP SLOPE
 This section presents the group and variable screening rules for SGS. They are derived using the SGS KKT conditions, formulated in terms of SLOPE and gSLOPE (by the sum rule of subdifferentials):
-$-\nabla f(\beta)\in\lambda\alpha\partial J_{\text{slope}}(\beta;v)+\lambda(1-\alpha)\partial J_{\text{gslope}}(\beta;w).$ (6)
+$\nabla f(\beta)\in\lambda\alpha\partial J_{\text{slope}}(\beta;v)+\lambda(1-\alpha)\partial J_{\text{gslope}}(\beta;w).$ (6)
 4.1 Group Screening
@@
 […]
-The problem reduces to that of the gSLOPE screening rule (Section 3), with input given by $c=([\nabla f(\beta)+\lambda\alpha\partial J_{\text{slope}}(\mathbf{0};v)]_{\mathcal{G},-0.5})_{\downarrow}$ and $\phi=\lambda(1-\alpha)w$.
+$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;-\lambda(1-\alpha)w_{\mathcal{Z}}\Bigr{)}\preceq\mathbf{0}.$ (7)
+The problem reduces to a form similar to the gSLOPE screening rule (Section 3), with inputs for Algorithm A1 given by $c=([\nabla f(\beta)+\lambda\alpha\partial J_{\text{slope}}(\mathbf{0};v)]_{\mathcal{G},-0.5})_{\downarrow}$ and $\phi=\lambda(1-\alpha)w$.
 To determine the form of the quantity $\partial J_{\text{slope}}(\mathbf{0};v)$, the term inside the $[\cdot]_{\mathcal{G}_{\mathcal{Z}},-0.5}$ operator needs to be as small as possible for Equation 7 to be satisfied.
-This term is found to be the soft thresholding operator, $S(\nabla f(\beta),\lambda\alpha):=\text{sign}(\nabla f(\beta))(|\nabla f(\beta)|-\lambda\alpha)_{+}$ (with the derivation presented in Appendix B.2).
-By using the soft-thresholding operator, a valuable connection between SGS and SGL is made, as this operator is used in the gradient update step for SGL [29].
-Such a connection has the potential to lead to new and more efficient optimization approaches for SGS that are more closely related to those used to solve SGL, similar to the recently developed coordinate descent algorithm for SLOPE [18].
-Using this function, the (non-approximated) strong group screening rule for SGS is given in Proposition B.1.
-Using a similar Lipschitz assumption as for the gSLOPE rule gives the gradient approximation strong group screening rule for SGS (Proposition 4.1).
+This term is found to be the soft thresholding operator, $S(\nabla f(\beta),\lambda\alpha v):=\text{sign}(\nabla f(\beta))(|\nabla f(\beta)|-\lambda\alpha v)_{+}$ by Lemma 4.1 (see Appendix B.2 for the proof).
+Lemma 4.1.
+In Equation 7, choosing $\partial J_{\text{slope}}(\mathbf{0};v)=S(\nabla f(\beta),\lambda\alpha v)$ minimises $[\nabla f(\beta)+\lambda\alpha\partial J_{\text{slope}}(\mathbf{0};v)]_{\mathcal{G}}$.
+By using the soft-thresholding operator, a valuable connection between SGS and SGL is found, as the operator is used in the gradient update step for SGL (Simon et al., 2013).
+This connection has the potential to lead to new and more efficient optimization approaches for SGS that are more closely related to those used to solve SGL, similar to the recent coordinate descent algorithm for SLOPE (Larsson et al., 2022).
+Using this operator, the (non-approximated) strong group screening rule for SGS is shown in Proposition B.1.
+Applying a similar Lipschitz assumption as for the gSLOPE rule gives the gradient approximation strong group screening rule for SGS (Proposition 4.2).
 Proposition 4.2 (Gradient approximation strong group screening rule for SGS).
-Let $\tilde{h}(\lambda):=([S(\nabla f(\hat{\beta}(\lambda)),\lambda\alpha)]_{\mathcal{G},-0.5})_{\downarrow}$.
-Taking $c=\tilde{h}(\lambda_{k})+\lambda_{k}(1-\alpha)w-\lambda_{k+1}(1-\alpha)w$ and $\phi=\lambda_{k+1}(1-\alpha)w$ as inputs for Algorithm A1, and assuming that for any $k\in\{1,\dots,l-1\}$,
-$\left|\tilde{h}_{g}(\lambda_{k+1})-\tilde{h}_{g}(\lambda_{k})\right|\leq(1-\alpha)w_{g}|\lambda_{k+1}-\lambda_{k}|,\forall g=1,\dots,m,$
-and $\mathcal{O}(\tilde{h}(\lambda_{k+1}))=\mathcal{O}(\tilde{h}(\lambda_{k}))$, then the algorithm returns a superset $\mathcal{S}_{g}(\lambda_{k+1})$ of $\mathcal{A}_{g}(\lambda_{k+1})$.
+Let $\tilde{h}(\lambda):=([S(\nabla f(\hat{\beta}(\lambda)),\lambda\alpha v)]_{\mathcal{G},-0.5})_{\downarrow}$.
+Taking $c=\tilde{h}(\lambda_{k})+\lambda_{k}(1-\alpha)w-\lambda_{k+1}(1-\alpha)w$ and $\phi=\lambda_{k+1}(1-\alpha)w$ as inputs for Algorithm A1, and assuming that for any $k\in[l-1]$,
+$\left|\tilde{h}_{g}(\lambda_{k+1})-\tilde{h}_{g}(\lambda_{k})\right|\leq(1-\alpha)w_{g}|\lambda_{k+1}-\lambda_{k}|,\;\forall g=[m],$
+and $\mathcal{O}(\tilde{h}(\lambda_{k+1}))=\mathcal{O}(\tilde{h}(\lambda_{k}))$, then the algorithm returns a superset $\mathcal{S}_{g}(\lambda_{k+1})$ of the active set $\mathcal{A}_{g}(\lambda_{k+1})$.
 4.2 Variable Screening
-Whilst the group screening for SGS returns a superset, $\mathcal{S}_{g}$, of the active group set, exploiting the sparse-group penalization of SGS allows for the screened set to be further reduced by also applying variable screening.
-The KKT conditions (Equation 6) for a zero variable, $j\in\mathcal{G}_{g}$, in an active group, $g\in\mathcal{A}_{g}$, are
-$-\nabla_{j}f(\beta)\in\lambda\alpha\partial J_{\text{slope}}(\mathbf{0};v_{j}).$ (8)
+By exploiting the sparse-group norm of SGS, the input dimensionality can be reduced further with a second layer of variable screening.
+The KKT conditions (Equation 6) for zero variables in active groups are
+$\nabla_{\mathcal{G}_{\mathcal{A}_{g}}}f(\beta)\in\lambda\alpha\partial J_{\text{slope}}(\mathbf{0};v_{\mathcal{G}_{\mathcal{A}_{g}}}).$ (8)
 The gSLOPE subdifferential term vanishes as the numerator is zero in Theorem 3.1.
-The problem reduces to that of SLOPE variable screening, applied only to the variables that belong to groups in $\mathcal{A}_{g}$ and scaled by $\alpha$.
-The gradient approximated rule is formalized in Proposition 4.2 (the non-approximated version is presented in Proposition B.2).
+The problem reduces to that of SLOPE screening, applied to the variables in groups in $\mathcal{A}_{g}$ and scaled by $\alpha$.
+The gradient approximated rule is shown in Proposition 4.3 (see Proposition B.2 for the non-approximated version).
 Proposition 4.3 (Gradient approximation strong variable screening rule for SGS).
-Let $\bar{h}(\lambda)=(\nabla f(\hat{\beta}(\lambda)))_{|\downarrow|}$.
-Taking $c=|\bar{h}(\lambda_{k+1})|+\lambda_{k}\alpha v-\lambda_{k+1}\alpha v$ and $\phi=\lambda_{k+1}\alpha v$ for only the variables in the groups in $\mathcal{A}_{g}(\lambda_{k+1})$ as inputs for Algorithm A1 , and assuming that for any $k\in\{1,\dots,l-1\}$,
+Let $\bar{h}(\lambda):=(\nabla f(\hat{\beta}(\lambda)))_{|\downarrow|}$.
+Taking $c=|\bar{h}(\lambda_{k+1})|+\lambda_{k}\alpha v-\lambda_{k+1}\alpha v$ and $\phi=\lambda_{k+1}\alpha v$ for the variables in the groups in $\mathcal{A}_{g}(\lambda_{k+1})$ as inputs for Algorithm A1, and assuming that for any $k\in[l-1]$,
 $\left|\bar{h}_{j}(\lambda_{k+1})-\bar{h}_{j}(\lambda_{k})\right|\leq\alpha v_{j}|\lambda_{k+1}-\lambda_{k}|,\forall j\in\mathcal{G}_{\mathcal{A}_{g}(\lambda_{k+1})},$
@@
 In practice $\mathcal{A}_{g}(\lambda_{k+1})$ is not available, as this is exactly what we are trying to superset with any screening rule.
-However, Proposition 4.1 states that it is contained within $\mathcal{S}_{g}(\lambda_{k+1})$, so that the screened set can be used instead and complemented with KKT checks to ensure there are no violations, which would be performed anyway.
-For the checks, the group KKT conditions (Equation 6) are checked.
-For any active groups or groups violating the check (indicating that they should be active), the variable KKT conditions (Equation 8) are checked on the variables within those groups, to identify those that need to be added to $\mathcal{E}_{v}$.
+However, Proposition 4.2 guarantees that it is contained in $\mathcal{S}_{g}(\lambda_{k+1})$ so that this can be used instead.
+To apply SGS in practice (Section 6.1), Proposition 4.4 describes the calculation of the first path value.
+Proposition 4.4 (SGS path start).
+For SGS, the path value at which the first variable enters the model is
+$\lambda=\max\{\operatorname{cumsum}(|\nabla$ $f(\mathbf{0})|_{\downarrow})\oslash$
+$\operatorname{cumsum}((1-\alpha)\tau\omega-\alpha v)\},$
+where $\tau$ and $\omega$ are expanded vectors of the group sizes ($\sqrt{p_{g}})$ and penalty weights ($w_{g}$) to $p$ dimensions.
+5 GROUP-BASED OWL
+The screening rule framework presented are also applicable to the wider class of OWL models. The Ordered Weighted $\ell_{1}$ (OWL) framework is defined as (Zeng and Figueiredo, 2014a)
+$\hat{\beta}\in\operatorname*{arg\,min}_{\beta\in\mathbb{R}^{p}}\left\{\nabla f(\beta)+\lambda J_{\text{owl}}(\beta;v)\right\},$
+where $J_{\text{owl}}(\beta;v)=\sum_{i=1}^{p}v_{i}|\beta|_{(i)}$, $|\beta|_{(1)}\geq\ldots\geq|\beta|_{(p)}$, and $v$ are non-negative non-increasing weights.
+SLOPE is a special case of OWL where the weights are taken to be the Benjamini-Hochberg critical values (Bogdan et al., 2015).
+Octagonal Shrinkage and Clustering Algorithm for Regression (OSCAR) (Bondell and Reich, 2008) is a further special case of OWL (often referred to as OWL with linear decay) where for a variable $i\in[p]$, the weights are taken to be $v_{i}=\sigma_{1}+\sigma_{2}(p-i)$, and $\sigma_{1},\sigma_{2}$ are to be set.
+In Bao et al.
+(2020) they are set to $\sigma_{1}=d_{i}\|\mathbf{X}^{\top}y\|_{\infty},\sigma_{2}=\sigma_{1}/p$, where $d_{i}=i\times e^{-2}$, $\mathbf{X}^{\top}\in\mathbb{R}^{n\times p}$ is the design matrix, and $y\in\mathbb{R}^{n}$ is the response vector.
+Group OSCAR (gOSCAR) and Sparse-group OSCAR (SGO) are defined using the frameworks provided by gSLOPE (Brzyski et al., 2019) and SGS (Feser and Evangelou, 2023), respectively, but instead use the OSCAR weights (see Appendix E.1).
 6 RESULTS
-In this section, the effectiveness of the screening rules for gSLOPE and SGS is illustrated through the analysis of both synthetic (Section 5.1) and real data (Section 5.2).
-Any mentions of $\mathcal{E},\mathcal{A}$ in reference to group and variable metrics refers to $\mathcal{E}_{g},\mathcal{A}_{g}$ and $\mathcal{E}_{v},\mathcal{A}_{v}$ respectively.
-For SGS, $\mathcal{E}_{g}$ is the groups with representation in $\mathcal{E}_{v}$ (the groups fitted).
-Computational information is presented in Appendix F.1.
+This section illustrates the effectiveness of the screening rules for gSLOPE and SGS using synthetic (Section 6.1) and real (Section 6.2) data.
+References to $\mathcal{E},\mathcal{A}$ for group and variable metrics denote $\mathcal{E}_{g},\mathcal{A}_{g}$ and $\mathcal{E}_{v},\mathcal{A}_{v}$.
+For SGS, $\mathcal{E}_{g}$ represents groups with members in $\mathcal{E}_{v}$.
 6.1 Synthetic Data Analysis
-Set up.
-For the synthetic data, a multivariate Gaussian design matrix, $\mathbf{X}\sim\mathcal{N}(\mathbf{0},\boldsymbol{\Sigma})\in\mathbb{R}^{400\times p}$, was used and the within-group correlation set to $\Sigma_{i,j}=\rho$, where $i$ and $j$ belong to the same group.
+Set up
+A multivariate Gaussian design matrix, $\mathbf{X}\sim\mathcal{N}(\mathbf{0},\boldsymbol{\Sigma})\in\mathbb{R}^{400\times p}$, was used and the within-group correlation set to $\Sigma_{i,j}=\rho$, where $i$ and $j$ belong to the same group.
 The correlation and number of features were varied between $\rho\in\{0,0.3,0.6,0.9\}$ and $p\in\{500,1625,2750,3875,5000\}$, producing $20$ simulation cases.
@@
 Groups of sizes between $3$ and $25$ were considered, of which $15\%$ were set to active.
-Within each group, $30\%$ of the variables were set to active with the signal sampled from $\beta\sim\mathcal{N}(0,5)$.
-gSLOPE and SGS were each fit along a log-linear path of $l=50$ regularization parameters using warm starts.
-The first path value $\lambda_{1}$ was set to the exact value at which the first group enters the model (given in Appendix A.3 for gSLOPE and Appendix B.4 for SGS).
-The final regularization value was set to $\lambda_{50}=0.05\lambda_{1}$.
-The data was $\ell_{2}$ standardized and for the linear model an intercept was fitted.
+Within each active group, $30\%$ of the variables were set to active with signal $\beta\sim\mathcal{N}(0,5)$.
+Figure 2: The proportion of groups/variables in $\mathcal{E},\mathcal{A}$, relative to the input, for both gSLOPE and SGS as a function of the path for the linear model with $p=2750,\rho=0.6,m=197$.
+The results are averaged over $100$ repetitions, with $95\%$ confidence intervals shown.
+gSLOPE and SGS were fit along a log-linear path of $50$ regularization parameters using warm starts, beginning at $\lambda_{1}$ (from Propositions 3.4 and 4.4), and ending at $\lambda_{50}=0.05\lambda_{1}$.
+The data was $\ell_{2}$ standardized and for the linear model an intercept was used.
 Both models had FDR-control parameters set to $0.05$, and $\alpha=0.95$ for SGS.
-Even though the models were fit using the adaptive three operator splitting algorithm (ATOS) [26], it should be noted that the screening rules are independent of the choice of fitting algorithm.
-Primarily, the results for the linear model are presented, and the corresponding plots for the logistic model are presented in Appendix F.3.2.
-Screening efficiency.
-By comparing the sizes of the fitting set $(\mathcal{E})$ to the active set ($\mathcal{A})$, we can understand the difference between the minimum dimensionality needed against the dimensionality provided by the screening rules.
-As expected, these sizes increase as $\lambda$ decreases (Figure 2).
-In particular, the gap between sets widens as more features enter the model.
-The size of $\mathcal{E}$ remains far below the size of the full input space even at the lowest value of $\lambda$, showing that the screening rules have a benefit along the whole path.
-This finding is found to be independent of $\rho$, $p$, and model fitted (see Appendix F.3).
-Figure 2: The number of groups/variables in $\mathcal{E},\mathcal{A}$ for both gSLOPE and SGS as a function of the regularization path for the linear model with $p=2750,\rho=0.6,m=197$.
-The results are averaged over $100$ repetitions, with the shaded regions corresponding to $95\%$ confidence intervals.
-Figure 3: The proportion of groups/variables in $\mathcal{E},\mathcal{A}$, relative to the full input, shown for gSLOPE and SGS.
-This is shown as a function of the correlation ($\rho$), averaged over all cases of the input dimension ($p$), with $100$ repetitions for each $p$, for both linear and logistic models, with standard errors shown.
-The performance of the screening rules are generally similar across linear and logistic models (Figure 3).
-As the within-group correlation increases, the size of both the screened and active sets decreases, causing the variables to move in the same direction within the same group.
-Due to the shape of the convex penalty of SLOPE, highly correlated features are clustered together [36].
-This leads to smaller active sets, and in turn, also smaller screened sets.
-Additionally, it can be seen that the size difference between the screened and active sets decreases with correlation.
-Across the different correlation values, the linear model had smaller set sizes than the logistic one.
-Runtime performance.
-A key metric of performance for a screening rule is the time taken to fit a path of models.
-Increasing the value of $p$ demonstrates the computational cost reduction of applying a screening rule (Figure 4).
-For the smaller values of $p$, the runtime tends to be similar whether screening is applied or not.
-However, once the input dimension increases, the benefit of screening can be seen clearly.
-Additionally, increasing the within-group correlation tends to lead to smaller active and screened sets, but the benefit of screening appears to be robust under different correlations.
-Figure 4: Runtime (in seconds) for fitting $50$ models along a path, shown for screening against no screening as a function of $p$, broken down into different correlation cases, for the linear model.
-The results are averaged over $100$ repetitions, with standard errors shown.
-Figure 5: Runtime (in seconds) for fitting $100$ models along a path for screening against no screening for gSLOPE and SGS, averaged across the nine collections of pathways for each dataset, with standard errors shown.
-A clear improvement with regards to runtime can be seen aggregating the results across all simulation cases (Table 1).
-In particular, for both models considered, screening halves the runtime for fitting $50$ models for SGS.
-5.2 Real data experiments
+The models were fit using the adaptive three operator splitting (ATOS) algorithm (Pedregosa and Gidel, 2018), although the screening rules can be applied with any fitting algorithm.
+Additional computational details are in Appendix F.1.
+Primarily, the results for the linear model are presented, and the results for the logistic model are in Appendix F.3.2. The simulations were repeated for group-based OSCAR models (Appendix E).
 Table 1: Runtime (in seconds) for fitting $50$ models along a path, shown for screening against no screening, for the linear and logistic models.
@@
 SGS Logistic $407\pm 2$ $859\pm 6$
-Table 2: The sizes of the sets $\mathcal{A}$ and $\mathcal{E}$ (with standard errors), averaged across $100$ path values and pathway collections for each of the two datasets.
-SGS var/grp shows the corresponding metrics for $\mathcal{A}_{v},\mathcal{E}_{v}$ and $\mathcal{A}_{g},\mathcal{E}_{g}$ for the bi-level screening procedure.
-Method Dataset $\operatorname{card}(\mathcal{A})$ $\operatorname{card}(\mathcal{E})$
-gSLOPE Cancer $66\pm 1$ $120\pm 2$
-gSLOPE Colitis $32\pm 1$ $56\pm 2$
-SGS grp Cancer $50\pm 1$ $115\pm 2$
-SGS grp Colitis $18\pm 1$ $87\pm 3$
-SGS var Cancer $71\pm 1$ $212\pm 4$
-SGS var Colitis $25\pm 1$ $209\pm 6$
-Datasets.
-The screening rules were applied to two gene expression datasets with binary responses.
-The datasets were previously analyzed to measure the predictive performance of sparse-group regression models on large datasets [14, 29].
-The cancer dataset contains gene data for $60$ breast cancer patients who had been treated with tamoxifen for 5 years, classified into binary labels depending on whether the cancer recurred [20].
-The colitis dataset uses transcriptional profiles in peripheral blood mononuclear cells of $127$ patients to determine which patients have an inflammatory bowel disease (ulcerative colitis or Crohn’s).
-Of the $127$ samples, $42$ of them are controls [7].
-The genes of both datasets were assigned to pathways (groups) that were obtained from the Molecular Signature Database.
-[^1] https://www.gsea-msigdb.org/gsea/msigdb/human/collections.jsp Nine collections of pathways were downloaded and as the pathways included different genes, nine unique design matrices for each dataset were created, each with a unique grouping structure.
-This gave $18$ design matrices in total to which the screening rules were applied.
-Both gSLOPE and SGS were fitted with their FDR-control parameters set to $0.01$ and for SGS $\alpha=0.99$.
-Each model was fitted along a path of $100$ regularization parameters, with $\lambda_{1}$ set as described in Appendices A.3 and B.4, and $\lambda_{100}=0.01\lambda_{1}$.
-The data was $\ell_{2}$ standardized and no intercept was used.
-Results.
-For both gSLOPE and SGS, the screening rules lead to considerably faster runtimes (Figure 5).
-In particular, the use of the screening rules significantly benefits the gSLOPE approach.
-With the reduced input feature space, any issues with convergence of the fitting algorithm are substantially decreased.
-For the cancer dataset, $186$ model fits (out of the $900$) did not reach convergence with no screening, compared to only $16$ with screening.
-A similar trend occurred for the colitis dataset.
+Screening Efficiency
+By comparing the sizes of the fitting set $(\mathcal{E})$ to the active set ($\mathcal{A})$, the screening rules are found to be efficient in providing dimensionality reduction close to the minimum possible (the active set size) (Figures 2 and A9).
+As expected, the sets increase in size as $\lambda$ decreases, and the difference in size between the sets remains stable along the path, decreasing towards the termination point.
+The size of the fitting set remains far below the input size across the whole path, showing the benefit of the screening.
+This is found to be true for any correlation, input dimensionality, and model considered (Appendix F.3).
+The screening rules perform well for linear and logistic models (Figure 3), showing robust dimensionality reduction for all correlation cases considered.
+As the correlation increases, the signal concentrates in fewer groups, causing the active group set to decrease in size.
+SLOPE models deal well with highly correlated features, as the sorted norm clusters them together (Zeng and Figueiredo, 2014b).
+The screening rules are found to efficiently reduce the input dimensionality on average across all cases considered (Table 2).
+Figure 3: The proportion of groups/variables in $\mathcal{E},\mathcal{A}$, relative to the full input, shown for gSLOPE and SGS.
+This is shown as a function of the correlation ($\rho$), averaged over all cases of the input dimension ($p$), with $100$ repetitions for each $p$, for both linear and logistic models, with standard errors shown.
+Runtime Performance
+A key metric of performance for a screening rule is the time taken to fit a path of models.
+Figure 4 shows the significant runtime improvements our screening rules provide as a function of increasing input dimensionality.
+The gain is observed to be substantial under all correlation cases.
+Applying screening improves the scaleability of applying gSLOPE and SGS models to larger datasets.
+The clear benefit and robustness of our screening method, with regards to runtime, can be seen by aggregating the results of all the simulation cases (Table 1).
+For both models, in the linear and logistic cases, screening substantially improves the computational cost, halving the runtime for SGS models.
+6.2 Real Data Experiments
+Datasets
+The screening rules were applied to seven real gene datasets, of different response types and dimensionality.
+Two of the datasets, carbotax and sheetz, had a continuous response so were fit using a linear model.
+For these, the groups were generated using K-means clustering (Lloyd, 1982).
+The remaining five (adenoma, cancer, celiac, colitis, and tumour), had binary labels, so a logistic model was used.
+For these, the design matrices contained gene expression data downloaded from NCBI’s GEO database (Edgar et al., 2002), so the genes could be assigned to pathways (groups) using the C3 regulatory target gene sets from MSigDB (Subramanian et al., 2005; Liberzon et al., 2011).
+All datasets were high-dimensional.
+See Appendix F.4 for full details.
+Both gSLOPE and SGS were fit with their FDR-control parameters set to $0.01$ and for SGS $\alpha=0.99$.
+Each model was applied along a path of $100$ regularization parameters, with $\lambda_{100}=0.01\lambda_{1}$.
+Table A1 describes the algorithmic parameters used for ATOS.
+Figure 4: Runtime (in seconds) for fitting $50$ models along a path, shown for screening against no screening as a function of $p$, broken down into different correlation cases, for the linear model.
+The results are averaged over $100$ repetitions, with standard errors shown.
+Results
+For both gSLOPE and SGS, the screening rules led to considerably faster runtimes for all datasets (Figure 5).
+The screening was more effective for SGS under a linear model and for gSLOPE under a logistic model.
+The active set sizes for gSLOPE tended to be smaller for the logistic models, allowing for larger feature reduction (Table 2).
+For SGS, this trend went the other way; the active sets were smaller for the continuous datasets.
+Another explanation for the reduced screening efficiency for gSLOPE under linear models was due to the grouping structure.
+For the continuous responses, the groups were generated using K-means clustering, leading to fewer groups and less opportunity for dimensionality reduction (Table A8).
+With fewer groups, gSLOPE is less likely to discard full groups, as they may contain some signal.
+The feature reduction provided by our screening rules led to the alleviation of convergence issues (Table F.5).
+SGS failed to converge for three datasets without screening and none with screening.
+gSLOPE experienced failed convergences across all datasets without screening, while with screening, it only failed to converge for three datasets, each showing fewer instances of failure.
 As gSLOPE applies no variable penalization, it is forced to fit all variables within a group.
 For datasets with large groups, such as those considered here, this leads to a problematic fitting process which can include many noisy variables.
-The analysis of the real data further illustrates the benefits of the bi-level screening to the runtime and performance of SGS.
-Figure 1(b) illustrates that for these large genetics datasets, the bi-level screening allows the input dimensionality for SGS to be reduced to a much greater extent than by just group screening.
-Both screening rules can drastically reduce the input dimensionality as seen in Table 2.
-For the colitis dataset, the SGS screening rule reduces the input dimensionality to just $2\%$ of the full space.
-This comes without the cost of solution consistency.
-The estimated model coefficients with and without screening were very close to each other, with their $\ell_{2}$ distances of order $10^{-8}$ (Appendices F.2 and F.4).
+Our screening rules help gSLOPE partially overcome this issue, leading to large computational savings and better solution optimality.
+The analysis of the real data further illustrates the benefits of the bi-level screening to the runtime and performance of SGS (for synthetic data, see Figure 1).
+Figure 6 illustrates that for the cancer and celiac datasets, the bi-level screening allows the input dimensionality for SGS to be reduced to a much greater extent than by just group screening (see Figure A17 for the other datasets).
+Both screening rules drastically reduce the input dimensionality on all the real datasets (Table 2).
+On average, for the logistic real data models, the SGS screening rules reduced the input to just $7\%$ of the total space.
+This comes without affecting solution consistency (Appendices F.2 and A16).
+Figure 5: The ratio of no screen time to screen time ($\uparrow$) of gSLOPE and SGS applied to the real datasets, for fitting $100$ path models, split into response type.
+The horizontal grey line represents no screening improvement.
+Table 2: The cardinality of the active ($\mathcal{A}$) and fitting ($\mathcal{E}$) sets for gSLOPE and SGS, averaged across all synthetic and real data cases, split into model type.
+For gSLOPE, the cardinality is for the group sets, and for SGS the variable ones.
+Dim.
+is the input dimensionality.
+For synthetic, it was $p=2750$ and $m=210$.
+Synthetic Real
+[…]
+gSLOPE Lin. $55\scriptstyle{\pm 1}$ $76\scriptstyle{\pm 1}$ $112\scriptstyle{\pm 3}$ $168\scriptstyle{\pm 3}$ $240$
+gSLOPE Log. $71\scriptstyle{\pm 1}$ $97\scriptstyle{\pm 1}$ $49\scriptstyle{\pm 2}$ $83\scriptstyle{\pm 3}$ $1634$
+SGS Lin. $178\scriptstyle{\pm 3}$ $364\scriptstyle{\pm 6}$ $378\scriptstyle{\pm 37}$ $1139\scriptstyle{\pm 87}$ $14488$
+SGS Log. $230\scriptstyle{\pm 3}$ $472\scriptstyle{\pm 6}$ $526\scriptstyle{\pm 5}$ $979\scriptstyle{\pm 13}$ $13734$
+For additional metrics and comparison to the SLOPE strong rule for the real datasets, see Figures A14, A15, and A16.
+Our rules for gSLOPE and SGS are found to offer similar levels of runtime improvements and feature reduction to that of SLOPE (Larsson et al., 2020).
+Figure 6: The proportion of variables in $\mathcal{S}_{v}$ relative to the full input for group-only and bi-level screening applied to SGS, plotted along the regularization path for the cancer and celiac datasets.
 6.3 KKT Violations
-The screening rules rely on assumptions that can fail.
-In these cases, KKT checks are implemented to ensure that no active variables are excluded from $\mathcal{E}_{v}$.
-A KKT violation occurs when a variable is not included in the screening set, $\mathcal{S}_{v}$, but should be according to the KKT optimality conditions (Equation 2), in which case it is added into the fitting set, $\mathcal{E}_{v}$.
+As with any strong screening rule, our approach depends on assumptions that may fail.
+When this happens, KKT checks are used to ensure no active variables are excluded, with violations added to $\mathcal{E}_{v}$.
+For SGS, KKT violations occur at a variable-level (Appendix B.4) and for gSLOPE at a group-level (Appendix A.4).
 KKT violations are very rare for gSLOPE (Figure 7), occurring on the simulated data infrequently toward the start of the path.
-On the real data, no KKT violations occurred for gSLOPE.
-(a) Group violations for gSLOPE and variable violations for SGS, fitted to linear models, averaged over all cases of $p$ and the correlation ($\rho$).
-(b) Variable violations for SGS applied to the cancer and colitis datasets, averaged over the nine pathway collections.
-Figure 6: The proportion of KKT violations relative to the full input space, as a function of the regularization path.
-For SGS, KKT violations are more common, as the additional layer of screening carries with it additional assumptions.
-Additionally, the subdifferential term in Equation 7 was chosen to be as small as possible, leading to tighter screened sets.
-The increased violations are reflective of this.
-On both the simulated and real datasets, SGS recorded an increasing number of KKT violations as the fitted models became denser (Figure 6).
-The shape of the increasing number of KKT violations mirrors the log-linear shape of the regularization path.
-The trend of increasing KKT violations with denser models is also observed for the SLOPE strong screening rule [17].
-Despite the increased number of KKT violations for SGS, which amounts to the additional computational cost of performing the KKT checks and repeated sub-fitting in Algorithm 1, the overall screening procedure still results in substantial runtime improvements, as evidenced in Tables 1 and 2.
-This shows that screening rules can provide improvements to computational cost, even with assumptions that are often violated.
+On the real data, only a single dataset had violations, and the number of violations was very small (Table F.5).
+For SGS, KKT violations are more common, but still infrequent (Figure 7 and Tables A9) due to additional assumptions in the second layer of screening.
+In Equation 7, minimizing the subdifferential term leads to tighter screened sets, contributing to these violations.
+In Figure 7, the number of violations increases as a function of the model density for SGS, mirroring the log-linear shape of the regularization path.
+This pattern is also seen in the strong rule for SLOPE (Larsson et al., 2020).
+Despite these violations and the computational cost of the checks and refitting (Algorithm 1), the overall screening process yields significant runtime improvements in all cases (Table 1 and Figure 5).
+Figure 7: The proportion of KKT violations relative to the full input for gSLOPE and SGS, under linear models, averaged over all synthetic data cases.
+This is shown as a function of the regularization path.
 7 DISCUSSION
-In this manuscript, we have developed strong screening rules for group-based SLOPE models: Group SLOPE and Sparse-group SLOPE, neither of which have any previous screening rules.
+In this manuscript, we have developed strong screening rules for group-based SLOPE models using our new sparse-group strong screening framework: Group SLOPE and Sparse-group SLOPE, neither of which have any previous screening rules.
+Our proposed screening rules differ from the existing SLOPE strong rule both in construction and in outcome.
 The screening rule for gSLOPE screens out irrelevant groups before fitting.
-The screening rules for SGS perform bi-level screening, based on our proposed sparse-group screening framework.
-The two proposed screening rules differ from the existing SLOPE screening rules both in construction and in final outcome.
-gSLOPE only performs screening through the groups, forcing it to keep all variables within the screened groups, even noisy ones, in the optimization process.
-SGS, similarly to SLOPE and in contrast to gSLOPE, performs variable screening after its group screening step.
-As additional assumptions are required for SGS to perform bi-level selection, its computational efficiency is not as effective as the one observed with the SLOPE screening rules [17].
-The superiority of SGS in variable selection and predictive performance when group information is available overcomes the additional computational burden [14].
+The screening rules for SGS perform bi-level screening.
+Our rules apply to the wider class of OWL models, including group-based OSCAR models.
+SLOPE models are finding increasing use in genetics and machine learning, with SGS found to have superior disease prediction performance over other penalized methods (Feser and Evangelou, 2023).
+Our screening rules will make the group-based versions more accessible by reducing their computational burden.
+This will allow practitioners to utilize their FDR properties, facilitating their widespread use across various fields.
 Through comprehensive analysis of synthetic and real data, we illustrate that the screening rules lead to dramatic improvements in the runtime of gSLOPE and SGS models, as well as for group-based OSCAR models (Appendix E).
 This is achieved without affecting model accuracy.
-This is particularly important in datasets where $p\gg n$, such as genetics ones, which is the main motivation behind SLOPE [4].
-The screening rules presented in this manuscript allow group-based SLOPE, and by extension, group-based OWL models, to achieve computational fitting times more in line with lasso-based models, making them more widely accessible.
-Limitations.
-One of the limitations of our proposed screening rules for SGS is the number of assumptions made.
-A future direction includes the exploration of alternative strong screening rules that require fewer assumptions.
-Alternatively, the development of safe rules, which guarantee that only inactive variables are discarded, could be an avenue for future exploration, potentially incorporating safe and strong rules together in a hybrid rule [32, 37].
-A comparison between safe and the proposed strong rules developed in this manuscript could provide further insight into which type of screening is most beneficial for non-separable penalties.
-Appendix
-[…]
+This is particularly important in datasets where $p\gg n$, such as genetics ones, which is the main motivation behind the proposal of SLOPE (Bogdan et al., 2015).
+The screening rules presented in this manuscript allow group-based SLOPE, and more generally group-based OWL models, to achieve computational fitting times more in line with their lasso-based counterparts.
+The screening rules also helped gSLOPE and SGS overcome convergence issues in large datasets, improving solution optimality.
+In our data studies, we have not discovered any scenario where our screening rules did not perform better than no screening.
+In each case, the rules greatly reduce the input dimensionality and speed up the computational runtime.
+Limitations
+Our screening rules, as any strong rule, rely on assumptions.
+For both gSLOPE and SGS, Lipschitz assumptions were used that are consistent with those used in the strong screening framework (Tibshirani et al., 2010).
+For gSLOPE, a Lipschitz assumption was used to derive Proposition 3.3, while for SGS, a separate Lipschitz assumption was made for each layer of screening (Propositions 4.2 and 4.3).
+Violations of these assumptions are checked for using the KKT conditions.
+The SGS KKT checks (Appendix B.4) led to an increased number of violations, as the checks were overly conservative.
+However, the overall amount of KKT violations for both models was still relatively small, suggesting that these assumptions are not overly restrictive.
+An attempt was made to derive less conservative checks (Appendix B.4.3), performed directly on the variables without an initial group check.
+These were found to be too lenient, incorrectly missing violations.
+Future work can consider alternative KKT checks.
+An additional future direction includes the development of safe rules, which guarantee that only inactive variables are discarded.
+These could be incorporated into a hybrid scheme together with our strong rules (Zeng et al., 2021; Wang and Breheny, 2022).
+Deriving safe rules would facilitate a comparison between them and our proposed strong rules, offering further insight into which type of screening is most effective for both non-separable and sparse-group norms.
+Acknowledgements
+We would like to thank the anonymous reviewers for their valuable comments.
+This work was supported by the Engineering and Physical Sciences Research Council (EPSRC) through the Modern Statistics and Statistical Machine Learning (StatML) CDT programme, grant no.
+EP/S023151/1.
+Checklist
+  1. 1.
+    For all models and algorithms presented, check if you include:
+    1. (a)
+      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes, we have defined our approach and the problem setting.]
+    2. (b)
+      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Yes, we have described the complexity of the fitting algorithm.]
+    3. (c)
+      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [Yes]
+  2. 2.
+    For any theoretical claim, check if you include:
+    1. (a)
+      Statements of the full set of assumptions of all theoretical results. [Yes. We have presented our assumptions in the results and have discussed the limitations of these in the discussion.]
+    2. (b)
+      Complete proofs of all theoretical results. [Yes. Proofs have been provided in the Appendix for all of our results.]
+    3. (c)
+      Clear explanations of any assumptions. [Yes, we have extensively discussed our assumptions and any corresponding limitations.]
+  3. 3.
+    For all figures and tables that present empirical results, check if you include:
+    1. (a)
+      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [Yes, code has been provided to reproduce any figure.]
+    2. (b)
+      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Yes, we have given the full simulation set up in Table A1.]
+    3. (c)
+      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times).
+[Yes, we have described our statistical measures and our error bars.]
+    4. (d)
+      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). [Yes]
+  4. 4.
+    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
+    1. (a)
+      Citations of the creator If your work uses existing assets. [Yes, we have cited our data and model sources.]
+    2. (b)
+      The license information of the assets, if applicable. [Yes]
+    3. (c)
+      New assets either in the supplemental material or as a URL, if applicable. [Not Applicable]
+    4. (d)
+      Information about consent from data providers/curators. [Yes, this is included in the code supplementary materials.]
+    5. (e)
+      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Not Applicable]
+  5. 5.
+    If you used crowdsourcing or conducted research with human subjects, check if you include:
+    1. (a)
+      The full text of instructions given to participants and screenshots. [Not Applicable]
+    2. (b)
+      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Not Applicable]
+    3. (c)
+      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Not Applicable]
+Strong Screening Rules for Group-based SLOPE Models:
+Supplementary Materials
 Appendix A GROUP SLOPE
 A.1 Penalty Weights
-The penalty weights for gSLOPE were derived to provide group FDR-control under orthogonal designs [6].
+The penalty weights for gSLOPE were derived to provide group FDR-control under orthogonal designs (Brzyski et al., 2019).
 For the FDR-control parameter $q_{g}\in(0,1)$, they are given by (where the indexing corresponds to the sorted groups)
 $w_{i}^{\text{max}}=\max_{j=1,\dots,m}\left\{\frac{1}{\sqrt{p_{j}}}F^{-1}_{\chi_{p_{j}}}(1-q_{g}i/m)\right\},\;\text{for}\;i=1,\dots,m,$
-where $F_{\chi_{p_{j}}}$ is the cumulative distribution function of a $\chi$ distribution with $p_{j}$ degrees of freedom. A relaxation to this sequence is applied in [6], to give
+where $F_{\chi_{p_{j}}}$ is the cumulative distribution function of a $\chi$ distribution with $p_{j}$ degrees of freedom. A relaxtion to this sequence is applied in Brzyski et al. (2019) to give
 $w_{i}^{\text{mean}}=\overline{F}^{-1}_{\chi_{p_{j}}}(1-q_{g}i/m),\;\text{where}\;\overline{F}_{\chi_{p_{j}}}(x):=\frac{1}{m}\sum_{j=1}^{m}F_{\chi_{p_{j}}}(\sqrt{p_{j}}x).$ (9)
@@
 Figure A1: The gSLOPE weights, $w$, shown for Figure 4 for $p=500,m=100,q_{g}=0.05$.
-A.2 Theory
+A.2 SLOPE Subdifferential
+The subdifferential for SLOPE was derived in Larsson et al. (2020) and as it is a vital part of our arguments, it is reproduced here for ease of reference.
+Define the function $R:\mathbb{R}^{p}\rightarrow\mathbb{N}^{p}$ that returns the ranks of the absolute values of its input and the set $\mathcal{C}_{i}(\beta)=\{j\in\{1,\ldots,p\}\;:\;|\beta_{i}|=|\beta_{j}|\}$.
+Then, the subdifferential is given by (Larsson et al., 2020)
+$\partial J_{\text{slope}}(\beta;v)=\begin{cases}\left\{x\in\mathbb{R}^{|\mathcal{C}_{i}|}:\text{cumsum}(|x|_{\downarrow}-v_{R(x)_{\mathcal{C}_{i}}})\preceq 0\right\}&\text{if}\;\beta_{\mathcal{C}_{i}}=0,\\ \Big{\{}x\in\mathbb{R}^{|\mathcal{C}_{i}|}:\text{cumsum}(|x|_{\downarrow}-v_{R(x)_{\mathcal{C}_{i}}})\preceq 0\\ \;\;\;\;\;\;\text{and}\;\;\sum_{j\in\mathcal{C}_{i}}(|x_{j}|-v_{R(s)_{j}})=0\\ \;\;\;\;\;\;\text{and}\;\;\text{sign}(\beta_{\mathcal{C}_{i}})=\text{sign}(x)\Big{\}}&\text{otherwise.}\end{cases}$
+The primary use of the subdifferential in this manuscript is the zero condition (Equation 5).
+A.3 Theory
 Proof of Theorem 3.1.
-The proof is similar to that of Theorem 2.7 in [6], where the subdifferential of gSLOPE is derived under equal groups.
+The proof is similar to that of Theorem 2.7 in Brzyski et al.
+(2019), where the subdifferential of gSLOPE is derived under equal groups.
 It is derived here under more general terms.
@@
     Active groups, $\mathcal{G}_{\mathcal{A}}$.
-Case 1: For inactive groups, we consider the subdifferential at zero. The subdifferential of a norm at zero is given by the dual norm of the unit ball [27],
+Case 1: For inactive groups, we consider the subdifferential at zero. The subdifferential of a norm at zero is given by the dual norm of the unit ball (Schneider and Tardivel, 2022),
 $\partial J_{\text{gslope}}(\mathbf{0};w)=\mathbf{B}_{J_{\text{gslope}}^{*}(\mathbf{0};w)}[0,1]=\{x:J_{\text{gslope}}^{*}(x;w)\leq 1\}.$
-The dual norm for gSLOPE is given by [6]
+The dual norm for gSLOPE is given by (Brzyski et al., 2019)
 $J_{\text{gslope}}^{*}(x;w)=J_{\text{slope}}^{*}([x]_{\mathcal{G},-0.5}).$
@@
 $\mathbf{B}_{J_{\text{gslope}}^{*}(\mathbf{0};w)}[0,1]=\{x:[x]_{\mathcal{G},-0.5}\in\mathbf{B}_{J_{\text{slope}}^{*}(\mathbf{0};w)}[0,1]\},$
-where $\mathbf{B}_{J_{\text{slope}}^{*}(\mathbf{0};w)}[0,1]=\{x\in\mathbb{R}^{m}:\text{cumsum}(|x|_{\downarrow}-w)\preceq\mathbf{0}\}$ is the unit ball of the dual norm to $J_{\text{slope}}$ [4].
+where $\mathbf{B}_{J_{\text{slope}}^{*}(\mathbf{0};w)}[0,1]=\{x\in\mathbb{R}^{m}:\text{cumsum}(|x|_{\downarrow}-w)\preceq\mathbf{0}\}$ is the unit ball of the dual norm to $J_{\text{slope}}$ (Bogdan et al., 2015).
 Using this, the subdifferential at zero for the inactive groups, $\mathcal{Z}$, is given by
@@
 $\sqrt{p_{g}}w_{g}\|\beta^{(g)}+d^{(g)}\|_{2}\geq\sqrt{p_{g}}w_{g}\|\beta^{(g)}\|_{2}+x^{(g)T}d^{(g)}.$
-Since the set $\{d^{(g)}:d\in D_{g}\}$ is open in $\mathbb{R}^{p_{g}}$ and contains zero, by Corollary G.1 in [6], it follows that $x^{(g)}\in\partial f_{g}(b^{(g)})$ for $f_{g}:\mathbb{R}^{p_{g}}\rightarrow\mathbb{R},f_{g}(x)=w_{g}\sqrt{p_{g}}\|x\|_{2}$.
+Since the set $\{d^{(g)}:d\in D_{g}\}$ is open in $\mathbb{R}^{p_{g}}$ and contains zero, by Corollary G.1 in Brzyski et al.
+(2019), it follows that $x^{(g)}\in\partial f_{g}(b^{(g)})$ for $f_{g}:\mathbb{R}^{p_{g}}\rightarrow\mathbb{R},f_{g}(x)=w_{g}\sqrt{p_{g}}\|x\|_{2}$.
 Now, for $g\leq s$, $f_{g}$ is differentiable in $\beta^{(g)}$, giving
@@
 Proof of Proposition 3.2.
-Suppose we have $\mathcal{B}\neq\emptyset$ after running the algorithm. Then, plugging in $h(\lambda_{k+1})=([\nabla f(\hat{\beta}(\lambda_{k+1}))]_{\mathcal{G},-0.5})_{\downarrow}$ gives
+Suppose we have $\mathcal{B}\neq\varnothing$ after running the algorithm. Then, plugging in $h(\lambda_{k+1})=([\nabla f(\hat{\beta}(\lambda_{k+1}))]_{\mathcal{G},-0.5})_{\downarrow}$ gives
 $\text{cumsum}\Bigl{(}\bigl{(}([\nabla f(\hat{\beta}(\lambda_{k+1}))]_{\mathcal{G},-0.5})_{\downarrow}\bigr{)}_{\mathcal{B}}-\lambda_{k+1}w_{\mathcal{B}}\Bigr{)}\prec\mathbf{0},$
@@
 Proof of Proposition 3.3.
-Since $\text{cumsum}(y)\succeq\text{cumsum}(x)\iff y\succeq x$ [17], we only need to show for a group $g$,
+Since $\text{cumsum}(y)\succeq\text{cumsum}(x)\iff y\succeq x$ (Larsson et al., 2020), we only need to show for a group $g$,
 $|h_{g}(\lambda_{k+1})|\leq|h_{g}(\lambda_{k})|+\lambda_{k}w_{g}-\lambda_{k+1}w_{g}.$
@@
 proving the result. ∎
-A.3 Path start derivation
+A.4 KKT Checks
+To check whether a group has been correctly discarded during the screening step, the KKT conditions for gSLOPE are checked. They are given by
+$\mathbf{0}$ $\in\nabla f(\beta)+\lambda\partial J_{\text{gslope}}(\beta;w)$
+$\implies-\nabla f(\beta)$ $\in\lambda\partial J_{\text{gslope}}(\beta;w).$
+Hence, we are checking whether the gradient of the loss function sits within the set of the gradient of the penalty.
+As we are only interested in identifying incorrectly discarded groups, we require only to check the subdifferential condition at zero.
+Hence, a violation occurs if a group is discarded but
+$-\nabla f(\beta)\notin\lambda\partial J_{\text{gslope}}(\mathbf{0};w_{\mathcal{G}_{\mathcal{Z}}})$
+$\implies$ $-\nabla f(\beta)\notin\left\{x\in\mathbb{R}^{\operatorname{card}\mathcal{G}_{\mathcal{Z}}}:[x]_{\mathcal{G}_{\mathcal{Z}},-0.5}\in\partial J_{\text{slope}}(0;\lambda w_{\mathcal{G}_{\mathcal{Z}}})\right\}$
+$\implies$ $[\nabla f(\beta)]_{\mathcal{G}_{\mathcal{Z}},-0.5}\notin\partial J_{\text{slope}}(0;\lambda w_{\mathcal{G}_{\mathcal{Z}}})$
+$\implies$ $\text{cumsum}(([\nabla f(\beta)]_{\mathcal{G}_{\mathcal{Z}},-0.5})_{\downarrow}-\lambda w_{\mathcal{G}_{\mathcal{Z}}})\succ 0.$
+A.5 Path Start Proof
+Proof of Proposition 3.4.
 The aim is to find the value of $\lambda$ at which the first group enters the model. When all features are zero, the gSLOPE KKT conditions (Equation 2) are
@@
 Rearranging this gives
-$\lambda\succeq\text{cumsum}\bigl{(}([\nabla f(\mathbf{0})]_{\mathcal{G},-0.5})_{\downarrow}\bigr{)}\oslash\text{cumsum}(w),$
-[…]
+$\lambda\succeq\text{cumsum}\bigl{(}([\nabla f(\mathbf{0})]_{\mathcal{G},-0.5})_{\downarrow}\bigr{)}\oslash\text{cumsum}(w).$
+Picking the maximum possible $\lambda$ such that this holds yields
 $\lambda_{1}=\max\left\{\text{cumsum}\bigl{(}([\nabla f(\mathbf{0})]_{\mathcal{G},-0.5})_{\downarrow}\bigr{)}\oslash\text{cumsum}(w)\right\}.$
-This can be verified by noting that $\lambda_{1}=J^{*}_{\text{gslope}}(\nabla f(\mathbf{0});w)$ [22].
-Now, $J^{*}_{\text{gslope}}(x;w)=J^{*}_{\text{slope}}([x]_{\mathcal{G},-0.5};w)$ [6].
-The dual norm of SLOPE is given by [23]
+This can be verified by noting that $\lambda_{1}=J^{*}_{\text{gslope}}(\nabla f(\mathbf{0});w)$ (Ndiaye et al., 2016b).
+Now, $J^{*}_{\text{gslope}}(x;w)=J^{*}_{\text{slope}}([x]_{\mathcal{G},-0.5};w)$ (Brzyski et al., 2019).
+The dual norm of SLOPE is given by (Negrinho and Martins, 2014)
 $J^{*}_{\text{slope}}(x;w)=\max\left\{\text{cumsum}(|x|_{\downarrow})\oslash\text{cumsum}(w)\right\}.$
-Therefore, $\lambda_{1}$ is as before.
+Therefore, $\lambda_{1}$ is as before. ∎
 Appendix B SPARSE-GROUP SLOPE
 B.1 Penalty Weights
-The penalty weights for SGS provide variable and group FDR-control simultaneously, under orthogonal designs [14]. They are given by (where the indexing corresponds to the sorted variables/groups)
-$v_{i}^{\text{max}}=\max_{j=1,\dots,m}\left\{\frac{1}{\alpha}F_{\mathcal{N}}^{-1}\left(1-\frac{q_{v}i}{2p}\right)-\frac{1}{3\alpha}(1-\alpha)a_{j}w_{j}\right\},\;i=1,\dots,p,$ (11)
-$w_{i}^{\text{max}}=\max_{j=1,\dots,m}\left\{\frac{F_{\text{FN}}^{-1}(1-\frac{q_{g}i}{m})-\alpha\sum_{k\in\mathcal{G}_{j}}v_{k}}{(1-\alpha)p_{j}}\right\},\;i=1,\dots,m,$ (12)
+The penalty weights for SGS provide variable and group FDR-control simultaneously, under orthogonal designs (Feser and Evangelou, 2023).
+They are given by (where the indexing corresponds to the sorted variables/groups)
+$v_{i}^{\text{max}}=\max_{j=1,\dots,m}\left\{\frac{1}{\alpha}F_{\mathcal{N}}^{-1}\left(1-\frac{q_{v}i}{2p}\right)-\frac{1}{3\alpha}(1-\alpha)a_{j}w_{j}\right\},\;i=1,\dots,p,$
+$w_{i}^{\text{max}}=\max_{j=1,\dots,m}\left\{\frac{F_{\text{FN}}^{-1}(1-\frac{q_{g}i}{m})-\alpha\sum_{k\in\mathcal{G}_{j}}v_{k}}{(1-\alpha)p_{j}}\right\},\;i=1,\dots,m,$
 where $F_{\chi_{p_{j}}}$ is the cumulative distribution function of a $\chi$ distribution with $p_{j}$ degrees of freedom, $F_{\mathcal{N}}$ is the cumulative distribution function of a folded Gaussian distribution, and $a_{j}$ is a quantity that requires estimation.
-The estimator $\hat{a}_{j}=\lfloor\alpha p_{j}\rfloor$ is proposed in [14].
-As with gSLOPE (Appendix A.1), a relaxation is possible, giving the weights
+The estimator $\hat{a}_{j}=\lfloor\alpha p_{j}\rfloor$ is proposed in Feser and Evangelou (2023).
+As with gSLOPE (Appendix A.1), a relaxtion is possible, giving the weights
 $v_{i}^{\text{mean}}=\overline{F}_{\mathcal{N}}^{-1}\left(1-\frac{q_{v}i}{2p}\right),\;\text{where}\;\overline{F}_{\mathcal{N}}(x):=\frac{1}{m}\sum_{j=1}^{m}F_{\mathcal{N}}\left(\alpha x+\frac{1}{3}(1-\alpha)a_{j}w_{j}\right),$ (11)
 $w_{i}^{\text{mean}}=\overline{F}_{\text{FN}}^{-1}\left(1-\frac{q_{g}i}{p}\right),\;\text{where}\;\overline{F}_{\text{FN}}(x):=\frac{1}{m}\sum_{j=1}^{m}F_{\text{FN}}\left((1-\alpha)p_{j}x+\alpha\sum_{k\in\mathcal{G}_{j}}v_{k}\right).$ (12)
-The mean weights defined in Equations 13 and 14 are used for all SGS numerical simulations in this manuscript (shown in Figure A2).
+In the manuscript, as recommended by Feser and Evangelou (2023) under general settings, the SGS variable mean (Equation 11) and gSLOPE group mean (Equation 9) weights are used for all SGS numerical simulations.
 Figure A2: The SGS weights, $(v,w)$, shown for Figure 4 for $p=500,m=100,q_{v}=0.05,q_{g}=0.05,\alpha=0.95$.
 B.2 Derivation of Soft Thresholding Operator
+Proof of Lemma 4.1.
 To determine the form of the quantity $\partial J_{\text{slope}}(\mathbf{0};v)$, consider that for Equation 7 to be satisfied, the term inside the []delimited-[][$\cdot$][ ⋅ ] operator needs to be as small as possible.
@@
   3. 3.
-    $\nabla_{i}f(\beta)\in[-\lambda\alpha v_{i},\lambda\alpha v_{i}]$: choose $y_{i}=\nabla_{i}f(\beta)/\alpha\lambda$.
+    $\nabla_{i}f(\beta)\in[-\lambda\alpha v_{i},\lambda\alpha v_{i}]$: choose $y_{i}=\nabla_{i}f(\beta)/\lambda\alpha v_{i}$.
 Hence, the term becomes
-$S(\nabla f(\beta),\lambda\alpha):=\text{sign}(\nabla f(\beta))(|\nabla f(\beta)|-\lambda\alpha)_{+},$
-which is the soft thresholding operator.
+$S(\nabla f(\beta),\lambda\alpha v):=\text{sign}(\nabla f(\beta))(|\nabla f(\beta)|-\lambda\alpha v)_{+},$
+which is the soft thresholding operator. ∎
 B.3 Theory
 Proposition B.1 (Strong group screening rule for SGS).
-Let $\tilde{h}(\lambda):=([S(\nabla f(\beta),\lambda\alpha)]_{\mathcal{G},-0.5})_{\downarrow}$.
+Let $\tilde{h}(\lambda):=([S(\nabla f(\beta),\lambda\alpha v)]_{\mathcal{G},-0.5})_{\downarrow}$.
 Then taking $c=\tilde{h}(\lambda_{k+1})$ and $\phi=(1-\alpha)\lambda_{k+1}w$ as inputs for Algorithm A1 returns a superset $\mathcal{S}_{g}(\lambda_{k+1})$ of the active set $\mathcal{A}_{g}(\lambda_{k+1})$.
 Proof of Proposition B.1.
-The proof is similar to that of Proposition 3.2. Suppose we have $\mathcal{B}\neq\emptyset$ after running the algorithm. Then,
+The proof is similar to that of Proposition 3.2. Suppose we have $\mathcal{B}\neq\varnothing$ after running the algorithm. Then,
 $\text{cumsum}(\tilde{h}_{\mathcal{B}}(\lambda_{k+1})-\lambda_{k+1}(1-\alpha)w_{\mathcal{B}})$ $\prec\mathbf{0}$
-$\implies\text{cumsum}\Bigl{(}\bigl{(}[S(\nabla f(\beta),\lambda_{k+1}\alpha)]_{\mathcal{G},-0.5})_{\downarrow}\bigr{)}_{\mathcal{B}}-\lambda_{k+1}(1-\alpha)w_{\mathcal{B}}\Bigr{)}$ $\prec,\mathbf{0}$
-so that by the SGS subdifferential (Equation 7), all groups in $\mathcal{B}$ are inactive. Hence, $\mathcal{S}_{g}(\lambda_{k+1})$ will contain the active set $\mathcal{A}_{g}(\lambda_{k+1})$. ∎
+$\implies\text{cumsum}\Bigl{(}\bigl{(}[S(\nabla f(\beta),\lambda_{k+1}\alpha v)]_{\mathcal{G},-0.5})_{\downarrow}\bigr{)}_{\mathcal{B}}-\lambda_{k+1}(1-\alpha)w_{\mathcal{B}}\Bigr{)}$ $\prec\mathbf{0},$
+so that by the SGS subdifferential (Equation 7) all groups in $\mathcal{B}$ are inactive. Hence, $\mathcal{S}_{g}(\lambda_{k+1})$ will contain the active set $\mathcal{A}_{g}(\lambda_{k+1})$. ∎
 Proof of Proposition 4.2.
@@
 Proof.
-Suppose we have $\mathcal{B}\neq\emptyset$ after running the algorithm. Then, we have
-$\text{cumsum}(\bar{h}_{\mathcal{B}}(\lambda_{k+1})-\lambda_{k+1}\alpha v_{\mathcal{B}})\prec\mathbf{0}\implies\text{cumsum}\Bigl{(}\bigl{(}|\nabla f(\hat{\beta}(\lambda_{k+1}))|_{\downarrow}\bigr{)}_{\mathcal{B}}-\lambda_{k+1}\alpha v_{\mathcal{B}}\Bigr{)}\prec\mathbf{0},$ (15)
-so that by the SGS subdifferential for non-zero groups (Equation 8), all variables in $\mathcal{B}$ are inactive.
+Suppose we have $\mathcal{B}\neq\varnothing$ after running the algorithm. Then, we have
+$\text{cumsum}(\bar{h}_{\mathcal{B}}(\lambda_{k+1})-\lambda_{k+1}\alpha v_{\mathcal{B}})\prec\mathbf{0}\implies\text{cumsum}\Bigl{(}\bigl{(}|\nabla f(\hat{\beta}(\lambda_{k+1}))|_{\downarrow}\bigr{)}_{\mathcal{B}}-\lambda_{k+1}\alpha v_{\mathcal{B}}\Bigr{)}\prec\mathbf{0},$
+so that by the SGS subdifferential for non-zero groups (Equation 8) all variables in $\mathcal{B}$ are inactive.
 Hence, $\mathcal{S}_{v}(\lambda_{k+1})$ will contain the active set $\mathcal{A}_{v}(\lambda_{k+1})$.
@@
 ∎
-B.4 Path start derivation
+B.4 KKT Checks
+For SGS, the KKT conditions are first checked at the group-level for inactive groups (Appendix B.4.1).
+Further variable checks are performed for violating groups and variables in active groups (to check whether the variables should also be active) (Appendix B.4.2).
+The violating variables from these secondary variable checks are added back into $\mathcal{E}_{v}$.
+B.4.1 Group Checks
+A group violation occurs if the KKT conditions do not hold at the group-level (Equation 7). That is, a violation occurs if a group is discarded but
+$\text{cumsum}\Bigl{(}\bigl{(}[\nabla f(\beta)+\lambda\alpha\partial J_{\text{slope}}(\mathbf{0};v)]_{\mathcal{G}_{\mathcal{Z}},-0.5}\bigr{)}_{\downarrow}-\lambda(1-\alpha)w_{\mathcal{Z}}\Bigr{)}\succ\mathbf{0}.$
+B.4.2 Variable Checks
+For the set of variables in a violating group (from Appendix B.4.1), denoted $\mathcal{G}_{\mathcal{K}_{g}}$, a variable violation occurs if Equation 8 does not hold. That is, if
+$\nabla_{\mathcal{G}_{\mathcal{K}_{g}}}f(\beta)\notin\lambda\alpha\partial J_{\text{slope}}(\mathbf{0};v_{\mathcal{G}_{\mathcal{K}_{g}}})\implies\text{cumsum}(|\nabla_{\mathcal{G}_{\mathcal{K}_{g}}}f(\beta)|-\lambda\alpha v_{\mathcal{G}_{\mathcal{K}_{g}}})\succ 0.$
+B.4.3 Alternative KKT Checks
+An alternative approach for SGS is to check the KKT conditions directly on the variables. The KKT conditions (Equation 6) can be rewritten as
+$-\nabla f(\beta)-\lambda(1-\alpha)\partial J_{\text{gslope}}(\beta;w)\in\lambda\alpha\partial J_{\text{slope}}(\beta;v).$
+A KKT violation occurs the zero subdifferential conditions are not satisfied
+$-\nabla f(\beta)-\lambda(1-\alpha)\partial J_{\text{gslope}}(\beta;w)\notin\lambda\alpha\partial J_{\text{slope}}(\mathbf{0};v)$
+$\implies$ $\text{cumsum}\left(\left|\nabla f(\beta)+\lambda(1-\alpha)\partial J_{\text{gslope}}(\beta;w)\right|_{\downarrow}-\lambda\alpha v\right)\succ\mathbf{0}.$
+Now, the objective is to make the term inside the sorted absolute value operator as small as possible, given that the subdifferential term is unknown.
+To do this, a similar derivation as in Section B.2 can be used to determine that the term must be the soft thresholding operator, so that a violation occurs if
+$\text{cumsum}\left(\left|S(\nabla f(\beta),\lambda(1-\alpha)\tau\omega)\right|_{\downarrow}-\lambda\alpha v\right)\succ\mathbf{0},$
+where $\tau$ and $\omega$ are expanded vectors of the group sizes ($\sqrt{p_{g}})$ and penalty weights ($w_{g}$) to $p$ dimensions, so that each variable within the same group is assigned the same value.
+However, as we have had to approximate the unknown subdifferential term, this check is not exact.
+In practice, we found that this check was not stringent enough (due to the approximation), leading to violations being missed.
+B.5 Path Start Proof
+Proof of Proposition 4.4.
 The aim is to find the value of $\lambda$ at which the first variable enters the model. When all features are zero, the SGS KKT conditions (Equation 6) are
@@
 $\text{cumsum}([x]_{\mathcal{G},-0.5}-w)\preceq\mathbf{0}\implies[x]_{\mathcal{G},-0.5}\preceq w\implies\|x^{(g)}\|_{2}\leq\sqrt{p_{g}}w_{g},\forall g\in\mathcal{G}.$
-This is satisfied at the upper limit at $x=\tau\omega$, where $\tau$ and $\omega$ are expanded vectors of the group sizes ($\sqrt{p_{g}})$ and penalty weights ($w_{g}$) to $p$ dimensions, so that each variable within the same group is assigned the same value.
-Hence,
+This is satisfied at the upper limit at $x=\tau\omega$. Hence,
 $\lambda_{1}=\max\left\{\text{cumsum}(|\nabla f(\mathbf{0})|_{\downarrow})\oslash\text{cumsum}((1-\alpha)\tau\omega-\alpha v)\right\}.$
 Appendix C SLOPE ALGORITHM
-Algorithm A1 Cumsum algorithm from [17]
+Algorithm A1 SLOPE subdifferential algorithm from Larsson et al. (2020)
   Input: $c\in\mathbb{R}^{p},\phi\in\mathbb{R}^{p}$, where ϕ1subscriptitalic-$\phi_{1}\geq\cdots\geq$ $\phi_{p}\geq 0$
@@
   Output: $\mathcal{S}$
-Appendix D Screening rule framework
+Appendix D SCREEENING RULE FRAMEWORK
 D.1 Group SLOPE Algorithm
@@
 Appendix E GROUP-BASED OSCAR
-This section presents an extension of the proposed screening rules for group-based OSCAR models.
-Firstly the OSCAR models are presented, and subsequently, the performance of the screening rules on synthetic data is presented.
-E.1 Model definition
-The Ordered Weighted $\ell_{1}$ (OWL) framework is defined as [35]
-$\hat{\beta}=\operatorname*{arg\,min}_{\beta}\left\{\nabla f(\beta)+\lambda J_{\text{owl}}(\beta;v)\right\},$
-where $J_{\text{owl}}(\beta;v)=\sum_{i=1}^{p}v_{i}|\beta|_{(i)}$, $|\beta|_{(1)}\geq\ldots\geq|\beta|_{(p)}$, and $v$ are non-negative non-increasing weights.
-SLOPE is a special case of OWL where the weights are taken to be the Benjamini-Hochberg critical values [4].
-Octagonal Shrinkage and Clustering Algorithm for Regression (OSCAR) [5] is a further special case of OWL (often referred to as OWL with linear decay) where for a variable $i$, the weights are taken to be $v_{i}=\sigma_{1}+\sigma_{2}(p-i)$, and $\sigma_{1},\sigma_{2}$ are to be set.
-In Bao et al.
-[3] they are set to $\sigma_{1}=d_{i}\|\mathbf{X}^{\top}y\|_{\infty},\sigma_{2}=\sigma_{1}/p$, where $d_{i}=i\times e^{-2}$.
-Group OSCAR (gOSCAR) and Sparse-group OSCAR (SGO) are defined using the frameworks provided by gSLOPE [6] and SGS [14], respectively, but instead using the weights (for a variable $i$ and a group $g$)
+This section provides supplementary materials for extending the proposed screening rules to group-based OSCAR models.
+E.1 Penalty Sequence
+The gOSCAR and SGO weights are defined by (for a variable $i\in[p]$ and group $g\in[m]$) (Figure A3)
 $v_{i}=\sigma_{1}+\sigma_{2}(p-i),\;w_{g}=\sigma_{1}+\sigma_{3}(m-g),\;\sigma_{3}=\sigma_{1}/m.$ (13)
-with the weights visualised in Figure A3.
-Figure A3: The SGO weights, $(v,w)$, shown for Figure 4 for $p=500,m=100,q_{v}=0.05,q_{g}=0.05,\alpha=0.95$.
+Figure A3: The SGO weights, $(v,w)$, for $p=500,m=100,q_{v}=0.05,q_{g}=0.05,\alpha=0.95$.
 E.2 Results
-Similar observations and conclusions made for the screening rules of gSLOPE and SGS are made for gOSCAR and SGO (Figures A4 - A8).
-Figure A4 illustrates the effectiveness of the bi-level screening of SGO, similar to the effectiveness observed for SGS.
-Figures A5 - A6 showcase the efficiency of the screening rules on the proportion of the selected groups/variables.
+Observations and conclusions made for the screening rules of gSLOPE and SGS are also found to be true for gOSCAR and SGO (Figures A4 - A8).
+Figure A4 illustrates the effectiveness of bi-selection of SGO, similar to the effectiveness observed for SGS.
+Figures A5 and A6 showcase the efficiency of the screening rules on the proportion of the selected groups/variables.
 The screening rules are found to be effective across different data characteristics, with the running time of the models significantly decreasing (Figure A7).
-Similarly to SGS, KKT violations are more common compared to the ones observed for gOSCAR (Figure A8).
-This is due to the additional assumptions made at the additional screening layer of SGO and SGS.
-Similarly to Figure 6(a), the shape of the increasing number of KKT violations mirrors the log-linear shape of the regularization path.
-Figure A4: The proportion of variables in $\mathcal{S}_{v}$ relative to the full input for SGO, shown for group and bi-level screening plotted as a function of the regularization path, applied to the synthetic data (Section 5.1).
+KKT violations for SGO are more common compared to gOSCAR (Figure A8), due to the additional assumptions made at the second screening layer (as with SGS).
+Similar to Figure 7, the shape of the increasing number of KKT violations mirrors the log-linear shape of the regularization path.
+Figure A4: The proportion of variables in $\mathcal{S}_{v}$ relative to the full input for SGO, shown for group and bi-level screening plotted as function of the regularization path, applied to the synthetic data (Section 6.1).
 The data are generated under a linear model for $p=500,5000$.
-The results are averaged over 100 repetitions and 95% confidence intervals are shown (the SGO equivalent of Figure 1 (a)).
+The results are averaged over 100 repetitions and 95% confidence intervals are shown (the SGO equivalent of Figure 1).
 Figure A5: The number of groups/variables in $\mathcal{E},\mathcal{A}$ for both gOSCAR and SGO as a function of the regularization path for the linear model with $p=2750,\rho=0.6,m=197$.
@@
 Figure A7: Runtime (in seconds) for fitting $50$ models along a path, shown for screening against no screening as a function of $p$, broken down into different correlation cases, for the linear model.
-The results are averaged over $100$ repetitions, with standard errors shown (the OSCAR equivalent of Figure 4).
+The results are averaged over $100$ repetitions, with standard errors shown (the gOSCAR/SGO equivalent of Figure 4).
 Figure A8: The proportion of KKT violations relative to the full input space, as a function of the regularization path.
-Group violations for gOSCAR and variable violations for SGS, fitted to linear models, averaged over all cases of $p$ and $\rho$ (the gOSCAR/SGO equivalent of Figure 6(a)).
+Group violations for gOSCAR and variable violations for SGS, under linear models, averaged over all cases of $p$ and $\rho$ (the gOSCAR/SGO equivalent of Figure 7).
 Appendix F RESULTS
@@
 The simulated experiments were executed on a high-performance computing cluster (x86-64 Linux GNU) and the real data analysis was conducted on a Apple Macbook Air (M1, 8GB).
-F.2 Model Comparison
+Code for all simulations is available in the Supplementary Material.
+For all models, ATOS was used with the algorithmic parameters given in Table A1.
+Table A1: Hyperparameters used for running ATOS in the synthetic and real data studies.
+Parameter Synthetic Data Real Data
+Max iterations 5000 10000
+Backtracking 0.7 0.7
+Max backtracking iterations 100 100
+Convergence tolerance $10^{-5}$ $10^{-5}$
+Convergence criteria $\|x-z\|_{2}$ $\|x-z\|_{2}$
+Standardization $\ell_{2}$ $\ell_{2}$
+Intercept Yes for linear Yes for linear
+Warm starts Yes Yes
+F.2 Solution Optimality
 This section presents the accuracy of the models with and without screening, by comparing the $\ell_{2}$ distances observed between the screened and non-screened fitted values.
-Synthetic data.
-For the linear model, the maximum $\ell_{2}$ distances observed between the screened and non-screened fitted values were of order $10^{-5}$ for gSLOPE and $10^{-8}$ for SGS (Table F.3).
+Synthetic Data
+For the linear model, the maximum $\ell_{2}$ distances observed between the screened and non-screened fitted values were of order $10^{-6}$ for gSLOPE and $10^{-9}$ for SGS (Table F.3).
 Across the different cases, 98000 models were fit in total for each approach (excluding the models for $\lambda_{1}$, where no screening is applied).
@@
 There was only one instance (out of the 98000) that this occurred for SGS, where $\mathcal{E}$ was missing a single variable contained in $\mathcal{A}$ (which had a non-screen fitted value of $\hat{\beta}=-0.004$).
-For the logistic model, the maximum $\ell_{2}$ distances observed between the screened and non-screened fitted values were of order $10^{-8}$ for both gSLOPE and SGS (Table F.3.2).
+For the logistic model, the maximum $\ell_{2}$ distances observed between the screened and non-screened fitted values were of order $10^{-8}$ for gSLOPE and $10^{-9}$ for SGS (Table F.3.2).
 Across the different cases, 98000 models were fit in total for each approach (excluding the models for $\lambda_{1}$, where no screening is applied).
 Of these model fits, there were no instances for gSLOPE or SGS where $\mathcal{E}$ was not a superset of $\mathcal{A}$.
-Real data.
-In the real data analysis, the estimated coefficients with and without screening were very close to each other (with $\ell_{2}$ to be of order $10^{-8}$) for both SGS and gSLOPE (Table F.4).
-No instances occurred where $\mathcal{E}$ was not a superset of $\mathcal{A}$.
+Real Data
+In the real data analysis, the estimated coefficients with and without screening were very close to each other for both SGS and gSLOPE (Table F.5).
+However, direct comparison is less meaningful here, as the models often failed to converge without screening, therefore not reaching the optimal solution.
 F.3 Additional Results from the Simulation Study
@@
 Standard errors are shown.
-Method Type Runtime (s)
-Runtime (s) screen
+Method Type Runtime
+Runtime screen (s)
 $\operatorname{card}(\mathcal{A}_{g})$ $\operatorname{card}(\mathcal{S}_{g})$ $\operatorname{card}(\mathcal{E}_{g})$ $\operatorname{card}(\mathcal{K}_{g})$ Num it
@@
 gSLOPE Logistic $814\pm 8$ $1409\pm 11$ $71\pm 1$ $97\pm 1$ $97\pm 1$ $0.014\pm 0.014$ $78\pm 1$ $83\pm 1$ $1\times 10^{-8}\pm 1\times 10^{-8}$
-SGS Linear $735\pm 15$ $1830\pm 34$ $61\pm 1$ $84\pm 1$ $91\pm 1$ $21\pm 1$ $91\pm 3$ $708\pm 12$ $2\times 10^{-9}\pm 3\times 10^{-9}$
-SGS Logistic $407\pm 2$ $859\pm 6$ $84\pm 1$ $107\pm 1$ $118\pm 1$ $22\pm 0.3$ $7\pm 0.2$ $51\pm 0.8$ $4\times 10^{-9}\pm 3\times 10^{-10}$
+SGS Linear $735\pm 15$ $1830\pm 34$ $61\pm 1$ $84\pm 1$ $91\pm 1$ $91\pm 3$ $708\pm 12$ $2\times 10^{-9}\pm 3\times 10^{-9}$
+SGS Logistic $407\pm 2$ $859\pm 6$ $84\pm 1$ $107\pm 1$ $118\pm 1$ $7\pm 0.2$ $51\pm 0.8$ $4\times 10^{-9}\pm 3\times 10^{-10}$
 F.3.1 Additional Results for the Linear Model
+Figure A9: The proportion of groups/variables in $\mathcal{E},\mathcal{A}$, relative to the full input, for gSLOPE and SGS, as a function of the regularization path for the linear model with $p=2750,\rho=0.6,m=197$.
+The results are averaged over $100$ repetitions.
 (a) $p=500,\rho=0$
@@
 Standard errors are shown.
-Method $p$ Runtime (s)
-Runtime (s) screen
+Method $p$ Runtime
+Runtime screen (s)
 $\operatorname{card}(\mathcal{A}_{g})$ $\operatorname{card}(\mathcal{S}_{g})$ $\operatorname{card}(\mathcal{E}_{g})$ $\operatorname{card}(\mathcal{K}_{g})$ Num it
@@
 gSLOPE $5000$ $1937\pm 63$ $2422\pm 66$ $102\pm 3$ $147\pm 4$ $147\pm 4$ $0.007\pm 0.012$ $594\pm 21$ $581\pm 19$ $2\times 10^{-7}\pm 3\times 10^{-7}$
-SGS $500$ $94\pm 1$ $133\pm 2$ $9\pm 0.2$ $11\pm 0.3$ $12\pm 0.3$ $3\pm 0.2$ $28\pm 1$ $74\pm 3$ $5\times 10^{-10}\pm 4\times 10^{-10}$
-SGS $1625$ $416\pm 8$ $1129\pm 19$ $25\pm 1$ $37\pm 1$ $41\pm 1$ $12\pm 0.3$ $62\pm 7$ $511\pm 21$ $2\times 10^{-9}\pm 2\times 10^{-9}$
-SGS $2750$ $639\pm 14$ $2137\pm 47$ $62\pm 2$ $82\pm 2$ $89\pm 3$ $20\pm 1$ $80\pm 6$ $791\pm 28$ $2\times 10^{-9}\pm 6\times 10^{-9}$
-SGS $3875$ $939\pm 31$ $2862\pm 96$ $93\pm 3$ $124\pm 3$ $136\pm 4$ $30\pm 1$ $112\pm 8$ $1049\pm 34$ $5\times 10^{-9}\pm 1\times 10^{-8}$
-SGS $5000$ $1586\pm 66$ $2891\pm 128$ $119\pm 4$ $164\pm 3$ $180\pm 4$ $39\pm 1$ $171\pm 11$ $1118\pm 37$ $2\times 10^{-9}\pm 4\times 10^{-9}$
+SGS $500$ $94\pm 1$ $133\pm 2$ $9\pm 0.2$ $11\pm 0.3$ $12\pm 0.3$ $28\pm 1$ $74\pm 3$ $5\times 10^{-10}\pm 4\times 10^{-10}$
+SGS $1625$ $416\pm 8$ $1129\pm 19$ $25\pm 1$ $37\pm 1$ $41\pm 1$ $62\pm 7$ $511\pm 21$ $2\times 10^{-9}\pm 2\times 10^{-9}$
+SGS $2750$ $639\pm 14$ $2137\pm 47$ $62\pm 2$ $82\pm 2$ $89\pm 3$ $80\pm 6$ $791\pm 28$ $2\times 10^{-9}\pm 6\times 10^{-9}$
+SGS $3875$ $939\pm 31$ $2862\pm 96$ $93\pm 3$ $124\pm 3$ $136\pm 4$ $112\pm 8$ $1049\pm 34$ $5\times 10^{-9}\pm 1\times 10^{-8}$
+SGS $5000$ $1586\pm 66$ $2891\pm 128$ $119\pm 4$ $164\pm 3$ $180\pm 4$ $171\pm 11$ $1118\pm 37$ $2\times 10^{-9}\pm 4\times 10^{-9}$
 F.3.2 Additional Results for the Logistic Model
@@
 The results are averaged over $100$ repetitions, with $95\%$ confidence intervals shown.
+This figure is the equivalent of Figure A10 for the logistic model.
 Figure A13: Runtime (in seconds) for screening against no screening as a function of $p$, broken down into different correlation cases, for the logistic model.
 The results are averaged over $100$ repetitions, with standard errors shown.
+This figure is the equivalent of Figure 4 for the logistic model.
 Table A6: Variable screening metrics for SGS using a logistic model for the simulation study presented in Section 6.1.
@@
 Standard errors are shown.
-Method $p$ Runtime (s)
-Runtime (s) screen
+Method $p$ Runtime
+Runtime screen (s)
 $\operatorname{card}(\mathcal{A}_{g})$ $\operatorname{card}(\mathcal{S}_{g})$ $\operatorname{card}(\mathcal{E}_{g})$ $\operatorname{card}(\mathcal{K}_{g})$ Num it
@@
 gSLOPE $5000$ $1454\pm 29$ $2727\pm 40$ $118\pm 3$ $168\pm 4$ $168\pm 4$ $0.022\pm 0.041$ $101\pm 2$ $109\pm 1$ $4\times 10^{-9}\pm 1\times 10^{-9}$
-SGS $500$ $118\pm 1$ $113\pm 1$ $28\pm 1$ $33\pm 1$ $38\pm 1$ $8\pm 0.2$ $6\pm 0.3$ $29\pm 1$ $8\times 10^{-9}\pm 7\times 10^{-10}$
-SGS $1625$ $248\pm 2$ $538\pm 9$ $50\pm 2$ $59\pm 2$ $64\pm 2$ $11\pm 0.4$ $7\pm 1$ $63\pm 3$ $5\times 10^{-9}\pm 1\times 10^{-9}$
-SGS $2750$ $374\pm 2$ $868\pm 12$ $85\pm 3$ $104\pm 2$ $115\pm 3$ $21\pm 1$ $7\pm 0.4$ $57\pm 2$ $3\times 10^{-9}\pm 5\times 10^{-10}$
-SGS $3875$ $558\pm 4$ $1280\pm 19$ $116\pm 3$ $148\pm 3$ $164\pm 3$ $30\pm 1$ $8\pm 0.4$ $54\pm 1$ $2\times 10^{-9}\pm 2\times 10^{-10}$
-SGS $5000$ $737\pm 5$ $1498\pm 19$ $141\pm 4$ $188\pm 3$ $209\pm 4$ $40\pm 1$ $8\pm 0.4$ $54\pm 1$ $1\times 10^{-9}\pm 2\times 10^{-10}$
-F.4 Additional results from the real data analysis
-Table A7: Variable screening metrics for SGS applied to real data in 5.2.
+SGS $500$ $118\pm 1$ $113\pm 1$ $28\pm 1$ $33\pm 1$ $38\pm 1$ $6\pm 0.3$ $29\pm 1$ $8\times 10^{-9}\pm 7\times 10^{-10}$
+SGS $1625$ $248\pm 2$ $538\pm 9$ $50\pm 2$ $59\pm 2$ $64\pm 2$ $7\pm 1$ $63\pm 3$ $5\times 10^{-9}\pm 1\times 10^{-9}$
+SGS $2750$ $374\pm 2$ $868\pm 12$ $85\pm 3$ $104\pm 2$ $115\pm 3$ $7\pm 0.4$ $57\pm 2$ $3\times 10^{-9}\pm 5\times 10^{-10}$
+SGS $3875$ $558\pm 4$ $1280\pm 19$ $116\pm 3$ $148\pm 3$ $164\pm 3$ $8\pm 0.4$ $54\pm 1$ $2\times 10^{-9}\pm 2\times 10^{-10}$
+SGS $5000$ $737\pm 5$ $1498\pm 19$ $141\pm 4$ $188\pm 3$ $209\pm 4$ $8\pm 0.4$ $54\pm 1$ $1\times 10^{-9}\pm 2\times 10^{-10}$
+F.4 Data Description
+  • carbotax: Carbotax study of ovarian tumor growth.
+    • Response (continuous): Relative tumor volume ($\log_{2}$ scale).
+    • Data matrix: Gene expression measurements. $10000$ factors were randomly sampled from a collection of $34964$.
+    • Grouping structure: Variables are grouped using k-means clustering (Lloyd, 1982).
+  • scheetz: Gene expression data in the mammalian eye.
+    • Response (continuous): Gene expression measurements for the Trim32 gene.
+    • Data matrix: Gene expression measurements for other genes.
+    • Grouping structure: Variables are grouped using k-means clustering (Lloyd, 1982).
+  • adenoma: Transcriptome profile data to identify formation of colorectal adenomas.
+    • Response (binary): Binary labels for whether sample came from adenoma or normal mucosa.
+    • Data matrix: Transcriptome profile measurements.
+    • Grouping structure: Genes are assigned to pathways (groups) using the C3 regulatory target gene sets.111gsea-msigdb.org/gsea/msigdb/human/collections.jsp. Accessed 08/2024.
+  • cancer: Breast cancer patients treated with tamoxifen for 5 years.
+    • Response (binary): Binary labels classifying whether the cancer had recurred.
+    • Data matrix: Gene expression data.
+    • Grouping structure: Genes are assigned to pathways (groups) using the C3 regulatory target gene sets.11footnotemark: 1
+  • celiac: Gene expression data of primary leucocytes to classify celiac disease.
+    • Response (binary): Binary labels as to whether a patient has celiac disease.
+    • Data matrix: Gene expression measurements from the primary leucocytes.
+    • Grouping structure: Genes are assigned to pathways (groups) using the C3 regulatory target gene sets.11footnotemark: 1
+  • colitis: Blood cells data for classifying whether a patient has colitis.
+    • Response (binary): Binary labels classifying whether a patient has colitis.
+    • Data matrix: Gene expression measurements.
+    • Grouping structure: Genes are assigned to pathways (groups) using the C3 regulatory target gene sets.11footnotemark: 1
+  • tumour: Gene expression data of pancreative cancer samples to identify tumorous tissue.
+    • Response (binary): Binary labels indicating if a sample is from tumour tissue.
+    • Data matrix: Gene expression measurements.
+    • Grouping structure: Genes are assigned to pathways (groups) using the C3 regulatory target gene sets.11footnotemark: 1
+Table A8: Dataset information for the six datasets used in the real data analysis.
+Dataset $p$ $n$ $m$ Group sizes Type Source
+carbotax $10000$ $101$ $100$ $[1,126]$ Linear Koussounadis et al. (2014)222downloaded from https://iowabiostat.github.io/data-sets/. Accessed 08/2024.
+scheetz $18975$ $120$ $379$ $[1,165]$ Linear Scheetz et al. (2006)0
+adenoma $17661$ $64$ $1849$ $[1,646]$ Logistic Sabates-Bellver et al. (2007)333downloaded from https://www.ncbi.nlm.nih.gov/. Accessed 09/2024.
+cancer $7057$ $60$ $1277$ $[1,292]$ Logistic Ma et al. (2004)0
+celiac $14294$ $132$ $1666$ $[1,570]$ Logistic Heap et al. (2009)0
+colitis $11999$ $127$ $1528$ $[1,497]$ Logistic Burczynski et al. (2006)0
+tumour $17661$ $52$ $1849$ $[1,646]$ Logistic Pei et al. (2009); Ellsworth et al. (2013); Li et al. (2016)0
+F.5 Additional Results from the Real Data Analysis
+Figure A14: The ratio of no screen time to screen time ($\uparrow$) of SLOPE, gSLOPE, and SGS applied to the real datasets, for fitting $100$ path models, split into response type.
+The horizontal grey line represents no screening improvement.
+Figure A15: The ratio of the fitting set ($\mathcal{E}_{v}$) to the input dimensionality ($p$) ($\downarrow$) of SLOPE, gSLOPE, and SGS applied to the real datasets, for fitting $100$ path models, split into response type.
+Figure A16: The ratio of the fitting set ($\mathcal{E}_{v}$) to the active set ($\mathcal{A}_{v}$) ($\downarrow$) of SLOPE, gSLOPE, and SGS applied to the real datasets, for fitting $100$ path models, split into response type.
+Figure A17: The proportion of variables in $\mathcal{S}_{v}$ relative to $p$ for group-only and bi-level screening applied to SGS, plotted along the regularization path for the carbotax, scheetz, adenoma, colitis, and tumour datasets.
+Table A9: Variable screening metrics for SGS applied to real data in Section 6.2.
 The number of variables in $\mathcal{A}_{v},\mathcal{S}_{v},\mathcal{E}_{v}$, and $\mathcal{K}_{v}$ are shown.
 The results are averaged across the nine pathway collections, with standard errors shown.
-Dataset Method Input
-$\operatorname{card}(\mathcal{A}_{v})$ $\operatorname{card}(\mathcal{S}_{v})$ $\operatorname{card}(\mathcal{E}_{v})$ $\operatorname{card}(\mathcal{K}_{v})$
-SGS Cancer $5651$ $71\pm 1$ $211\pm 4$ $212\pm 4$ $1\pm 0.04$
-SGS Colitis $10259$ $25\pm 1$ $208\pm 6$ $209\pm 6$ $1\pm 0.07$
-Table A8: General and group screening metrics for SGS and gSLOPE applied to real data in 5.2.
+Method Dataset $\operatorname{card}(\mathcal{A}_{v})$ $\operatorname{card}(\mathcal{S}_{v})$ $\operatorname{card}(\mathcal{E}_{v})$ $\operatorname{card}(\mathcal{K}_{v})$
+SGS carbotax $144\pm 6$ $361\pm 10$ $364\pm 10$ $2\pm 0.50$
+SGS scheetz $612\pm 74$ $1908\pm 174$ $1915\pm 175$ $7\pm 1.28$
+SGS adenoma $2174\pm 23$ $2451\pm 4$ $2463\pm 5$ $12\pm 0.22$
+SGS cancer $86\pm 3$ $273\pm 7$ $273\pm 7$ $1\pm 0.05$
+SGS celiac $175\pm 10$ $545\pm 22$ $546\pm 22$ $1\pm 0.07$
+SGS colitis $68\pm 4$ $409\pm 7$ $411\pm 7$ $2\pm 0.18$
+SGS tumour $128\pm 5$ $1189\pm 59$ $1200\pm 59$ $12\pm 0.54$
+Table A10: General and group screening metrics for SGS and gSLOPE applied to real data in Section 6.2.
 General metrics: the runtime (in seconds) for screening against no screening, the number of fitting iterations for screening against no screening (with the number of occasions of failed convergence given in brackets), and the $\ell_{2}$ distance between the fitted values obtained with screening and no screening.
@@
 The results are averaged across the nine pathway collections, with standard errors shown.
-Method Dataset Input
+Method Dataset Runtime
+Runtime screen (s)
 $\operatorname{card}(\mathcal{A}_{g})$ $\operatorname{card}(\mathcal{S}_{g})$ $\operatorname{card}(\mathcal{E}_{g})$ $\operatorname{card}(\mathcal{K}_{g})$ Num it
@@
 to no screen
-gSLOPE Cancer $592$ $66\pm 1$ $120\pm 2$ $120\pm 2$ $0\pm 0$ $1719\pm 63(16)$ $4453\pm 103(186)$ $7\times 10^{-8}\pm 7\times 10^{-9}$
-gSLOPE Colitis $655$ $32\pm 1$ $56\pm 2$ $56\pm 2$ $0\pm 0$ $1215\pm 73(28)$ $3904\pm 127(140)$ $3\times 10^{-7}\pm 5\times 10^{-8}$
-SGS Cancer $592$ $50\pm 1$ $167\pm 2$ $115\pm 2$ $67\pm 0.8$ $72\pm 3(0)$ $127\pm 4(0)$ $1\times 10^{-8}\pm 1\times 8^{-10}$
-SGS Colitis $655$ $18\pm 0.6$ $136\pm 4$ $87\pm 3$ $61\pm 2$ $156\pm 9(0)$ $1348\pm 62(0)$ $4\times 10^{-8}\pm 6\times 10^{-9}$
+gSLOPE carbotax $865$ $3456$ $71\pm 3$ $124\pm 2$ $124\pm 2$ $0.05\pm 0.05$ $2367\pm 176(1)$ $6433\pm 303(10)$ $1\times 10^{-9}\pm 2\times 10^{-10}$
+gSLOPE scheetz $4731$ $7735$ $153\pm 6$ $212\pm 6$ $212\pm 6$ $0\pm 0$ $6252\pm 313(20)$ $7183\pm 332(38)$ $8\times 10^{-10}\pm 8\times 10^{-10}$
+gSLOPE adenoma $415$ $12610$ $71\pm 6$ $77\pm 4$ $77\pm 4$ $0\pm 0$ $3956\pm 378(14)$ $7864\pm 394(78)$ $9\times 10^{-6}\pm 1\times 10^{-6}$
+gSLOPE cancer $109$ $5003$ $46\pm 3$ $81\pm 5$ $81\pm 5$ $0\pm 0$ $764\pm 74(0)$ $6250\pm 404(40)$ $4\times 10^{-7}\pm 8\times 10^{-8}$
+gSLOPE celiac $205$ $7530$ $34\pm 75$ $62\pm 9$ $62\pm 9$ $0\pm 0$ $834\pm 177(0)$ $4764\pm 428(27)$ $4\times 10^{-6}\pm 1\times 10^{-5}$
+gSLOPE colitis $148$ $11567$ $69\pm 5$ $114\pm 8$ $114\pm 8$ $0\pm 0$ $749\pm 62(0)$ $9090\pm 206(72)$ $4\times 10^{-6}\pm 7\times 10^{-7}$
+gSLOPE tumour $222$ $33405$ $24\pm 1$ $83\pm 4$ $83\pm 4$ $0\pm 0$ $892\pm 67(0)$ $9224\pm 218(86)$ $4\times 10^{-7}\pm 5\times 10^{-8}$
+SGS carbotax $63$ $2067$ $40\pm 1$ $138\pm 1$ $119\pm 1$ $168\pm 60(0)$ $3572\pm 350(3)$ $8\times 10^{-10}\pm 3\times 10^{-10}$
+SGS scheetz $358$ $5476$ $105\pm 4$ $203\pm 5$ $186\pm 6$ $10\pm 4(0)$ $5039\pm 360(28)$ $7\times 10^{-8}\pm 2\times 10^{-8}$
+SGS adenoma $1234$ $6143$ $601\pm 6$ $789\pm 1$ $735\pm 1$ $1172\pm 69(0)$ $4224\pm 338(20)$ $6\times 10^{-7}\pm 5\times 10^{-8}$
+SGS cancer $151$ $211$ $70\pm 2$ $295\pm 3$ $182\pm 4$ $124\pm 15(0)$ $147\pm 14(0)$ $2\times 10^{-8}\pm 2\times 10^{-9}$
+SGS celiac $339$ $2035$ $120\pm 7$ $460\pm 10$ $288\pm 10$ $55\pm 12(0)$ $1259\pm 112(0)$ $9\times 10^{-5}\pm 2\times 10^{-5}$
+SGS colitis $271$ $1821$ $52\pm 3$ $368\pm 3$ $230\pm 3$ $54\pm 10(0)$ $1411\pm 188(0)$ $1\times 10^{-6}\pm 3\times 10^{-7}$
+SGS tumour $781$ $7451$ $94\pm 3$ $733\pm 10$ $479\pm 14$ $33\pm 7(0)$ $6006\pm 287(23)$ $7\times 10^{-7}\pm 2\times 10^{-7}$